### v1

# On strong Euler-homogeneity and Saito-holonomicity for complex hypersurfaces

Abstract
Based on previous work on linear free divisors by Granger et al., we develop necessary and sufficient conditions for a general divisor to be strongly Euler-homogeneous in terms of some Fitting ideals.
We also define the notions of weak and strong Saito-holonomicity for a general divisor in such a way that they extend the definitions of Saito-holonomicity and weak and strong Koszul-freeness. Then, we characterize these properties using the Fitting ideals previously defined. Finally, we generalize some results regarding Koszul-freeness and strong Euler-homogeneity to the non-free case.
2020 Mathematics Subject Classification. 32S25, 32S05.
Keywords: logarithmic vector fields, Euler-homogeneity, Saito-holonomicity.
Declarations: The author is supported by a Junta de Andalucía PIF fellowship num. PREDOC_00485 and by PID2020-114613GB-I00. He has no further relevant interests to disclose.
Acknowledgements: The author would like to thank his PhD advisors Luis Narváez Macarro and Alberto Castaño Domínguez for their support and guidance while writing this paper.
1 Introduction
It is well-known that a polynomial $f\in\mathbb{C}[x_{1},\ldots,x_{n}]$ is quasihomogeneous of degree $d$ with respect to the weights $w_{1},\ldots,w_{n}\in\mathbb{Z}_{+}$ (that is, $f(x_{1}^{w_{1}},\ldots,x_{n}^{w_{n}})$ is homogeneous of degree $d$) if and only if it satisfies the Euler equation
$w_{1}x_{1}\frac{\partial f}{\partial x_{1}}+\ldots+w_{n}x_{n}\frac{\partial f}{\partial x_{n}}=df,$
which can be rewritten in terms of derivations as $\delta(f)=f$ with $\delta=\frac{1}{d}\sum_{i=1}^{n}w_{i}x_{i}\partial_{i}$.
If we consider holomorphic functions $f\in\mathbb{C}\{x_{1},\ldots,x_{n}\}$, we may ask which ones satisfy an equation of the form $\delta(f)=f$ with $\delta=\sum_{i=1}^{n}a_{i}\partial_{i}$ for some $a_{i}\in\mathbb{C}\{x_{1},\ldots,x_{n}\}$. This is what we call Euler-homogeneity.
Any quasihomogeneous polynomial is Euler-homogeneous but, clearly, there are holomorphic functions that are not polynomials and still verify this property (take for example $f=e^{x_{1}}$ with $\delta=\partial_{1}$). However, K. Saito proved in 1971 that, for isolated singularities, quasihomogeneous polynomials are in essence all the examples. Precisely, he proved that, if $f$ defines an isolated singularity in a neighbourhood of 00, then $f$ is Euler-homogeneous if and only if there exists a coordinate system in which $f$ becomes a quasihomogeneous polynomial [16, Theorem 4.1]. For non-isolated singularities, this is no longer true, even for polynomials. The divisor $x_{1}x_{2}(x_{1}+x_{2})(x_{1}+x_{2}x_{3})=0$ in $\mathbb{C}^{3}$, usually called the four lines, is a classical counterexample [3, Section 4].
Thus, it is interesting to see when this generalization of quasihomogeneity occurs. We are going to study this question from a geometric point of view, asking for those hypersurfaces having a local equation satisfying this property.
Let us consider a complex analytic manifold $X$ and let $D\subset X$ be a divisor (only reduced divisors, i.e. hypersurfaces, and reduced local equations of them will be considered, but sometimes we will omit this for conciseness). Denote by $\mathcal{O}_{X}$ the sheaf of holomorphic functions on $X$ and by $\operatorname{Der}_{X}$ the $\mathcal{O}_{X}$-module of $\mathbb{C}$-derivations.
Definition 1.1.
A germ of holomorphic function $f\in\mathcal{O}_{X,p}$ is called strongly Euler-homogeneous at $p$ if there exists a germ of derivation $\delta$ vanishing at $p$ (i.e. $\delta\in\mathfrak{m}_{X,p}\operatorname{Der}_{X,p}$, where $\mathfrak{m}_{X,p}$ denotes the maximal ideal of the local ring $\mathcal{O}_{X,p}$) such that $\delta(f)=f$. A divisor $D$ is said to be strongly Euler-homogeneous at $p\in D$ if some (or any) reduced local equation of $D$ at $p$ is strongly Euler-homogeneous and it is called strongly Euler-homogeneous on a subset $E\subset D$ if so is it at any $p\in E$. When $E=D$, we simply say that $D$ is strongly Euler-homogeneous.
Remark 1.2.
In the above definition, the word “strongly” refers to the extra condition that the germ of derivation vanishes at the given point. This is to ensure that the property is intrinsic to the divisor and does not depend on the choice of local equation.
In order to study this property, logarithmic derivations (or vector fields) along $D$ turn out to be very useful. This concept was extensively studied by K. Saito in [17] and refers to those derivations leaving invariant the defining ideal of $D$. These derivations form an $\mathcal{O}_{X}$-module, denoted by $\operatorname{Der}_{X}(-\log D)$. Locally, a derivation $\delta\in\operatorname{Der}_{X,p}$ belongs to $\operatorname{Der}_{X,p}(-\log D)$ if $\delta(f)\in(f)$ for some (or any) reduced local equation $f$ of $D$ at $p$. In this case, we will also say that $\delta$ is a logarithmic derivation for $f$.
Consider a reduced local equation $f$ of $D$ at $p$ and a generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$. Once a local coordinate system $x_{1},\ldots,x_{n}$ is chosen, we define the Saito matrix with respect to $\mathcal{S}$ as $A=(\delta_{i}(x_{j}))_{i,j}$, the $m\times n$ matrix whose entries are the coefficients in the expression of $\delta_{1},\ldots,\delta_{m}$ as linear combinations of $\partial_{1},\ldots,\partial_{n}$. Denoting by $\overline{\alpha}:=(\alpha_{1},\ldots,\alpha_{m})^{t}$, where $\alpha_{i}\in\mathcal{O}_{X,p}$ is such that $\delta_{i}(f)=\alpha_{i}f$, we define the extended Saito matrix with respect to $\mathcal{S}$ and $f$ as the $m\times(n+1)$ matrix $\tilde{A}=(A\mid-\overline{\alpha})$. Strictly speaking, these matrices also depend on the coordinate system, but we will omit this for the sake of brevity.
When $\operatorname{Der}_{X}(-\log D)$ is a locally free $\mathcal{O}_{X}$-module, we say that $D$ is a free divisor. If $D$ is free, then $\operatorname{Der}_{X,p}(-\log D)$ is free of rank $n$ as an $\mathcal{O}_{X,p}$-module for every $p\in X$. By Saito’s criterion [17, Theorem 1.8], the freeness of $\operatorname{Der}_{X,p}(-\log D)$ is equivalent to the existence of some $\delta_{1},\ldots,\delta_{n}\in\operatorname{Der}_{X,p}(-\log D)$ such that if $f\in\mathcal{O}_{X,p}$ is a reduced local equation at $p$, then there exists a unit $u\in\mathcal{O}_{X,p}$ with $f=u\det(A)$, where $A$ is the Saito matrix with respect to $\{\delta_{1},\ldots,\delta_{n}\}$. In this case, $\{\delta_{1},\ldots,\delta_{n}\}$ is automatically a basis of $\operatorname{Der}_{X,p}(-\log D)$.
In Section 2, inspired by [8, Lemma 7.5], we characterize strong Euler-homogeneity at a point $p\in D$ in terms of the rank of the matrices $A$ and $\tilde{A}$ evaluated at $p$ (Proposition 2.6). Then, we define some analytic closed subsets $D_{i}\subset X$ (resp. $\tilde{D}_{i}\subset X$) as the vanishing locus of some Fitting ideals and see that they can be explicitly described as the set of points at which the rank of $A$ (resp. $\tilde{A}$) is less or equal than $i$. This allows us to give a criterion for strong Euler-homogeneity in a neighbourhood of $p$ in terms of these sets. We also give the algebraic translation of this criterion and a necessary condition for a divisor to be strongly Euler-homogeneous outside $D_{0}$.
K. Saito defined in [17, Definition 3.8] holonomic divisors as those for which, at each point, the logarithmic stratification is locally finite and characterized them in terms of the dimensions of the sets $D_{i}$ (he called them $A_{r}$). Later, Calderón-Moreno defined in [2, Definition 4.1.1] the notion of Koszul-freeness for free divisors in terms of regular sequences in the graded ring $\operatorname{gr}\mathcal{D}_{X}$. Although this was not initially noticed, these two notions turned out to be the same for free divisors [8, Theorem 7.4]. In particular, Koszul-freeness can also be characterized in terms of these sets.
Still in the free setting, in [15, Definition 1.13] and [10, Remark 7.3], weak and strong Koszul-freeness were introduced (see Definition 3.1). While studying these properties, we asked ourselves if they could be extended to non-free divisors. This is what led us to define, by analogy, weak and strong Saito-holonomicity (Definition 3.3). And, surprisingly, we found out that they also have a characterization in terms of the dimensions of the sets $\tilde{D}_{i}$ (Theorem 3.6). This comprises the first half of Section 3.
The rest of Section 3 is devoted to study these properties and their relation with strong Euler-homogeneity and another interesting property: being of linear Jacobian type. In [15, Proposition 1.11 and Theorem 4.7], it was shown that being strongly Koszul-free and being of linear Jacobian type are equivalent for free divisors and that, under Koszul-freeness hypothesis, strong Koszul-freeness is equivalent to strong Euler-homogeneity (among other properties, including the so-called Logarithmic Comparison Theorem). We use the developed criteria of strong Euler-homogeneity and (strong) Saito-holonomicity to generalize, when possible, these results.
Finally, in Section 4, we extend the characterizations of strong Euler-homogeneity and (weak and strong) Saito-holonomicity, given in the convergent setting, to the ring of formal power series. The reason is that formal coordinate changes are much more flexible, as we do not have to care about convergence, and permits a particularly nice description of the module of formal logarithmic derivations, (by virtue of the formal structure theorem, see [9, Theorem 5.4]). This will be specially useful in [6] to advance in the comprehension of a conjecture posed in [3] stating that free divisors satisfying the Logarithmic Comparison Theorem are always strongly Euler-homogeneous.
2 Strong Euler-homogeneity around a point
For the sake of simplicity, the results will be developed for a germ of divisor in $(\mathbb{C}^{n},0)$, but they can be easily generalized for any divisor in a complex smooth manifold. So, from now on, let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$.
Let $(\mathcal{O},\mathfrak{m}):=(\mathcal{O}_{\mathbb{C}^{n},0},\mathfrak{m}_{\mathbb{C}^{n},0})$ be the local ring of convergent power series, which can be identified with $\mathbb{C}\{x_{1},\ldots,x_{n}\}$. Let $\operatorname{Der}:=\operatorname{Der}_{\mathbb{C}^{n},0}$ be the $\mathcal{O}$-module of $\mathbb{C}$-derivations of $\mathcal{O}$, which is free with a basis given by the partial derivatives $\{\partial_{1},\ldots,\partial_{n}\}$ and let $\operatorname{Der}_{f}:=\operatorname{Der}_{\mathbb{C}^{n},0}(-\log D)$ be the $\mathcal{O}$-module of logarithmic derivations for $f$. Recall that $\operatorname{Der}_{g}=\operatorname{Der}_{f}$ if $g$ is another reduced local equation of $D$ (i.e. $g=uf$ for some unit $u$).
We first introduce some analytic sets that will be used throughout the rest of the text and will play a crucial role in the characterization of strong Euler-homogeneity and Saito-holonomicity properties.
Let $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ be a generating set of $\operatorname{Der}_{f}$ (note that $m\geq n$ by coherence, since at a smooth point there cannot be less than $n$ generators). Let $A=(\delta_{i}(x_{j}))_{i,j}$ be the Saito matrix with respect to $\mathcal{S}$. Let $\alpha_{i}\in\mathcal{O}$ be such that $\delta_{i}(f)=\alpha_{i}f$ for $i=1,\ldots,m$ and let $\tilde{A}=(A|-\overline{\alpha})$ be the extended Saito matrix with respect to $\mathcal{S}$ and $f$.
For an $\mathcal{O}$-module $M$, let $\operatorname{Fitt}_{i}(M)$ denote the $i$-th Fitting ideal of $M$. Consider the Jacobian ideal $\mathcal{J}_{f}=\mathcal{O}(\partial_{1}(f),\ldots,\partial_{n}(f),f)$ and the $\mathcal{O}$-module $\mathcal{J}_{f}/(f)$ generated by the classes of the partial derivatives $[\partial_{1}(f)],\ldots,[\partial_{n}(f)]$. We set $I_{i}:=\operatorname{Fitt}_{n-i}\left(\mathcal{J}_{f}/(f)\right)$ and $D_{i}:=V(I_{i+1})$.
The rows of $A$ are a generating set of the syzygies of $[\partial_{1}(f)],\ldots,[\partial_{n}(f)]$. A free presentation of $\mathcal{J}_{f}/(f)$ as an $\mathcal{O}$-module is then:
$\mathcal{O}^{m}\stackrel{{\scriptstyle\cdot A}}{{\longrightarrow}}\mathcal{O}^{n}\stackrel{{\scriptstyle\underline{e}_{i}\mapsto[\partial_{i}(f)]}}{{\xrightarrow{\hskip 28.45274pt}}}\mathcal{J}_{f}/(f)\to 0.$
Thus, $I_{i}$ is the ideal generated by the minors of order $i$ of $A$ for $1\leq i\leq n$, $I_{i}=\mathcal{O}$ for $i\leq 0$ and $I_{i}=0$ for $i\geq n$. The set $D_{i}$ is then the vanishing locus of all minors of order $(i+1)$, that is, the set of those points at which $A$ has rank less or equal than $i$:
$D_{i}=\{p\in(\mathbb{C}^{n},0)\mid\operatorname{rank}A(p)\leq i\}.$
Let us see some properties of these sets:
Proposition 2.1.
The following properties hold:
  • (1)
    $D_{0}\subset D_{1}\subset\ldots\subset D_{n-1}\subset D_{n}$.
  • (2)
    $D_{n}=(\mathbb{C}^{n},0)$ and $D_{n-1}=D$.
  • (3)
    If $p\in D\setminus D_{k}$ for some $20\leq k\leq n-$ 2, then $(D,p)\cong(D^{\prime},0)\times(\mathbb{C}^{k+1},0)$, where $(D^{\prime},0)$ is a germ of divisor in $(\mathbb{C}^{n-k-1},0)$.
    In particular, if $p\in D\setminus D_{0}$, then $(D,p)\cong(D^{\prime},0)\times(\mathbb{C},0)$. In this case, we say that $D$ is a product with a smooth factor at $p$.
Proof.
As $I_{i+1}\subset I_{i}$ for all $i$, $D_{i-1}=V(I_{i})\subset V(I_{i+1})=D_{i}$, which proves $(1)$. Since $A$ is an $m\times n$ matrix with $m\geq n$, it is clear that $\operatorname{rank}A(p)\leq n$ for all $p$, so $D_{n}=(\mathbb{C}^{n},0)$. In [12, Proposition 2.2], it is shown in a more general setting that $D=\{p\in(\mathbb{C}^{n},0)\mid\dim\operatorname{Der}_{f}(p)\leq n-1\}$, where $\operatorname{Der}_{f}(p)$ stands for the $\mathbb{C}$-vector subspace of $T_{p}(\mathbb{C}^{n},0)$ generated by $\delta_{1}(p),\ldots,\delta_{m}(p)$. It is clear that $\dim\operatorname{Der}_{f}(p)=\operatorname{rank}A(p)$, so $D=D_{n-1}$ and we get $(2)$.
Let $p\in D\setminus D_{k}$, so $\operatorname{rank}A(p)>k$. Let us see by induction on $k$ that, in a neighbourhood of $p$, at least $k+1$ partial derivatives of $f$ belong to the ideal generated by $f$ and the rest of them.
For $k=0$, at least one of the entries $a_{ij}$ (that corresponds to the coefficient of $\partial_{j}$ in $\delta_{i}$) of $A$ cannot vanish at $p$, so it is a unit in $\mathcal{O}_{X,p}$. Thus, we can isolate $\partial_{j}(f)$ in $\delta_{i}(f)\in(f)$ and see that $\partial_{j}(f)$ belongs to the ideal generated by $f$ and $\partial_{k}(f)$, $k\neq j$.
Suppose the result is true for $k\geq 0$ and let $\operatorname{rank}A(p)>k+1$. As before, we can find $i_{0},j_{0}$ such that $a_{i_{0}j_{0}}(p)\neq 0$. Now set $\delta^{\prime}_{k}=\delta_{k}-\dfrac{a_{kj_{0}}}{a_{i_{0}j_{0}}}\delta_{i_{0}}$ for $k\neq i_{0}$ and $\delta^{\prime}_{i_{0}}=\delta_{i_{0}}$. Let $A^{\prime}$ be the Saito matrix with respect to $\{\delta^{\prime}_{1},\ldots,\delta^{\prime}_{m}\}$ (which is still a generating set) and note that $a^{\prime}_{kj_{0}}(p)=0$ for $k\neq i_{0}$. Since $\operatorname{rank}A^{\prime}(p)=\operatorname{rank}A(p)>k+1$ and $a^{\prime}_{i_{0}j_{0}}(p)=a_{i_{0}j_{0}}(p)\neq 0$, the rank of $A^{\prime}(p)$ is exactly $1+\operatorname{rank}B(p)$, where $B$ is the submatrix of $A^{\prime}$ formed by the entries $a^{\prime}_{ij}$ with $i\neq i_{0}$, $j\neq j_{0}$. Thus, $\operatorname{rank}B(p)>k$ and, by induction hypothesis, at least $k+1$ partial derivatives of $f$ belong to the ideal generated by $f$ and the rest of them except $\partial_{j_{0}}(f)$. But, reasoning as in the case $k=0$, $\partial_{j_{0}}(f)$ also belongs to this ideal, so we conclude that at least $k+2$ partial derivatives of $f$ belong to the ideal generated by $f$ and the rest of them.
Now, by [17, Lemma 3.5], there exist a unit $u$ and a local coordinate change such that $g=uf$ (which is another equation of $D$) only depends on $n-k-1$ variables. If we set $D^{\prime}=V(g)\subset\mathbb{C}^{n-k-1}$, we get $(3)$.
Remark 2.2.
  • (a)
    For free divisors, the equality $D_{n-1}=D$ can be deduced from Saito’s criterion.
  • (b)
    Strong Euler-homogeneity is preserved under smooth factors. That is, if $D$ is a product with a smooth factor at $p$, so that $(D,p)\cong(D^{\prime},0^{\prime})\times(\mathbb{C},0)$, then $D$ is strongly Euler-homogeneous at $p$ if and only if so is $D^{\prime}$ at 00 [9, Lemma 3.2].
Now, call $\tilde{I}_{i}:=\operatorname{Fitt}_{n+1-i}\left(\mathcal{J}_{f}\right)$ and $\tilde{D}_{i}:=V(\tilde{I}_{i+1})$. The rows of $\tilde{A}$ generate the syzygies of $\partial_{1}(f),\ldots,\partial_{n}(f),f$, so a free presentation of $\mathcal{J}_{f}$ as an $\mathcal{O}$-module is:
$\mathcal{O}^{m}\stackrel{{\scriptstyle\cdot\tilde{A}}}{{\longrightarrow}}\mathcal{O}^{n+1}\stackrel{{\scriptstyle\begin{subarray}{c}\underline{e}_{i}\mapsto\partial_{i}(f),\hskip 2.84544pti=1,\ldots,n\\ \underline{e}_{n+1}\mapsto f\end{subarray}}}{{\xrightarrow{\hskip 56.9055pt}}}\mathcal{J}_{f}\to 0.$
Therefore, $\tilde{I}_{i}$ is the ideal generated by the minors of order $i$ of $\tilde{A}$ for $11\leq i\leq n+$ 1, $\tilde{I}_{i}=\mathcal{O}$ for $i\leq 0$ and $\tilde{I}_{i}=0$ for $i\geq n+1$. As before, $\tilde{D}_{i}$ is the set of those points at which $\tilde{A}$ has rank less or equal than $i$:
$\tilde{D}_{i}=\{p\in(\mathbb{C}^{n},0)\mid\operatorname{rank}\tilde{A}(p)\leq i\}.$
Remark 2.3.
  • (a)
    Note that, if $g$ is another equation of $D$, then $\mathcal{J}_{g}=\mathcal{J}_{f}$ and $\mathcal{J}_{f}/(f)=\mathcal{J}_{g}/(g)$. Also, Fitting ideals do not depend on the choice of presentation. Thus, $D_{i}$ (resp. $\tilde{D}_{i}$) and $I_{i}$ (resp. $\tilde{I}_{i}$), which in principle might depend on the equation of $D$ or on the Saito matrix with respect to the chosen basis, only depend on $D$ itself.
  • (b)
    Note that the ideal generated by the $\alpha_{i}$ can be written as the colon ideal $(\partial_{1}(f),\ldots,\partial_{n}(f)):f$ (from now on, we will just write $\partial f:f$ for short). Whereas $I_{i}$ and $\tilde{I}_{i}$ do not depend on the choice of $f$, the ideal $\partial f:f$ does depend on such a choice. In fact, if $g=uf$ for some unit $u$, then
    $\partial g:g=(\alpha_{1}+\delta_{1}(u)u^{-1},\ldots,\alpha_{m}+\delta_{m}(u)u^{-1}).$
The sets $\tilde{D}_{i}$ and ideals $\tilde{I}_{i}$ are related with the previous ones in the following way:
Proposition 2.4.
The following properties hold:
  • (1)
    $\tilde{I}_{i+1}\subset I_{i+1}+(\alpha_{1},\ldots,\alpha_{m})I_{i}\subset I_{i}\subset\tilde{I}_{i}$   for all $i=1,\ldots,n-1$.
  • (2)
    $\tilde{D}_{i}\subset D_{i}\subset\tilde{D}_{i+1}$   for all $i=0,\ldots,n-2$.
  • (3)
    $\tilde{D}_{n-1}=D_{n-2}=\operatorname{Sing}D$.
Proof.
The minors of order $i+1$ of $\tilde{A}$ are the ones of $A$ together with those involving the last column $-\overline{\alpha}$, which are (expanding them by this column) linear combinations of minors of order $i$ of $A$ with coefficients in the $\alpha_{i}$. This justifies the first inclusion of $(1)$. The second inclusion is due to the fact that $I_{i+1}\subset I_{i}$ for all $i$. The last inclusion holds because, $A$ being a submatrix of $\tilde{A}$, the minors of order $i$ of $A$ are also minors of order $i$ of $\tilde{A}$. Since the map $V$ is inclusion-reversing, $(2)$ is immediate from $(1)$.
Let us prove $(3)$. Since $\dim(\operatorname{Sing}D)\leq n-2$ (or $\operatorname{ht}\mathcal{J}_{f}\geq 2$), by [7, Theorem 4.1], we have that $\sqrt{\mathcal{J}_{f}}=\sqrt{\operatorname{Fitt}_{1}(\mathcal{J}_{f})}=\sqrt{\tilde{I}_{n}}$. Thus,
$\tilde{D}_{n-1}=V(\tilde{I}_{n})=V\left(\sqrt{\tilde{I}_{n}}\right)=V\left(\sqrt{\mathcal{J}_{f}}\right)=V(\mathcal{J}_{f})=\operatorname{Sing}D.$
By $(2)$, $D_{n-2}\subset\tilde{D}_{n-1}=\operatorname{Sing}D$. Now, suppose that there exists $p\in\operatorname{Sing}D\setminus D_{n-2}$. Then, by $(3)$ in Proposition 2.1, $(D,p)\cong(\mathbb{C}^{n-1},0)$. But this is a contradiction because $(\mathbb{C}^{n-1},0)$ is non-singular, so $D_{n-2}=\operatorname{Sing}D$ and we have $(3)$.
Remark 2.5.
Since $D_{n-1}=D$ and $\tilde{D}_{n-1}=\operatorname{Sing}D$, we have that $D_{i}$ and $\tilde{D}_{i}$ are closed analytic subsets of $D$ for $10\leq i\leq n-$ 1.
Now, we are going to give algebraic and geometric characterizations of strong Euler-homogeneity for points in a neighbourhood of 00.
As $\tilde{D}_{n-1}=D_{n-2}$, we may wonder if it is always true that $\tilde{D}_{i}=D_{i-1}$ for all $i=0,\ldots,n-1$ (where $D_{-1}=V(I_{0})=V(\mathcal{O})=\varnothing$). We are going to see that this property is, in fact, equivalent to local strong Euler-homogeneity on a neighbourhood of the origin.
This criterion is based in [8, Lemma 7.5], where the authors give a characterization of local strong Euler-homogeneity for linear free divisors, for which there is always a strong Euler vector field at the origin. We have adapted this result to the case in which we do not know a priori whether there exists an Euler vector field.
First, we give a characterization for strong Euler-homogeneity at a point $p$ in terms of the rank of the matrices $A(p)$ and $\tilde{A}(p)$:
Proposition 2.6.
$D$ is strongly Euler-homogeneous at the point $p\in D$ if and only if $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)+1$.
Proof.
Let us note that, as $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ is a generating set of $\operatorname{Der}_{f}$, in a neighbourhood of the origin, $D$ is strongly Euler-homogeneous at $p$ if and only if there exists $a_{1},\ldots,a_{m}\in\mathcal{O}$ such that $\chi=\sum_{i=1}^{m}a_{i}\delta_{i}$ vanishes at $p$ and $\chi(f)=\gamma f$ with $\gamma(p)\neq 0$ (that is, $\sum_{i=1}^{m}a_{i}(p)\alpha_{i}(p)\neq 0$). Moreover, the $\mathbb{C}$-vector space $\tilde{N}=\{\bar{x}\in\mathbb{C}^{m}\mid\tilde{A}(p)^{t}\bar{x}=\bar{0}\}$ is always contained in the $\mathbb{C}$-vector space $N=\{\bar{x}\in\mathbb{C}^{m}\mid A(p)^{t}\bar{x}=\bar{0}\}$, with equality if and only if $\dim\tilde{N}=m-\operatorname{rank}\tilde{A}(p)=m-\operatorname{rank}A(p)=\dim N$, that is, if and only if $\operatorname{rank}A(p)=\operatorname{rank}\tilde{A}(p)$.
Observe that we always have $\operatorname{rank}A(p)\leq\operatorname{rank}\tilde{A}(p)\leq\operatorname{rank}A(p)+1$. Let us suppose that $\operatorname{rank}\tilde{A}(p)\neq\operatorname{rank}A(p)+1$, so $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)$ and $\tilde{N}=N$. Now, if a vector field $\chi=\sum_{i=1}^{m}a_{i}\delta_{i}$ verifies $\chi(p)=0$, then $\bar{a}(p)=(a_{1}(p),\ldots,a_{m}(p))^{t}$ belongs to $N=\tilde{N}$, so $\sum_{i=1}^{m}a_{i}(p)\alpha_{i}(p)=0$ and there cannot be a strong Euler vector field at $p$.
Reciprocally, if $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)+1$, then $\tilde{N}$ is strictly contained in $N$. Therefore, there exists $\bar{a}=(a_{1},\ldots,a_{m})^{t}\in\mathbb{C}^{m}$ such that $\sum_{i=1}^{m}a_{i}\delta_{i}(p)=0$ and $\sum_{i=1}^{m}a_{i}\alpha_{i}(p)\neq 0$. Taking $\chi=\sum_{i=1}^{m}a_{i}\delta_{i}$ we get the desired conclusion.
As a consequence of this result, we get a criterion for strong Euler-homogeneity in terms of the analytic sets $D_{i}$ and $\tilde{D}_{i}$:
Theorem 2.7 (Geometric criterion for strong Euler-homogeneity).
Let $D$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$. Then:
  • (1)
    $D$ is always strongly Euler-homogeneous on $D_{0}\setminus\tilde{D}_{0}$ and non-strongly Euler-homogeneous at any point of $\tilde{D}_{0}$. Thus, $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\tilde{D}_{0}=\varnothing$.
  • (2)
    $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\tilde{D}_{i}=D_{i-1}\text{ for }i=1,\ldots,n-2$.
  • (3)
    $D$ is strongly Euler-homogeneous if and only if $\tilde{D}_{0}=\varnothing$ and $\tilde{D}_{i}=D_{i-1}\text{ for }i=1,\ldots,n-2$.
    In particular, for $n=2$, $D$ is strongly Euler-homogeneous if and only if $\tilde{D}_{0}=\varnothing$.
Proof.
For $p\in\tilde{D}_{0}$, $\operatorname{rank}A(p)=\operatorname{rank}\tilde{A}(p)=0$, so $D$ is not strongly Euler-homogeneous at $p$. For $p\in D_{0}\setminus\tilde{D}_{0}$, $A(p)$ vanishes but $\tilde{A}(p)$ does not, so the last column of $\tilde{A}(p)$ is not zero. This means that $\operatorname{rank}A(p)=0$ and $\operatorname{rank}\tilde{A}(p)=1$, so $D$ is strongly Euler-homogeneous at $p$ and we get $(1)$.
Let us prove $(2)$. First, suppose that $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$. As we always have $D_{i-1}\subset\tilde{D}_{i}\subset D_{i}$ for $i=1,\ldots,n-1$ (Proposition 2.4), it is enough to prove $\tilde{D}_{i}\subset D_{i-1}$. Let $p\in\tilde{D}_{i}$ with $21\leq i\leq n-$ 2 (in particular, $p\in\tilde{D}_{n-2}\subset D$). If $p\in D_{0}$, then $p\in D_{i-1}$ (Proposition 2.1) and we are done. Otherwise, $p\in D\setminus D_{0}$, so $D$ is strongly Euler-homogeneous at $p$ and $\operatorname{rank}A(p)+1=\operatorname{rank}\tilde{A}(p)\leq i$, meaning $\operatorname{rank}A(p)\leq i-1$ and $p\in D_{i-1}$. Reciprocally, suppose $D$ is not strongly Euler-homogeneous at $p\in D\setminus D_{0}$ and let $i_{0}=\operatorname{rank}A(p)=\operatorname{rank}\tilde{A}(p)$. Note that $11\leq i_{0}\leq n-$ 1 as $p\in D\setminus D_{0}=D_{n-1}\setminus D_{0}$. We have $\operatorname{rank}\tilde{A}(p)\leq i_{0}$ and $\operatorname{rank}A(p)>i_{0}-1$, which implies $p\in\tilde{D}_{i_{0}}\setminus D_{i_{0}-1}$. But we have already proved that $\tilde{D}_{n-1}=D_{n-2}$, so $i_{0}\neq n-1$. Thus, we get $\tilde{D}_{i_{0}}\neq D_{i_{0}-1}$ for some $21\leq i_{0}\leq n-$ 2.
Finally, $(3)$ is straightforward from $(1)$ and $(2)$.
Remark 2.8.
One might ask why we should care about strong Euler-homogeneity only on $D_{0}$ or $D\setminus D_{0}$. Here is one reason:
Suppose that we want to prove that a divisor satisfying a certain property $P$ is always strongly Euler-homogeneous. Assume we know how to prove it for plane curves and that the property $P$ is preserved under smooth factors (as it happens with strong Euler-homogeneity). Then, we can apply an inductive argument: suppose the result is true for ambient dimension $k-1$ and consider a divisor $D$ in ambient dimension $k$ satisfying $P$. As $D$ is a product with a smooth factor at every point outside $D_{0}$, by induction hypothesis, we deduce that $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$. Thus, we will only have to show that $P$ implies strong Euler-homogeneity on $D_{0}$ (or $\tilde{D}_{0}=\varnothing$ by Theorem 2.7). And any equivalent or necessary condition for strong Euler-homogeneity outside $D_{0}$ could be useful to this purpose.
This strategy will be used in the proof of Theorem 3.11 and in [6] to show the main results.
Hilbert’s Nullstellensatz gives us an immediate algebraic translation of the developed criterion:
Corollary 2.9 (Algebraic criterion for strong Euler-homogeneity).
Let $D$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$. Then:
  • (1)
    $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\tilde{I}_{1}=\mathcal{O}$.
  • (2)
    $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,\ldots,n-2$.
  • (3)
    $D$ is strongly Euler-homogeneous if and only if $\tilde{I}_{1}=\mathcal{O}$ and $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,\ldots,n-2$.
Remark 2.10.
We have formulated this criterion in terms of ideals of minors of $A$ and $\tilde{A}$, which is more practical for our purpose. In terms of Fitting ideals, this criterion reads as:
  • $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\sqrt{\operatorname{Fitt}_{n}(\mathcal{J}_{f})}=\sqrt{\operatorname{Fitt}_{n}\left(\mathcal{J}_{f}/(f)\right)}\hskip 5.69046pt(=\mathcal{O})$.
  • $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\sqrt{\operatorname{Fitt}_{i}(\mathcal{J}_{f})}=\sqrt{\operatorname{Fitt}_{i}\left(\mathcal{J}_{f}/(f)\right)}$ for all $i=2,\ldots,n-1$.
  • $D$ is strongly Euler-homogeneous if and only if $\sqrt{\operatorname{Fitt}_{i}(\mathcal{J}_{f})}=\sqrt{\operatorname{Fitt}_{i}\left(\mathcal{J}_{f}/(f)\right)}$ for all $i=2,\ldots,n$.
We present now two examples in which we apply the developed criterion to determine if they are strongly Euler-homogeneous or not. Both of them are strongly Euler-homogeneous at the origin but the first one is also strongly Euler-homogeneous outside the origin and the second one is not. Calculations have been made with Macaulay2 [11]:
Example 2.11.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z+yt)$. A minimal generating set of $\operatorname{Der}_{f}$ is:
$\begin{array}[]{l}\delta_{1}=x\partial_{x}+y\partial_{y}+z\partial_{z},\\ \delta_{2}=(z+yt)\partial_{t},\\ \delta_{3}=y(x+z)\partial_{y}-z(x+z)\partial_{z}-2t(x+z)\partial_{t}\\ \delta_{4}=y^{2}(y+z)\partial_{y}+y(x^{2}-z^{2})\partial_{z}+(-x^{2}+yz+2z^{2})\partial_{t},\\ \delta_{5}=yz(y+z)\partial_{y}+z(x^{2}-z^{2})\partial_{z}+(x^{2}t-2z^{2}t+z^{2})\partial_{t}.\\ \end{array}$
As $\delta_{1}(f)=6f$, $\delta_{2}(f)=yf$, $\delta_{3}(f)=-zf$, $\delta_{4}(f)=(xy+3y^{2}+yz)f$ and $\delta_{5}(f)=(x^{2}+xz+3yz-z^{2})f$, the extended Saito matrix is:
$\tilde{A}=\left(\begin{array}[]{cccc|c}x&y&z&0&-6\\ 0&0&0&z+yt&-y\\ 0&y(x+z)&-z(x+z)&-2t(x+z)&z\\ 0&y^{2}(y+z)&y(x^{2}-z^{2})&-x^{2}+yz+2z^{2}&-xy-3y^{2}-yz\\ 0&yz(y+z)&z(x^{2}-z^{2})&x^{2}t-2z^{2}t+z^{2}&-x^{2}-xz-3yz+z^{2}\\ \end{array}\right).$
By Corollary 2.9, $D$ is strongly Euler-homogeneous if and only if $\tilde{I}_{1}=\mathcal{O}$ (this holds because the last column of $\tilde{A}$ contains a unit) and $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,2$. We have:
$\begin{array}[]{l}\sqrt{\tilde{I}_{2}}=\sqrt{I_{1}}=(x,y,z),\\[5.69046pt] \sqrt{\tilde{I}_{3}}=\sqrt{I_{2}}=(x,z,yt).\\[5.69046pt] \end{array}$
We conclude that $D$ is strongly Euler-homogeneous.
Example 2.12.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z^{2}+yt)$ (note that the only difference with the previous example is that the $z$ in the last term of the product is squared). A minimal generating set of $\operatorname{Der}_{f}$ is:
$\begin{array}[]{l}\delta_{1}=x\partial_{x}+y\partial_{y}+z\partial_{z}+t\partial_{t},\\ \delta_{2}=(z^{2}+yt)\partial_{t},\\ \delta_{3}=y(x+z)\partial_{y}-z(x+z)\partial_{z}-3t(x+z)\partial_{t}\\ \delta_{4}=y^{2}(y+z)\partial_{y}+y(x^{2}-z^{2})\partial_{z}-(2x^{2}z+y^{2}t+3yzt)\partial_{t},\\ \delta_{5}=yz(y+z)\partial_{y}+z(x^{2}-z^{2})\partial_{z}+(x^{2}t-2z^{2}t+z^{2})\partial_{t}.\\ \end{array}$
We have $\delta_{1}(f)=7f$, $\delta_{2}(f)=yf$, $\delta_{3}(f)=-(x+2z)f$, $\delta_{4}(f)=(xy+2y^{2}-2yz)f$ and $\delta_{5}(f)=(2x^{2}+xz+2yz-2z^{2}+3yt)f$. The extended Saito matrix is then:
$\tilde{A}=\left(\begin{array}[]{cccc|c}x&y&z&t&-7\\ 0&0&0&z^{2}+yt&-y\\ 0&y(x+z)&-z(x+z)&-3t(x+z)&x+2z\\ 0&y^{2}(y+z)&y(x^{2}-z^{2})&-(2x^{2}z+y^{2}t+3yzt)&-xy-2y^{2}+2yz\\ 0&yz(y+z)&z(x^{2}-z^{2})&2x^{2}t-yzt+3yt^{2}&-2x^{2}-xz-2yz+2z^{2}-3yt\\ \end{array}\right)$.
As before, $D$ is strongly Euler-homogeneous if and only if $\tilde{I}_{1}=\mathcal{O}$ (which holds because of the $-7$ in the last column) and $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,2$. But we have:
$\begin{array}[]{l}\sqrt{\tilde{I}_{2}}=(x,y,z)\subsetneq\sqrt{I_{1}}=(x,y,z,t),\\[5.69046pt] \sqrt{\tilde{I}_{3}}=\sqrt{I_{2}}=(t(y+z),t(x+z),z(z-t),z(y+t),z(x+t),xy-zt).\\[5.69046pt] \end{array}$
This means that $D_{0}\subsetneq\tilde{D}_{1}$ and $D$ is not strongly Euler-homogeneous at any point of $\tilde{D}_{1}\setminus D_{0}=\{(0,0,0,t)\mid t\neq 0\}$, where both matrices $A$ and $\tilde{A}$ have rank $1$.
The following lemma states that the vanishing locus of $\partial f:f$ is always inside the singular locus. This will give us a necessary condition for a divisor to be strongly Euler-homogeneous outside $D_{0}$.
Lemma 2.13.
$V(\partial f:f)\subset\operatorname{Sing}D$.
Proof.
Recall that, with the previous notation, $\partial f:f=(\alpha_{1},\ldots,\alpha_{m})$. As $f$ is integral over the ideal generated by its partial derivatives [18, Satz 5.2], there exists a logarithmic derivation $\eta$ and an integer $r\geq 1$ such that $\eta(f)=f^{r}$. Let us write $\eta=\sum_{i=1}^{m}c_{i}\delta_{i}$, so that we get $f^{r-1}=\sum_{i=1}^{m}c_{i}\alpha_{i}$ when we apply $\eta$ to $f$. If $r=1$, then $V(\alpha_{1},\ldots,\alpha_{m})=\varnothing$ and there is nothing to prove. Otherwise, it is clear that $V(\alpha_{1},\ldots,\alpha_{m})\subset D$. As $D=D_{n-1}$, if $p\in V(\alpha_{1},\ldots,\alpha_{m})$, then $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)\leq n-1$, so $p\in\tilde{D}_{n-1}=\operatorname{Sing}D$.
Corollary 2.14 (Necessary condition for strong Euler-homogeneity outside $D_{0}$).
If $D$ is a germ of divisor in $(\mathbb{C}^{n},0)$ that is strongly Euler-homogeneous on $D\setminus D_{0}$, then $\tilde{D}_{0}=V(\partial f:f)$ and $\sqrt{\tilde{I}_{1}}=\sqrt{\partial f:f}$.
Proof.
It is clear that $\tilde{D}_{0}\subset V(\alpha_{1},\ldots,\alpha_{m})$. For the converse, let $p\in V(\alpha_{1},\ldots,\alpha_{m})$. As the last column of $\tilde{A}$ vanishes at $p$, we have that $\tilde{A}(p)$ and $A(p)$ have the same rank. By Lemma 2.13, $p\in\operatorname{Sing}D=D_{n-2}$, so $\operatorname{rank}A(p)\leq n-2$. Let $i_{0}=\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)\leq n-2$. Then, $i_{0}$ cannot be greater than zero, as then $p\in\tilde{D}_{i_{0}}\setminus D_{i_{0}-1}$ and $D$ would not be strongly Euler-homogeneous outside $D_{0}$ (Theorem 2.7). Therefore, $\operatorname{rank}\tilde{A}(p)=0$ and $p\in\tilde{D}_{0}$. The equality $\sqrt{\tilde{I}_{1}}=\sqrt{\partial f:f}$ follows from the first one by Hilbert’s Nullstellensatz.
This result tells us that, when $D$ is strongly Euler-homogeneous outside $D_{0}$, the radical of $\partial f:f$ does not depend on the choice of $f$. The following example shows that, in order for this to be true, the hypothesis of being strongly Euler-homogeneous on $D\setminus D_{0}$ is necessary:
Example 2.15.
Let $D=V(f)$ in $(\mathbb{C}^{3},0)$ with $f=xyz(x^{3}+xyz+y^{3}z^{3})$. This divisor is not strongly Euler-homogeneous outside $D_{0}$, as $\sqrt{I_{1}}=(x,y,z)\neq(x,yz)=\sqrt{\tilde{I_{2}}}$. And, if we set $g=(1+y)f$, we have $\sqrt{\partial g:g}=(x,y)\neq(5x^{2}+2yz,5y^{2}z^{2}+2x)=\sqrt{\partial f:f}$.
3 Saito-holonomic type properties
Let us begin this section by recalling the different notions of Koszul-freeness defined in [15, Definitions 1.10 and 1.13] and [10, Remark 7.3]. We consider the graded ring $\operatorname{gr}\mathcal{D}_{X,p}$ associated with the filtration induced by the order. For a differential operator $P\in\mathcal{D}_{X,p}$, we will denote by $\sigma(P)$ the principal symbol of $P$ in $\operatorname{gr}\mathcal{D}_{X,p}$. It is well-known that $\operatorname{gr}\mathcal{D}_{X,p}$ can be identified with the commutative ring $\mathcal{O}_{X,p}[\underline{\xi}]$, where $\xi_{i}=\sigma(\partial_{i})$.
Definition 3.1.
Let $D\subset X$ be a free divisor. We say $D$ is:
  • weakly Koszul-free at $p\in D$ if, for some (or any) basis $\{\delta_{1},\ldots,\delta_{n}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$, the sequence
    $\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{n})-\alpha_{n}s,\text{ with }\alpha_{i}\in\mathcal{O}_{X,p}\text{ being such that }\delta_{i}(f)=\alpha_{i}f,$
    is regular in $\operatorname{gr}\mathcal{D}_{X,p}[s]$.
  • Koszul-free at $p\in D$ if, for some (or any) basis $\{\delta_{1},\ldots,\delta_{n}\}$ of $\operatorname{Der}_{X,p}(-\log D)$, the sequence
    $\sigma(\delta_{1}),\ldots,\sigma(\delta_{n}),$
    is regular in $\operatorname{gr}\mathcal{D}_{X,p}$.
  • strongly Koszul-free at $p\in D$ if, for some (or any) basis $\{\delta_{1},\ldots,\delta_{n}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$, the sequence
    $f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{n})-\alpha_{n}s,\text{ with }\alpha_{i}\in\mathcal{O}_{X,p}\text{ being such that }\delta_{i}(f)=\alpha_{i}f,$
    is regular in $\operatorname{gr}\mathcal{D}_{X,p}[s]$.
We say $D$ is weakly Koszul-free, Koszul-free or strongly Koszul-free if so is it at each $p\in D$.
Our purpose is to define analogous properties in the general case in which our divisor is not necessarily free. In order to do this, we need first to make some algebraic considerations:
In a Cohen-Macaulay local ring, a sequence of $r$ elements is regular if and only if the height of the ideal they generate is $r$ [14, Theorem 17.4]. The ring $\mathcal{O}_{X,p}[\underline{\xi}]=\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ (resp. $\mathcal{O}_{X,p}[\underline{\xi},s]$) is not local but we will see that, for homogeneous elements with respect to $\underline{\xi}$ (resp. with respect to $\underline{\xi},s$), this is still true. In order to do so, we will study the extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}\{\underline{x},\underline{\xi}\}$
If a ring map $A\to B$ is faithfully flat, then a sequence of elements is regular in $A$ if and only if the corresponding sequence is regular in $B$. Moreover, if $A$ and $B$ are Noetherian, then the height of an ideal $I\subset A$ (denoted by $\operatorname{ht}I$) coincides with the height of its extension $B\cdot I$ in $B$ [13, (4.C), (5.D) and (13.B.3)]. The extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}\{\underline{x},\underline{\xi}\}$ is not faithfully flat but, for homogeneous elements and ideals, these statements still hold. For the sake of completeness, we include the proof here:
Proposition 3.2.
Let $a_{1},\ldots,a_{r}\in\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ be homogeneous elements with respect to $\underline{\xi}$ and consider the ring extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}\{\underline{x},\underline{\xi}\}$. Then:
  • (1)
    $a_{1},\ldots,a_{r}$ is a regular sequence in $\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ if and only if so is it in $\mathbb{C}\{\underline{x},\underline{\xi}\}$.
  • (2)
    If $I=(a_{1},\ldots,a_{r})$, then $\operatorname{ht}(I)=\operatorname{ht}(\mathbb{C}\{\underline{x},\underline{\xi}\}\cdot I)$.
Consequently, $a_{1},\ldots,a_{r}$ is a regular sequence in $\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ if and only if $\operatorname{ht}(a_{1},\ldots,a_{r})=r$.
Proof.
Let $A=\mathbb{C}\{\underline{x}\}[\underline{\xi}]$. As the extension $\mathbb{C}\{\underline{x},\underline{\xi}\}\subset\mathbb{C}[[\underline{x},\underline{\xi}]]$ is faithfully flat, we just need to prove the statements for the extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}[[\underline{x},\underline{\xi}]]$. Let $M=(\underline{x},\underline{\xi})$ be the maximal homogeneous ideal of $A$, so that the $M$-adic completion of $A$ is $\hat{A}=\mathbb{C}[[\underline{x},\underline{\xi}]]$, and let $S=1+M$. By [13, (24.A) and (24.B)], the canonical map $A\to\hat{A}$ factors through $S^{-1}A$ and the extension $S^{-1}A\to\hat{A}$ is faithfully flat. Thus, it remains to show the results for the extension $A\to S^{-1}A$:
It is easy to see that, if $J$ is a homogeneous ideal, $a\in A$, $s\in S$ and $as\in J$, then $a\in J$. From this, we obtain that $(S^{-1}J)^{c}=J$, where the superindex $c$ means the contraction of the ideal. This allows us to deduce that $a_{1},\ldots,a_{r}$ are a regular sequence in $\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ if and only if so are $a_{1}/1,\ldots,a_{r}/1$ in $S^{-1}A$, which gives $(1)$.
Let us prove $(2)$. Consider a homogeneous (with respect to $\underline{\xi}$) prime ideal $\mathfrak{p}$. Note that $S$ cannot meet $\mathfrak{p}$ because the 00-th degree part of an element in the intersection would be a unit belonging to $\mathfrak{p}$. Consequently, we have $\operatorname{ht}(\mathfrak{p})=\operatorname{ht}(S^{-1}\mathfrak{p})$ [1, VIII.9, Corollaire after Proposition 7]. As minimal primes over homogeneous ideals are homogeneous [1, VIII.64, Lemme 1.c)], we deduce that $\operatorname{ht}(I)=\operatorname{ht}(S^{-1}I)$, which finishes the proof.
The last statement follows from the previous ones and the fact that $\mathbb{C}\{\underline{x},\underline{\xi}\}$ is a Cohen-Macaulay local ring.
Now, we are ready to generalize Koszul type properties:
Definition 3.3.
Let $D\subset X$ be a divisor. We say:
  • $D$ is weakly Saito-holonomic at $p\in D$ if, for some (or any) generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$,
    $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n\text{ in }\operatorname{gr}\mathcal{D}_{X,p}[s],$
    where $\alpha_{i}\in\mathcal{O}_{X,p}$ is such that $\delta_{i}(f)=\alpha_{i}f.$
  • $D$ is Saito-holonomic at $p\in D$ if, for some (or any) generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$,
    $\operatorname{ht}(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))=n\text{ in }\operatorname{gr}\mathcal{D}_{X,p}.$
  • $D$ is strongly Saito-holonomic at $p\in D$ if, for some (or any) generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$,
    $\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1\text{ in }\operatorname{gr}\mathcal{D}_{X,p}[s],$
    where $\alpha_{i}\in\mathcal{O}_{X,p}$ is such that $\delta_{i}(f)=\alpha_{i}f.$
We say $D$ is weakly Saito-holonomic, Saito-holonomic or strongly Saito-holonomic if so is it at each $p\in D$.
Remark 3.4.
  • (a)
    For free divisors, these definitions coincide with those of Koszul-freeness given in Definition 3.1 by virtue of Proposition 3.2.
  • (b)
    We will see that the classical definition of holonomicity given by K. Saito in terms of the logarithmic stratification [17, Definition 3.8] is exactly our definition of Saito-holonomicity, which justifies the name we have chosen for this property, in agreement with [20, Definition 2.5].
By coherence, if $D$ satisfies any of the properties in Definition 3.3 at $p$, then so does it in a neighbourhood of $p$. This will allow us to characterize divisors in $(\mathbb{C}^{n},0)$ satisfying them in a neighbourhood of 00 by just looking at the dimensions of the germs $D_{i}$ and $\tilde{D}_{i}$ defined in the previous section. In order to do this, we need first a technical lemma:
Lemma 3.5.
Let $E$ be a non-empty analytic set in $(\mathbb{C}^{n},0)$ given by $k$ equations with $0\leq k\leq n$ and let $B(x)$ be an $m\times r$ matrix with entries in $\mathcal{O}$. Let $W=\{(x,y)\in(E,0)\times(\mathbb{C}^{r},0)\mid B(x)y=0\}$ and let $B_{i}=\{p\in(E,0)\mid\operatorname{rank}B(p)\leq i\}$. Then, $\dim W\leq r-k$ if and only if $B_{i}=\varnothing$ for all $i<k$ and $\dim B_{i}\leq i-k$ for all $k\leq i\leq s$, where $s=\min\{m,r\}$. In particular, as $B_{s}=E$, this can only hold when $s\geq n$.
Proof.
Let us consider the canonical projection $\pi:W\to E$. Note that it is surjective because $(p,0)\in W$ for every $p\in E$, so $\pi(\pi^{-1}(B_{i}))=B_{i}$ for all $i$. Let $0\leq i\leq s$ be such that $B_{i}\neq\varnothing$, so that $\dim B_{i}\geq 0$. For every $p\in B_{i}$, the dimension of $\pi^{-1}(p)$ is at least $r-i$ (because the dimension of $\{y\in(\mathbb{C}^{r},0)\mid B(p)y=0\}$ is $r-\operatorname{rank}B(p)$ and $\operatorname{rank}B(p)\leq i$), so $\dim\pi^{-1}(B_{i})\geq\dim B_{i}+r-i$. Now, if $\dim W\leq r-k$, then
$r-k\geq\dim W\geq\dim\pi^{-1}(B_{i})\geq\dim B_{i}+r-i\hskip 5.69046pt(\geq r-i).$
In particular, $i\geq k$ (so $B_{i}=\varnothing$ for $i<k$) and $\dim B_{i}\leq i-k$.
Note that $B_{s}=E$. Since $E$ is given by $k$ equations, $\dim E\geq n-k$ so we get $n-k\leq\dim E=\dim B_{s}\leq s-k$ and $s\geq n$.
To prove the converse, consider an irreducible component $W^{\prime}$ of $W$. We can distinguish two cases:
  • If $W^{\prime}\subset\pi^{-1}(B_{0})$, in particular, $B_{0}\neq\varnothing$, so $k=0$ and $\dim B_{0}=0$ (by hypothesis). As $p\in B_{0}$ implies $\operatorname{rank}B(p)=0$ and $\dim\pi^{-1}(p)=r$, we get $\dim W^{\prime}\leq\dim\pi^{-1}(B_{0})=r=r-k$.
  • If $W^{\prime}\not\subset\pi^{-1}(B_{0})$, as $W^{\prime}\subset W=\pi^{-1}(E)=\pi^{-1}(B_{s})$, there exists $i\leq s$ for which $W^{\prime}\subset\pi^{-1}(B_{i})\setminus\pi^{-1}(B_{i-1})$. In particular, $B_{i}\setminus B_{i-1}$ (which is the set of points $p\in E$ such that $\operatorname{rank}B(p)=i$) is a non-empty open set of $B_{i}$ where the fibre dimension of $\pi:W^{\prime}\to B_{i}$ attains the minimum value $\lambda(\pi)=r-i$. Since $W^{\prime}$ is irreducible, we have that $\dim W^{\prime}=\lambda(\pi)+\dim\pi(W^{\prime})\leq r-i+\dim B_{i}\leq r-k$, where the last inequality holds by hypothesis.
We conclude that every irreducible component of $W^{\prime}$ verifies $\dim W^{\prime}\leq r-k$ and, thus, $\dim W\leq r-k$.
In what follows, we will need to consider the following map of graded algebras:
$\begin{array}[]{cccccc}\varphi:&\mathcal{O}[\xi_{1},\ldots,\xi_{n},s]&\longrightarrow&\operatorname{\mathcal{R}}(\mathcal{J}_{f})\\ &\xi_{i}&\mapsto&\partial_{i}(f)\cdot t\\ &s&\mapsto&f\cdot t\end{array}$
where $\operatorname{\mathcal{R}}(\mathcal{J}_{f})=\bigoplus_{i=0}^{\infty}\mathcal{J}_{f}^{i}t^{i}$ is the Rees algebra of $\mathcal{J}_{f}$. As $\varphi$ is surjective and $\operatorname{\mathcal{R}}(\mathcal{J}_{f})$ is a domain, $\ker\varphi$ is a prime ideal. Since the dimension of $\mathcal{O}[\xi_{1},\ldots,\xi_{n},s]$ is $12n+$ 1 and that of $\operatorname{\mathcal{R}}(\mathcal{J}_{f})$ is $n+1$ (cf. [19, Theorem 5.1.4]), we deduce that $\operatorname{ht}(\ker\varphi)=n$.
The homogeneous part of degree $1$ of $\ker\varphi$ is generated by the set $\{\sigma(\delta_{i})-\alpha_{i}s,i=1,\ldots,m\}$ (in the notation of Definition 3.3) and we denote it by $\ker^{(1)}\varphi$. A divisor $D$ is said to be of linear Jacobian type if $\ker\varphi=\ker^{(1)}\varphi$.
Theorem 3.6.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$. Then:
  • (1)
    $D$ is weakly Saito-holonomic if and only if $\dim\tilde{D}_{i}\leq i$ for all $i=0,\ldots,n-3$.
  • (2)
    $D$ is Saito-holonomic if and only if $\dim D_{i}\leq i$ for all $i=0,\ldots,n-3$.
  • (3)
    $D$ is strongly Saito-holonomic if and only if $\tilde{D}_{0}=\varnothing$ and $\dim\tilde{D}_{i}\leq i-1$ for all $i=1,\ldots,n-2$.
Proof.
Let $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ be a generating set of $\operatorname{Der}_{f}$ and let $A$ be the Saito matrix with respect to $\mathcal{S}$. Let $\alpha_{i}\in\mathcal{O}$ such that $\delta_{i}(f)=\alpha_{i}f$ for all $i=1,\ldots,m$ and let $\tilde{A}$ be the extended Saito matrix. By coherence, it is enough to show these properties hold at 00:
  • (1)
    $D$ is weakly Saito-holonomic at 00 if and only if $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n$ in $\mathcal{O}[\underline{\xi},s]$. As $(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)$ is a homogeneous ideal (with respect to $\xi_{1},\ldots,\xi_{n},s$), by Proposition 3.2, this is equivalent to $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n$ in $\mathbb{C}\{\underline{x},\underline{\xi},s\}$ or even $\dim W=n+1$, where $W=V(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)\subset(\mathbb{C}^{2n+1},0)$. As $(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=\ker^{(1)}\varphi\subset\ker\varphi$ and $\ker\varphi$ is a prime ideal of $\mathcal{O}[\underline{\xi},s]$ of height $n$, the inequality $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)\leq n$ (or $\dim W\geq n+1$) always holds. Thus, $D$ is weakly Saito-holonomic at 00 if and only if $\dim W\leq n+1$. Note that we can write
    $W=\{(x,(\xi,s))\in(\mathbb{C}^{n},0)\times(\mathbb{C}^{n+1},0)\mid\tilde{A}(x)(\xi,s)^{t}=0\}.$
    Now, we just need to apply Lemma 3.5. With the notation of the lemma, $E=\mathbb{C}^{n},k=0$, $B(x)=\tilde{A}(x)$, $r=n+1$, $s=\min\{m,n+1\}\in\{n,n+1\}$ (recall that $m\geq n$) and $B_{i}=\tilde{D}_{i}$. Thus, $\dim W\leq n+1$ if and only if $\dim\tilde{D}_{i}\leq i$ for all $i=0,\ldots,s$. The conditions $\dim\tilde{D}_{n}\leq n$ and $\dim\tilde{D}_{n+1}\leq n+1$ (if needed) always holds because $\tilde{D}_{n}\subset\tilde{D}_{n+1}\subset\mathbb{C}^{n}$. Moreover, since $\tilde{D}_{n-2}\subset\tilde{D}_{n-1}=\operatorname{Sing}D$ and the singular locus of $D$ is of dimension at most $n-2$, we have $\dim\tilde{D}_{n-2}\leq\dim\tilde{D}_{n-1}\leq n-2<n-1$. We conclude that $D$ is weakly Saito-holonomic if and only if $\dim\tilde{D}_{i}\leq i$ for all $i=0,\ldots,n-3$.
  • (2)
    $D$ is Saito-holonomic at 00 if and only if $\operatorname{ht}(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))=n$ in $\mathcal{O}[\underline{\xi}]$, or in $\mathbb{C}\{\underline{x},\underline{\xi}\}$ by Proposition 3.2. Since $(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))\subset(\xi_{1},\ldots,\xi_{n})$, we get $\operatorname{ht}(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))\leq n$. Thus, we just need to consider the other inequality, which is equivalent to $\dim W\leq n$, where $W=V(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))=\{(x,\xi)\in(\mathbb{C}^{n},0)\times(\mathbb{C}^{n},0)\mid A(x)\xi^{t}=0\}$. We proceed as in $(1)$ with $E=\mathbb{C}^{n},k=0,B(x)=A(x)$, $r=n$, $s=\min\{m,r\}=n$ and $B_{i}=D_{i}$ to obtain that $\dim W\leq n$ if and only if $\dim D_{i}\leq i$ for all $0\leq i\leq n$. Since $D_{n}=\mathbb{C}^{n}$, $D_{n-1}=D$ and $D_{n-2}=\operatorname{Sing}D$, the last three conditions are automatically satisfied.
  • (3)
    $D$ is strongly Saito-holonomic at 00 if and only if $\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1$ in $\mathcal{O}[\underline{\xi},s]$ or, reasoning as before, if $\dim W=n$, where $W=V(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=\{(x,(\xi,s))\in D\times(\mathbb{C}^{n+1},0)\mid\tilde{A}(x)(\xi,s)^{t}=0\}$. Note that
    $\dim W\geq\dim V(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)-1\geq n,$
    where the last inequality was argued in $(1)$. Thus, $D$ is strongly Saito-holonomic if and only if $\dim W\leq n$, where $W=V(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=\{(x,(\xi,s))\in D\times(\mathbb{C}^{n+1},0)\mid\tilde{A}(x)(\xi,s)^{t}=0\}$. We apply again Lemma 3.5 with $E=D,k=1,B(x)=\tilde{A}(x)$, $r=n+1$ and $s=\min\{m,n+1\}\in\{n,n+1\}$ to obtain that $\dim W\leq n$ if and only if $B_{0}=\varnothing$ and $\dim B_{i}\leq i-1$ for all $i=1,\ldots,s$. Now, note that $B_{i}=\{p\in D\mid\operatorname{rank}\tilde{A}(p)\leq i\}=\tilde{D}_{i}$ for $i\leq n-1$, since $\tilde{D}_{i}\subset D$ for $i\leq n-1$ (observe that in $\tilde{D}_{i}$ we allow $p$ to be in principle any point of $(\mathbb{C}^{n},0)$, while in $B_{i}$ it must be a point of $D$). Since $B_{n-1}=\tilde{D}_{n-1}=\operatorname{Sing}D$ and $B_{n}\subset B_{n+1}=D$, the condition $\dim B_{i}\leq i-1$ is always satisfied for $i\in\{n-1,n,n+1\}$. We conclude that $D$ is strongly Saito-holonomic if and only if $\tilde{D}_{0}=\varnothing$ and $\dim\tilde{D}_{i}\leq i-1$ for $i=1,\ldots,n-2$.
Remark 3.7.
  • (a)
    As $\tilde{D}_{i}\subset D_{i}\subset\tilde{D}_{i+1}$, it is now clear that strongly Saito-holonomic $\Rightarrow$ Saito-holonomic $\Rightarrow$ weakly Saito-holonomic.
  • (b)
    If $D$ is Saito-holonomic or weakly Saito-holonomic in a punctured neighbourhood of a point $p$, then so is it at $p$. This is because outside the fibre at $p$, the set $W$ (in the notation of the proof of Theorem 3.6) has the required dimension ($n$ or $n+1$, respectively) and the fibre at $p$ (of maximal dimension $n$ or $n+1$, respectively) cannot increase that dimension.
    This argument cannot be applied to strong Saito-holonomicity, since we need $W$ to be $n$-dimensional and the fibre at $p$ could be of dimension $n+1$. In fact, any non-quasihomogeneous germ of plane curve (which is a non-strongly Euler-homogeneous isolated singularity) is a counterexample (see Theorem 3.11).
  • (c)
    The characterization given by K. Saito for holonomic divisors in terms of the dimension of the sets $D_{i}$ ($A_{r}$ in his notation) coincides with the one given in Theorem 3.6 for Saito-holonomic divisors (see [17, Definition 3.12 and Lemma 3.13]). Thus, our definition of holonomicity is equivalent to the one given by K. Saito. This generalizes [8, Theorem 7.4], which states that holonomic free divisors are exactly Koszul-free divisors. The names for weak and strong Saito-holonomicity were chosen by analogy with Koszul type properties in the free case.
Some straightforward consequences of Theorem 3.6 in low dimension are:
Corollary 3.8.
Let $D$ be a germ of divisor in $(\mathbb{C}^{n},0)$.
  • (1)
    If $n=2$, then $D$ is Saito-holonomic [2, Corollary 4.2.2].
  • (2)
    If $n=3$ and $D$ is strongly Euler-homogeneous, then it is weakly Saito-holonomic.
Proof.
For $n=2$, the conditions for Saito-holonomicity in Theorem 3.6 are automatically satisfied. For $n=3$, $D$ is weakly Saito-holonomic if and only if $\dim\tilde{D}_{0}\leq 0$. But strong Euler-homogeneity implies $\tilde{D}_{0}=\varnothing$, so this holds.
As it happens with strong Euler-homogeneity and Koszul-freeness [4, Proposition 1.10], Saito-holonomicity is well-behaved with respect to smooth factors:
Proposition 3.9.
Let $D^{\prime}$ be a germ of divisor in $(\mathbb{C}^{n-1},0)$ and let $D=D^{\prime}\times\mathbb{C}\subset(\mathbb{C}^{n},0)$. Then $D$ is weakly Saito-holonomic, Saito-holonomic or strongly Saito-holonomic if and only if so is $D^{\prime}$.
Proof.
Let $D^{\prime}=V(g)$, where $g\in\mathcal{O}_{n-1}=\mathbb{C}\{x_{2},\ldots,x_{n}\}$ and let $f(x_{1},\ldots,x_{n})=g(x_{2},\dots,x_{n})$, so that $D=V(f)$. If $\mathcal{S}^{\prime}=\{\delta_{1},\ldots,\delta_{m}\}$ is a generating set for $\operatorname{Der}_{g}$, then $\mathcal{S}=\{\partial_{1},\delta_{1},\ldots,\delta_{m}\}$ is a generating set for $\operatorname{Der}_{f}$ (where $\partial_{1}(f)=0$). Thus, if $A^{\prime}$ and $\tilde{A}^{\prime}$ are the Saito matrices for $D^{\prime}$ with respect to $\mathcal{S}^{\prime}$ (and $g$), then the Saito matrices for $D$ with respect to $\mathcal{S}$ (and $f$) are:
$A=\left(\begin{array}[]{@{}c|c@{}}1&\begin{matrix}0&\ldots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\end{matrix}&A^{\prime}\end{array}\right),\hskip 14.22636pt\tilde{A}=\left(\begin{array}[]{@{}c|c@{}}1&\begin{matrix}0&\ldots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\end{matrix}&\tilde{A}^{\prime}\end{array}\right).$
It is clear that $\operatorname{rank}A(p)=1+\operatorname{rank}A^{\prime}(p)$ and $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}\tilde{A}^{\prime}(p)+1$. Consequently, we have the relations $D_{i}=D^{\prime}_{i-1}\times\mathbb{C}$ and $\tilde{D}_{i}=\tilde{D}^{\prime}_{i-1}\times\mathbb{C}$ for $i\geq 1$. Observe that $D_{i}=\varnothing$ if and only if $D^{\prime}_{i-1}=\varnothing$ and similarly for $\tilde{D}_{i}$ and $\tilde{D}^{\prime}_{i-1}$. Note also that $D_{0}=\tilde{D}_{0}=\varnothing$. Thus, if they are not empty, $\dim D_{i}\leq k$ if and only if $\dim D^{\prime}_{i-1}\leq k-1$ and $\dim\tilde{D}_{i}\leq k$ if and only if $\dim\tilde{D}^{\prime}_{i-1}\leq k-1$. This proves the result for weakly Saito-holonomic and Saito-holonomic divisors if we take $k=i$ by virtue of Theorem 3.6. Taking $k=i-1$ and noticing that if $\dim\tilde{D}_{1}=\dim(\tilde{D}_{0}^{\prime}\times\mathbb{C})\leq 0$, then $\tilde{D}_{0}^{\prime}=\varnothing$, we prove the result for strong Saito-holonomicity.
A direct algebraic translation of Theorem 3.6 is the following:
Corollary 3.10.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$. Then:
  • (1)
    $D$ is weakly Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\mathcal{J}_{f})\geq i$ for all $i=3,\ldots,n$.
  • (2)
    $D$ is Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\mathcal{J}_{f}/(f))\geq i+1$ for all $i=2,\ldots,n-1$.
  • (3)
    $D$ is strongly Saito-holonomic if and only if $\hskip 2.84544pt(\tilde{I}_{1}=)\operatorname{Fitt}_{n}(\mathcal{J}_{f})=\mathcal{O}$ and $\operatorname{ht}\operatorname{Fitt}_{i}(\mathcal{J}_{f})\geq i+1$ for all $i=2,\ldots,n-1$.
Proof.
To prove $(1)$ and $(3)$ we use that $\tilde{D}_{i}=V(\tilde{I}_{i+1})=V(\operatorname{Fitt}_{n-i}(\mathcal{J}_{f}))$, so $\dim\tilde{D}_{i}\leq k$ if and only if $\operatorname{ht}\operatorname{Fitt}_{n-i}(\mathcal{J}_{f})\geq n-k$. We proceed similarly for $(2)$ with $D_{i}=V(I_{i+1})=V(\operatorname{Fitt}_{n-i-1}(\mathcal{J}_{f}/(f)))$.
In the free case, we know that, if a divisor is Koszul-free, then it is strongly Koszul-free if and only if it is strongly Euler-homogeneous [15, Theorem 4.7]. We are now able to generalize this result to the non-free case:
Theorem 3.11.
Let $D\subset X$ be a Saito-holonomic divisor. Then, it is strongly Saito-holonomic if and only if it is strongly Euler-homogeneous.
Proof.
These properties being local, we may assume $D$ is a germ of divisor in $(\mathbb{C}^{n},0)$. If $D$ is strongly Euler-homogeneous, then $\tilde{D}_{0}=\varnothing$ and $D_{i-1}=\tilde{D}_{i}$ for all $i=1,\ldots,n-1$. Since $D$ is also Saito-holonomic, $\dim D_{i}\leq i$ for all $i=0,\ldots,n-3$, so $\dim\tilde{D}_{i}=\dim D_{i-1}\leq i-1$ for all $i=1,\ldots,n-2$ and $D$ is strongly Saito-holonomic.
To prove the converse, we will proceed by induction on $n$. For $n=2$, by Theorem 3.6, a divisor is strongly Saito-holonomic if and only if $\tilde{D}_{0}=\varnothing$. By Theorem 2.7, this happens if and only if it is strongly Euler-homogeneous. Suppose the result is true for $n-1$ and consider a strongly Saito-holonomic divisor $D$ in $(\mathbb{C}^{n},0)$. Since strong Saito-holonomicity is well behaved with respect to smooth factors (Proposition 3.9), by Remark 2.8, we only have to show that $\tilde{D}_{0}=\varnothing$. But this is guaranteed for a strongly Saito-holonomic divisor by Theorem 3.6.
Remark 3.12.
There are plenty of results in [20] in which the divisor is assumed to be Saito-holonomic and strongly Euler-homogeneous. By Theorem 3.11, all of them can be reformulated in terms of strong Saito-holonomicity.
In the free case, it is known that being of linear Jacobian type is equivalent to being strongly Koszul-free [15, Proposition 1.11]. In the non-free case, we can only say the following:
Proposition 3.13.
Let $D\subset X$ be a divisor of linear Jacobian type. Then, it is strongly Saito-holonomic.
Proof.
We may suppose $D$ is a germ of divisor in $(\mathbb{C}^{n},0)$. As $D$ is of linear Jacobian type, we have that $I:=\ker^{(1)}\varphi=\ker\varphi$. But $\ker\varphi$ is a prime ideal of height $n$ and $f\notin\ker\varphi$, so $\operatorname{ht}(I+(f))=\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1$. Thus, $D$ is strongly Saito-holonomic.
Example 3.14.
The converse is not true. Consider the hyperplane arrangement $D=V(f)$, where $f=xyz(x+t)(y+t)(z+t)(x+y+t)(x+z+t)(y+z+t)$ in $(\mathbb{C}^{4},0)$. It is well-known that all hyperplane arrangements are Saito-holonomic and obviously strongly Euler-homogeneous. Thus, they are strongly Saito-holonomic by Theorem 3.11. However, $D$ is an example of a hyperplane arrangement that is not of differential linear type (that is, the $\mathcal{D}[s]$-annihilator of $f^{s}$ is not generated by operators of order one) [20, Example 5.7], so it cannot be of linear Jacobian type [5, Proposition 1.15].
Nevertheless, if we impose an extra condition on $\ker^{(1)}\varphi$, we have the converse:
Proposition 3.15.
Let $D$ be a germ of strongly Saito-holonomic divisor in $(\mathbb{C}^{n},0)$. If $\ker^{(1)}\varphi$ is unmixed (that is, all its associated primes have the same height), then $D$ is of linear Jacobian type.
Proof.
As $D$ is strongly Saito-holonomic, in particular, it is weakly Saito-holonomic. Thus, $\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1$ and $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n$. That is, $\operatorname{ht}((f)+I)=n+1$ and $\operatorname{ht}I=n$, where $I=\ker^{(1)}\varphi$. A divisor is always of linear Jacobian type at smooth points, so the quotient $\ker\varphi/\ker^{(1)}\varphi$ is supported at the singular locus. In particular, for any $a\in\ker\varphi$, there exists $k\geq 1$ such that $f^{k}a\in I$. Thus, it remains to prove that $f$ is not a zero divisor modulo $I$. If it were, then $f$ would lie in some associated prime $\mathfrak{p}$ of $I$. In particular, $(f)+I\subset\mathfrak{p}$. But then $\operatorname{ht}\mathfrak{p}\geq\operatorname{ht}((f)+I)=n+1$, which is a contradiction with the fact that $I$ is unmixed.
Remark 3.16.
  • (a)
    The unmixedness theorem, which holds in any Cohen-Macaulay ring, states that any ideal of height $h$ that can be generated by $h$ elements is unmixed [13, 16.C]. Thus, Proposition 3.15 applies when $\ker^{(1)}\varphi$ can be generated by $n$ elements. This is what happens in the free case.
    This result suggests the introduction of stronger versions of the Saito-holonomic type properties, in which we also require that the ideal in question is unmixed. Those notions would coincide with the previous ones for free divisors, for which the unmixedness hypothesis is automatic. However, we have not explored the relations between them.
  • (b)
    Proposition 3.15 complements [20, Corollary 3.23] where, instead of unmixedness, tameness is assumed.
Let us revisit Examples 2.11 and 2.12 to study their Saito-holonomic type properties:
Example 3.17.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z+yt)$. We have that
$\begin{array}[]{l}\tilde{D}_{0}=\varnothing,\\[5.69046pt] \tilde{D}_{1}=D_{0}=V(x,y,z),\\[5.69046pt] \tilde{D}_{2}=D_{1}=V(x,z,yt).\\[5.69046pt] \end{array}$
As $\dim D_{0}=1$, $D$ is not Saito-holonomic. Consequently, it is not strongly Saito-holonomic. However, $\dim\tilde{D}_{i}\leq i$ for $i=0,1$, so it is weakly Saito-holonomic.
Recall that $D$ is strongly Euler-homogeneous. This example tells us that Saito-holonomicity is necessary in Theorem 3.11. Weak Saito-holonomicity is not enough to guarantee the equivalence between strong Euler-homogeneity and strong Saito-holonomicity.
Example 3.18.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z^{2}+yt)$. We have that:
$\begin{array}[]{l}\tilde{D}_{0}=\varnothing\\[5.69046pt] \tilde{D}_{1}=V(x,y,z),\hskip 8.5359ptD_{0}=(x,y,z,t),\\[5.69046pt] \tilde{D}_{2}=D_{1}=V(t(y+z),t(x+z),z(z-t),z(y+t),z(x+t),xy-zt).\end{array}$
As $\dim D_{i}=i$ for $i=0,1$, $D$ is Saito-holonomic (and, therefore, also weakly Saito-holonomic). Since $D$ is not strongly Euler-homogeneous, Theorem 3.11 tells us that $D$ is not strongly Saito-holonomic either. And, indeed, we confirm this by noticing that $\dim\tilde{D}_{1}=1>0$.
4 From convergent to formal criteria
Let us denote by $\widehat{\mathcal{O}}=\mathbb{C}[[x_{1},\ldots,x_{n}]]$ the $\mathfrak{m}$-adic completion of $\mathcal{O}$, which is the local ring of formal power series with maximal ideal $\widehat{\mathfrak{m}}$. Let $\widehat{\operatorname{Der}}$ be the $\mathfrak{m}$-adic completion of $\operatorname{Der}$, which coincides with the $\widehat{\mathcal{O}}$-module of $\mathbb{C}$-derivations of $\widehat{\mathcal{O}}$, freely generated by $\partial_{1},\ldots,\partial_{n}$.
Let $f\in\mathcal{O}$ and denote by $\widehat{\operatorname{Der}}_{f}$ the $\mathfrak{m}$-adic completion of $\operatorname{Der}_{f}$, which consists of those formal derivations $\delta\in\widehat{\operatorname{Der}}$ such that $\delta(f)\in\widehat{\mathcal{O}}(f)$. In [9, Theorem 5.4], the authors show that there exists a minimal generating set of the formal logarithmic derivations which takes a particularly nice form in some formal coordinate system, resulting in simple Saito matrices. This is the so-called formal structure theorem.
The objective of this section is to extend the characterizations developed in previous sections, in which the Saito matrices play an important role, to the case of formal power series. This, together with the formal structure theorem, will be useful in [6] to obtain the principal results.
Take a generating set $\mathcal{T}=\{\hat{\delta}_{1},\ldots,\hat{\delta}_{m}\}$ of $\widehat{\operatorname{Der}}_{f}$ and let $B=(\hat{\delta}_{i}(x_{j}))$ be the (formal) Saito matrix with respect to $\mathcal{T}$. Let $\hat{\alpha}_{i}\in\widehat{\mathcal{O}}$ be such that $\hat{\delta}_{i}(f)=\hat{\alpha}_{i}f$ and define similarly $\tilde{B}=\left(B\hskip 2.84544pt\Big{|}-\overline{\hat{\alpha}}\right)$, the (formal) extended Saito matrix with respect to $\mathcal{T}$ and $f$.
Consider the formal Jacobian ideal $\widehat{\mathcal{J}}_{f}=\widehat{\mathcal{O}}(\partial_{1}(f),\ldots,\partial_{n}(f),f)$. Let $J_{i}=\operatorname{Fitt}_{n-i}(\widehat{\mathcal{J}}_{f}/(f))$. As in the convergent case, a free presentation of $\widehat{\mathcal{J}}_{f}/(f)$ as an $\widehat{\mathcal{O}}$-module is given by the matrix $B$, so $J_{i}$ is the ideal of $\widehat{\mathcal{O}}$ generated by the minors of order $i$ of $B$ (for $1\leq i\leq n$). Similarly, if we set $\tilde{J}_{i}=\operatorname{Fitt}_{n+1-i}(\widehat{\mathcal{J}}_{f})$, then $\tilde{J}_{i}$ is the ideal of $\widehat{\mathcal{O}}$ generated by the minors of order $i$ of $\tilde{B}$.
Once more, these ideals do not depend on the choice of basis, coordinates or equation. Note that any convergent generating set $\mathcal{S}$ of $\operatorname{Der}_{f}$ is still an $\widehat{\mathcal{O}}$-generating set of $\widehat{\operatorname{Der}}_{f}$, so we deduce that $J_{i}$ and $\tilde{J}_{i}$ are also the ideals of $\widehat{\mathcal{O}}$ generated by the minors of order $i$ of the Saito matrices with respect to $\mathcal{S}$ and $f$. That is, $J_{i}=I_{i}^{e}$ and $\tilde{J}_{i}=\tilde{I}_{i}^{e}$ (where the superindex $e$ means the extended ideal in $\widehat{\mathcal{O}}$).
In particular, by $(2)$ in Proposition 2.4, we have that $\sqrt{J_{n-1}}=\sqrt{\tilde{J}_{n}}=\sqrt{\widehat{\mathcal{J}}_{f}}$.
Now, we can state the algebraic criterion for strong Euler-homogeneity in terms of $J_{i}$ and $\tilde{J}_{i}$:
Proposition 4.1 (Algebraic formal criterion for strong Euler-homogeneity).
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$. Then:
  • (1)
    $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\tilde{J}_{1}=\widehat{\mathcal{O}}$.
  • (2)
    $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\sqrt{\tilde{J}_{i+1}}=\sqrt{J_{i}}$ for $i=1,\ldots,n-2$.
  • (3)
    $D$ is strongly Euler-homogeneous if and only if $\tilde{J}_{1}=\widehat{\mathcal{O}}$ and $\sqrt{\tilde{J}_{i+1}}=\sqrt{J_{i}}$ for $i=1,\ldots,n-2$.
Proof.
Due to Noetherianity and faithful flatness of $\widehat{\mathcal{O}}$ over $\mathcal{O}$, given two ideals $I,J\subset\mathcal{O}$, we have that $\sqrt{I}=\sqrt{J}$ if and only if $\sqrt{I^{e}}=\sqrt{J^{e}}$ in $\widehat{\mathcal{O}}$. Thus, $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ if and only if $\sqrt{\tilde{J}_{i+1}}=\sqrt{J_{i}}$. The result follows from Corollary 2.9.
Remark 4.2.
As before, this criterion can be read in terms of Fitting ideals. We just need to replace $\mathcal{J}_{f}$ by $\widehat{\mathcal{J}}_{f}$ and $\mathcal{J}_{f}/(f)$ by $\widehat{\mathcal{J}}_{f}/(f)$ in Remark 2.10.
The necessary condition for having strong Euler-homogeneity outside $D_{0}$ can also be translated to the formal setting:
Corollary 4.3.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$ that is strongly Euler-homogeneous outside $D_{0}$. Let $g$ be a “formal equation” of $D$ (i.e. a formal unit multiplied by $f$). Then $\sqrt{\tilde{J}_{1}}=\sqrt{\partial g:g}$.
Proof.
Let $\{\hat{\eta}_{1},\ldots,\hat{\eta}_{m}\}$ be a generating set of $\widehat{\operatorname{Der}}_{f}$ such that $\hat{\eta}_{i}(g)=\hat{\beta}_{i}g$ for all $i=1,\ldots,m$, so that $\partial g:g=(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})$. Recall that the $J_{i}$ (resp. $\tilde{J}_{i}$) do not depend on the chosen presentation, so these ideals are the ones generated by the minors of order $i$ of the Saito matrix with respect to $\{\hat{\eta}_{1},\ldots,\hat{\eta}_{m}\}$ (resp. the extended Saito matrix with respect to $\{\hat{\eta}_{1},\ldots,\hat{\eta}_{m}\}$ and $g$). Thus, $(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})\subset\tilde{J}_{1}$ and it is clear that $\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}\subset\sqrt{\tilde{J}_{1}}$.
Let us prove the other inclusion. The proof starts as in Lemma 2.13: $g$ is integral over the ideal of its partial derivatives (convergent series with the trivial metric are exactly formal power series, so the proof of [18, Satz 5.2] is still valid), so either some $\hat{\beta}_{i}$ is a unit or a power of $g$ belongs to $(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})$.
In the first case, it is clear that $(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})=\tilde{J}_{1}=\widehat{\mathcal{O}}$ and we are done. In the second one, as $D_{n-1}=D$ and $f$ is reduced, $\sqrt{I_{n}}=(f)$ and $\sqrt{J_{n}}=\sqrt{I_{n}^{e}}=(f)^{e}=(g)\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$.
Note that, reasoning as in $(1)$ in Proposition 2.4, we have $\tilde{J}_{i}\subset J_{i}+(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})$, so $\sqrt{\tilde{J}_{i}}\subset\sqrt{J_{i}+(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}=\sqrt{\sqrt{J_{i}}+\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}}$ for all $i$. In particular, as $\sqrt{J_{n}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$, we also have $\sqrt{\tilde{J}_{n}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$. But, as $D$ is strongly Euler-homogeneous outside $D_{0}$, we have that $\sqrt{J_{i}}=\sqrt{\tilde{J}_{i+1}}$ for $i\leq n-1$ (Proposition 4.1). This implies $\sqrt{J_{n-1}}=\sqrt{\tilde{J}_{n}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$ and so
$\sqrt{\tilde{J}_{n-1}}\subset\sqrt{\sqrt{J_{n-1}}+\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}.$
By reverse induction, we finally get $\sqrt{\tilde{J}_{1}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$.
Remark 4.4.
As in the convergent case, we deduce that, when $D$ is strongly Euler-homogeneous outside $D_{0}$, the radical of the ideal $\partial g:g$ does not depend on the choice of formal equation $g$.
We have argued that the Fitting ideals of the $\widehat{\mathcal{O}}$-modules $\widehat{\mathcal{J}}_{f}$ and $\widehat{\mathcal{J}}_{f}/(f)$ are the extended ones of those of $\mathcal{J}_{f}$ and $\mathcal{J}_{f}/(f)$. Since the height of the extension is the same, we deduce from Corollary 3.10 that we can also check (weak or strong) Saito-holonomicity using formal data:
Corollary 4.5.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$. Then:
  • (1)
    $D$ is weakly Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\widehat{\mathcal{J}}_{f})\geq i$ for all $i=3,\ldots,n$.
  • (2)
    $D$ is Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\widehat{\mathcal{J}}_{f}/(f))\geq i+1$ for all $i=2,\ldots,n-1$.
  • (3)
    $D$ is strongly Saito-holonomic if and only if $\hskip 2.84544pt(\tilde{J}_{1}=)\operatorname{Fitt}_{n}(\widehat{\mathcal{J}}_{f})=\widehat{\mathcal{O}}$ and $\operatorname{ht}\operatorname{Fitt}_{i}(\widehat{\mathcal{J}}_{f})\geq i+1$ for all $i=2,\ldots,n-1$.