### v3

# On strong Euler-homogeneity and Saito-holonomicity for complex hypersurfaces. Applications to a conjecture on free divisors

Abstract
We first develop some criteria for a general divisor to be strongly Euler-homogeneous in terms of the Fitting ideals of certain modules. We also study new variants of Saito-holonomicity, generalizing Koszul-free type properties and characterizing them in terms of the same Fitting ideals.
Thanks to these advances, we are able to make progress in the understanding of a conjecture from 2002: a free divisor satisfying the Logarithmic Comparison Theorem (LCT) must be strongly Euler-homogeneous. Previously, it was known to be true only for ambient dimension $n\leq 3$ or assuming Koszul-freeness. We prove it in the following new cases: assuming strong Euler-homogeneity on a punctured neighbourhood of a point; assuming the divisor is weakly Koszul-free; for $n=4$; for linear free divisors in $n=5$.
Finally, we refute a conjecture stating that all linear free divisors satisfy LCT and are strongly Euler-homogeneous.
2020 Mathematics Subject Classification. 32S25, 32S05, 14F40.
Keywords: free divisor, Logarithmic Comparison Theorem, logarithmic vector field, Euler-homogeneity, Saito-holonomicity.
Declarations: The author is supported by a Junta de Andalucía PIF fellowship num. PREDOC_00485, by PID2020-114613GB-I00 and by PID2024-156912NB-I00. He has no further relevant interests to disclose.
1 Introduction
It is well-known that a polynomial $f\in\mathbb{C}[x_{1},\ldots,x_{n}]$ is quasihomogeneous of degree $d$ with respect to the weights $w_{1},\ldots,w_{n}\in\mathbb{Z}_{+}$ (that is, $f(x_{1}^{w_{1}},\ldots,x_{n}^{w_{n}})$ is homogeneous of degree $d$) if and only if it satisfies the Euler equation
$w_{1}x_{1}\frac{\partial f}{\partial x_{1}}+\ldots+w_{n}x_{n}\frac{\partial f}{\partial x_{n}}=df,$
which can be rewritten in terms of derivations as $\delta(f)=f$ with $\delta=\frac{1}{d}\sum_{i=1}^{n}w_{i}x_{i}\partial_{i}$.
If we consider holomorphic functions $f\in\mathbb{C}\{x_{1},\ldots,x_{n}\}$, we may ask which ones satisfy an equation of the form $\delta(f)=f$ with $\delta=\sum_{i=1}^{n}a_{i}\partial_{i}$ for some $a_{i}\in\mathbb{C}\{x_{1},\ldots,x_{n}\}$. In algebraic terms, we are asking whether $f$ belongs to the ideal generated by its partial derivatives. This is what we call Euler-homogeneity.
Any quasihomogeneous polynomial is Euler-homogeneous but, clearly, there are holomorphic functions that are not polynomials and still verify this property (take for example $f=e^{x_{1}}$ with $\delta=\partial_{1}$). However, K. Saito proved in 1971 that, for isolated singularities, quasihomogeneous polynomials are in essence all the examples. Precisely, he proved that, if $f$ defines an isolated singularity in a neighbourhood of 0, then $f$ is Euler-homogeneous if and only if there exists a coordinate system in which $f$ becomes a quasihomogeneous polynomial [Saito71, Theorem 4.1]. For non-isolated singularities, this is no longer true, even for polynomials. The divisor $x_{1}x_{2}(x_{1}+x_{2})(x_{1}+x_{2}x_{3})=0$ in $\mathbb{C}^{3}$, usually called the four lines, is a classical counterexample [2002, Section 4].
Thus, it is interesting to see when this generalization of quasihomogeneity occurs. We are going to address this question from a geometric point of view, studying those hypersurfaces having, at each point, a local equation satisfying this property.
Let us consider a complex analytic manifold $X$ of dimension $n$ and let $D\subset X$ be a divisor (only reduced divisors, i.e. hypersurfaces, and reduced local equations of them will be considered, but sometimes we will omit this for conciseness). Denote by $\mathcal{O}_{X}$ the sheaf of holomorphic functions on $X$, by $\operatorname{Der}_{X}$ the $\mathcal{O}_{X}$-module of $\mathbb{C}$-derivations of $\mathcal{O}_{X}$ and by $\mathscr{D}_{X}$ the sheaf of differential operators on $X$.
Definition 1.1.
A germ of holomorphic function $f\in\mathcal{O}_{X,p}$ is called strongly Euler-homogeneous at $p$ if there exists a germ of derivation $\delta$ vanishing at $p$ (i.e. $\delta\in\mathfrak{m}_{X,p}\operatorname{Der}_{X,p}$, where $\mathfrak{m}_{X,p}$ denotes the maximal ideal of the local ring $\mathcal{O}_{X,p}$) such that $\delta(f)=f$. A divisor $D$ is said to be strongly Euler-homogeneous at $p\in D$ if some (or any) reduced local equation of $D$ at $p$ is strongly Euler-homogeneous and it is called strongly Euler-homogeneous on a subset $E\subset D$ if it is so at any $p\in E$. When $E=D$, we simply say that $D$ is strongly Euler-homogeneous.
Remark 1.2.
In the above definition, the word “strongly” refers to the extra condition that the germ of derivation vanishes at the given point. This is to ensure that the property is intrinsic to the divisor and does not depend on the choice of local equation.
In order to study this property, logarithmic derivations (or vector fields) along $D$ turn out to be very useful. This concept was extensively studied by K. Saito in [Saito80] and refers to those derivations leaving invariant the defining ideal of $D$. These derivations form an $\mathcal{O}_{X}$-module, denoted by $\operatorname{Der}_{X}(-\log D)$. Locally, a derivation $\delta\in\operatorname{Der}_{X,p}$ belongs to $\operatorname{Der}_{X,p}(-\log D)$ if $\delta(f)\in(f)$ for some (or any) reduced local equation $f$ of $D$ at $p$. In this case, we will also say that $\delta$ is a logarithmic derivation for $f$.
Consider a reduced local equation $f$ of $D$ at $p$ and a generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$. Once a local coordinate system $x_{1},\ldots,x_{n}$ is chosen, we define the Saito matrix with respect to $\mathcal{S}$ as $A=(\delta_{i}(x_{j}))_{i,j}$, the $m\times n$ matrix whose entries are the coefficients in the expression of $\delta_{1},\ldots,\delta_{m}$ as linear combinations of $\partial_{1},\ldots,\partial_{n}$. Writing $\overline{\alpha}:=(\alpha_{1},\ldots,\alpha_{m})^{t}$, where $\alpha_{i}\in\mathcal{O}_{X,p}$ is such that $\delta_{i}(f)=\alpha_{i}f$, we define the extended Saito matrix with respect to $\mathcal{S}$ and $f$ as the $m\times(n+1)$ matrix $\tilde{A}=(A\mid-\overline{\alpha})$. Strictly speaking, these matrices also depend on the coordinate system, but we will omit this for the sake of brevity.
When $\operatorname{Der}_{X}(-\log D)$ is a locally free $\mathcal{O}_{X}$-module, we say that $D$ is a free divisor. If $D$ is free, then $\operatorname{Der}_{X,p}(-\log D)$ is free of rank $n$ as an $\mathcal{O}_{X,p}$-module for every $p\in X$. By Saito’s criterion [Saito80, Theorem 1.8], the freeness of $\operatorname{Der}_{X,p}(-\log D)$ is equivalent to the existence of some $\delta_{1},\ldots,\delta_{n}\in\operatorname{Der}_{X,p}(-\log D)$ such that if $f\in\mathcal{O}_{X,p}$ is a reduced local equation at $p$, then there exists a unit $u\in\mathcal{O}_{X,p}$ with $f=u\det(A)$, where $A$ is the Saito matrix with respect to $\{\delta_{1},\ldots,\delta_{n}\}$. In this case, $\{\delta_{1},\ldots,\delta_{n}\}$ is automatically a basis of $\operatorname{Der}_{X,p}(-\log D)$ and we will also say that $f$ is free. This close relation between the equation and logarithmic derivations makes free divisors easier to work with.
After setting some preliminary results in Section 2, we begin Section 3 by defining some analytic closed subsets $D_{i}\subset X$ (resp. $\tilde{D}_{i}\subset X$) as the vanishing locus of certain Fitting ideals and seeing that they can be explicitly described as the set of points at which the rank of $A$ (resp. $\tilde{A}$) is less or equal than $i$. This allows us to give a criterion for strong Euler-homogeneity in a neighbourhood of a point $p$ in terms of these sets. We also give the algebraic translation of this criterion and a necessary condition for a divisor to be strongly Euler-homogeneous outside $D_{0}$.
K. Saito defined in [Saito80, Definition 3.8] holonomic divisors as those for which, at each point, the logarithmic stratification is locally finite and characterized them in terms of the dimensions of the sets $D_{i}$ (he called them $A_{r}$). Later, Calderón-Moreno defined in [Cald1999, Definition 4.1.1] the notion of Koszul-freeness for free divisors in terms of regular sequences in the graded ring $\operatorname{gr}\mathscr{D}_{X}$. Although this was not initially noticed, these two notions turned out to be the same for free divisors [GMNS, Theorem 7.4]. In particular, Koszul-freeness can also be characterized in terms of these sets.
Still in the free setting, in [NarvDual, Definition 1.13] and [SymmetryGS, Remark 7.3], weak and strong Koszul-freeness were introduced (see Definition 4.1). While studying these properties, we asked ourselves if they could be extended to non-free divisors. This is what led us to define, by analogy, weak and strong Saito-holonomicity (Definition 4.3). And we found out that they have a characterization in terms of the dimensions of the sets $\tilde{D}_{i}$ (Theorem 4.6). This comprises the first half of Section 4.
The rest of Section 4 is devoted to study these new versions of Saito-holonomicity and their relation with strong Euler-homogeneity and another interesting property: being of linear Jacobian type. In [NarvDual, Proposition 1.11 and Theorem 4.7], it was shown that being strongly Koszul-free and being of linear Jacobian type are equivalent for free divisors and that, under Koszul-freeness hypothesis, strong Koszul-freeness is equivalent to strong Euler-homogeneity (among other properties). We use the developed criteria for strong Euler-homogeneity and (strong) Saito-holonomicity to generalize these results.
Formal coordinate changes are much more flexible and they permit a particularly nice description of the module of formal logarithmic derivations. Thus, it is interesting to extend the criteria developed in Sections 3 and 4 to the ring of formal power series. This is done in Section 5, whose results will be very helpful in the following section.
Section 6 is devoted to the study of a conjecture on free divisors. But in order to state it, we must first introduce the Logarithmic Comparison Theorem.
Denote by $U=X\setminus D$ the complement of the divisor $D$ in $X$ and let $j:U\to X$ be the inclusion. Let $\Omega^{\bullet}_{X}(*D)$ be the meromorphic de Rham complex, the complex of sheaves (of $\mathbb{C}$-vector spaces) of meromorphic differential forms with poles along $D$, together with the exterior differential.
Grothendieck’s Comparison Theorem [Groth, Theorem 2] states that the natural morphism
$\Omega^{\bullet}_{X}(*D)\to Rj_{*}\mathbb{C}_{U}$
is an isomorphism in the derived category of sheaves of $\mathbb{C}$-vector spaces.
It is a general fact that the hypercohomology of $Rj_{*}\mathbb{C}_{U}$ is isomorphic to $H^{*}(U;\mathbb{C})$, the singular cohomology of $U$. Thus, Grothendieck’s Comparison Theorem tells us that the meromorphic de Rham complex can be used to compute the cohomology of the complement of a divisor. However, the sheaves of $\Omega^{\bullet}_{X}(*D)$ are not coherent $\mathcal{O}_{X}$-modules, so it is interesting to see when one can replace the meromorphic de Rham complex by a quasi-isomorphic complex whose objects are coherent.
K. Saito introduced in [Saito80] logarithmic forms (along $D$) as those meromorphic forms $\omega$ such that $f\omega$ and $fd\omega$ are holomorphic, where $f$ is a local equation of $D$. Logarithmic forms constitute a complex of sheaves of $\mathbb{C}$-vector spaces with the exterior differential (in fact, a subcomplex of $\Omega_{X}^{\bullet}(*D)$), which is called the logarithmic de Rham complex and it is denoted by $\Omega_{X}^{\bullet}(\log D)$. As the $\Omega^{i}_{X}(\log D)$ are coherent $\mathcal{O}_{X}$-modules, it is natural to ask when the inclusion $\Omega^{\bullet}_{X}(\log D)\hookrightarrow\Omega^{\bullet}_{X}(*D)$ is a quasi-isomorphism, so that $\Omega^{\bullet}_{X}(\log D)$ could serve our purpose. This question leads us, by analogy with Grothendieck’s Comparison Theorem, to the following definition:
Definition 1.3.
We say that $D$ satisfies the Logarithmic Comparison Theorem (LCT) if the inclusion
$\Omega^{\bullet}_{X}(\log D)\hookrightarrow\Omega^{\bullet}_{X}(*D)$
is a quasi-isomorphism.
It is clear that, if LCT holds for $D$, then for each $p\in D$, the map $\Omega^{\bullet}_{X,p}(\log D)\hookrightarrow\Omega^{\bullet}_{X,p}(*D)$ is a quasi-isomorphism. However, it is not necessarily true that, if $\Omega^{\bullet}_{X,p}(\log D)\hookrightarrow\Omega^{\bullet}_{X,p}(*D)$ is a quasi-isomorphism, then LCT holds in a neighbourhood of $p$. This motivates the local definition: we say that $D$ satisfies LCT at $p$ or that the germ $(D,p)$ verifies LCT if $D$ satisfies LCT in a neighbourhood of $p$.
P. Deligne proved that LCT holds for normal crossing divisors [Deligne70, Chap. II, Lemme 6.9], a fact that plays a crucial role in the development of mixed Hodge theory. Later, F.J. Castro Jiménez, L. Narváez Macarro and D. Mond showed in [CNM96] that LCT also holds for any locally quasihomogeneous free divisor. H. Terao had conjectured in 1977 that, in fact, any hyperplane arrangement (which are obviously locally quasihomogeneous), free or not, satisfies LCT. This was finally proved in 2024 by D. Bath [BathHypArr]. For a recent detailed survey on LCT, see [SurveyLCT].
In 2002, F.J. Calderón Moreno, D. Mond, L. Narváez Macarro and F.J. Castro Jiménez proved that, in fact, LCT is equivalent to local quasihomogeneity for plane curves [2002, Theorem 1.3]. They also showed that this is not true in higher dimensions, providing the four lines in $\mathbb{C}^{3}$ as a counterexample. Nevertheless, this divisor is still strongly Euler-homogeneous.
In fact, all known examples of divisors satisfying LCT are strongly Euler-homogeneous. The same authors conjectured that this is always the case for free divisors.
Conjecture 1.4.
If $D$ is a free divisor in a complex analytic manifold $X$ of dimension $n$ that satisfies the Logarithmic Comparison Theorem, then it is strongly Euler-homogeneous.
Very little is known about the veracity of this conjecture. Apart from plane curves, for which both properties are equivalent, Conjecture 1.4 is also known to be true for $n=3$. M. Granger and M. Schulze proved it in [GS, Theorem 1.6] with the help of the formal structure theorem for logarithmic vector fields. Later, in 2015, L. Narváez Macarro showed that, for Koszul-free divisors, both properties (and some others) are also equivalent [NarvDual, Theorem 4.7] (see also [Tor04, Corollary 1.8]). In the non-free setting, recent work by D. Bath and M. Saito confirmed that LCT also implies strong Euler-homogeneity for isolated singularities [BathSaitoTLCT, Corollary 1].
One might also ask for the converse. D. Bath and M. Saito also characterized when a quasihomogeneous isolated singularity verifies LCT in terms of the unipotent monodromy part of the vanishing cohomology [BathSaitoTLCT, Remark 1.4a]. This provides several counterexamples to the reverse implication, at least for non-free divisors. However, as far as we are concerned, no free counterexample is known.
In Section 6, we prove Conjecture 1.4 in some new cases. It is divided into three subsections. In Subsection 6.1 we study free divisors that are strongly Euler-homogeneous in a punctured neighbourhood of a point, characterizing when they are strongly Euler-homogeneous at the given point in terms of the existence of non-topologically nilpotent singular logarithmic derivations (Theorem 6.1). As a consequence, we prove a weak version of the conjecture in arbitrary dimension: if $D$ is a germ of free divisor in $(\mathbb{C}^{n},0)$ satisfying LCT and being strongly Euler-homogeneous outside the origin, then it is so at the origin (Theorem 6.3). In Subsection 6.2, we use this result to prove that Conjecture 1.4 also holds in arbitrary dimension for weakly Koszul-free divisors (Theorem 6.5). Subsection 6.3 presents the proof of the conjecture for the $4$-dimensional case (Theorem 6.10), after developing an intrinsic version of the formal structure theorem (Theorem 6.8).
Finally, M. Granger, D. Mond, A. Nieto and M. Schulze posed a conjecture in [GMNS] stating that all linear free divisors satisfy LCT and are strongly Euler-homogeneous, and they proved it for ambient dimension $n\leq 4$. We present in Section 7 a counterexample for $n=5$. Nonetheless, from the previous results, we are also able to show that Conjecture 1.4 holds for linear free divisors in ambient dimension $n=5$.
2 Preliminaries
Within this section, we fix the notation and sum up the main results of [dVR24, Sections 2 and 3], which are going to be used throughout the rest of the text. Due to their different nature, we have divided them into three subsections.
Since the properties we are dealing with are local, it will be enough to consider a germ of free divisor $D=V(f)$ in $(\mathbb{C}^{n},0)$. Let $(\mathcal{O},\mathfrak{m}):=(\mathcal{O}_{\mathbb{C}^{n},0},\mathfrak{m}_{\mathbb{C}^{n},0})$, which can be identified with the convergent power series ring $\mathbb{C}\{x_{1},\ldots,x_{n}\}$. Let $\operatorname{Der}:=\operatorname{Der}_{\mathbb{C}^{n},0}$ be the $\mathcal{O}$-module of $\mathbb{C}$-derivations of $\mathcal{O}$, which is free with a basis given by the partial derivatives $\{\partial_{1},\ldots,\partial_{n}\}$ and let $\operatorname{Der}_{f}:=\operatorname{Der}_{\mathbb{C}^{n},0}(-\log D)$ be the $\mathcal{O}$-module of logarithmic derivations for $f$. Recall that $\operatorname{Der}_{g}=\operatorname{Der}_{f}$ if $g$ is another reduced local equation of $D$ (i.e. $g=uf$ for some unit $u$).
2.1 About products with smooth factors
An important concept that will be useful is that of a product (with a smooth factor, but we will omit this for the sake of brevity):
Definition 2.1.
Let $(D,p)$ be a germ of a divisor in a complex analytic manifold $X$ of dimension $n$. We will say that $(D,p)$ is a product with a smooth factor (from now on, just product) if there exists some germ of divisor $(D^{\prime},0)\subset(\mathbb{C}^{n-1},0)$ such that $(D,p)$ is biholomorphic to $(D^{\prime},0)\times(\mathbb{C},0)$. In other words, if $f\in\mathcal{O}_{X,p}$ is a reduced local equation of $D$ at $p$, then there exists a coordinate change such that $f=ug$ for some unit $u$ and some convergent power series $g$ in n1n-1 variables. In this case, we will also say that $f$ is a product (with a smooth factor) at $p$.
The following is an equivalent algebraic condition:
Lemma 2.2.
Let $f$ be a reduced local equation of a divisor $D\subset X$ at a point $p\in D$. Then, $f$ is a product at $p$ if and only if it admits a non-singular logarithmic derivation, that is, if $\operatorname{Der}_{X,p}(-\log D)$ is not contained in $\mathfrak{m}_{p}\operatorname{Der}_{X,p}$.
Proof.
It can be easily deduced from [Saito80, Lemma 3.5].
Remark 2.3.
In particular, $f\in\mathcal{O}$ is a product (at 0) if and only if $\operatorname{Der}_{f}\not\subset\mathfrak{m}\operatorname{Der}$.
The next lemma (see [CNM96, Lemma 2.2] and [GS, Lemma 3.2]) will allow us, by applying an inductive argument, to just consider the case in which the local equation of the divisor is not a product:
Lemma 2.4.
Let $f\in\mathcal{O}$ be reduced. Suppose $f$ is a product, so that there exists $g\in\mathcal{O}$ depending only on n1n-1 variables in some coordinate system with $f=ug$ for some unit $u$. Then, freeness and strong Euler-homogeneity are equivalent for $f$ and $g$. Moreover, the divisor defined by $f$ in $(\mathbb{C}^{n},0)$ satisfies LCT if and only if the divisor defined by $g$ in $(\mathbb{C}^{n-1},0)$ does.
2.2 From convergent to formal definitions
If we see convergent power series as formal power series we have a wider range of coordinate changes, which gives us more flexibility. Thus, we will also consider $\widehat{\mathcal{O}}=\mathbb{C}[[x_{1},\ldots,x_{n}]]$, the $\mathfrak{m}$-adic completion of $\mathcal{O}$, which is the local ring of formal power series with maximal ideal $\widehat{\mathfrak{m}}=(x_{1},\ldots,x_{n})$.
Let us denote by $\widehat{\operatorname{Der}}$ the $\mathfrak{m}$-adic completion of $\operatorname{Der}$, which coincides with the $\widehat{\mathcal{O}}-$module of $\mathbb{C}$-derivations of $\widehat{\mathcal{O}}$ and is free with a basis given by the partial derivatives. As in the convergent case, for a formal power series $g\in\widehat{\mathcal{O}}$, we will denote by $\operatorname{Der}_{g}$ the $\widehat{\mathcal{O}}-$module of formal logarithmic derivations for $g$, formed by those $\delta\in\widehat{\operatorname{Der}}$ for which $\delta(g)\in(g)$. If a convergent power series $f\in\mathcal{O}$ is seen inside $\widehat{\mathcal{O}}$, then this module is precisely the $\mathfrak{m}$-adic completion of the $\mathcal{O}$-module $\operatorname{Der}_{f}$ and we will denote it by $\widehat{\operatorname{Der}}_{f}$.
Remark 2.5.
As in the convergent case, if $g,h\in\widehat{\mathcal{O}}$ are such that $h=ug$ for some formal unit $u$, then $\operatorname{Der}_{h}=\operatorname{Der}_{g}$.
The notions of being a product and strongly Euler-homogeneous can be extended in a natural way to formal power series:
Definition 2.6.
Let $f\in\widehat{\mathcal{O}}$. We say that $f$ is:
  • (1)
    a product if $\operatorname{Der}_{f}\not\subset\mathfrak{m}\widehat{\operatorname{Der}}$.
  • (2)
    strongly Euler-homogeneous if there exists a formal derivation $\delta\in\mathfrak{m}\widehat{\operatorname{Der}}$ such that $\delta(f)=f$.
The next result (see [dVR24, Propositions 2.1, 2.2 and 2.3]) ensures that these definitions (and that of being reduced) coincide when convergent power series are seen as formal ones. Thus, we will just have to prove that $f$ satisfies the required property as a formal power series (which, in general, will be easier):
Proposition 2.7.
Let $f\in\mathcal{O}$. Then,
  • (1)
    $f$ is reduced as a convergent power series if and only if it is so as a formal power series.
  • (2)
    $f$ is strongly Euler-homogeneous as a convergent power series if and only if it is so as a formal power series.
  • (3)
    $f$ is a product as a convergent power series if and only if it is so as a formal power series.
2.3 About singular derivations
Once a coordinate system is chosen, each derivation $\delta\in\operatorname{Der}$ can be uniquely decomposed as a sum $\delta=\sum_{i=-1}^{\infty}\delta_{i}$, where $\delta_{i}=\sum_{j=1}^{n}a_{ij}\partial_{j}$ with $a_{ij}$ being homogeneous of degree $i+1$ for all $j=1,\ldots,n$. Moreover, $\delta_{0}$, which is called the linear part of $\delta$, can be written as $\underline{x}A\overline{\partial}$, where $A$ is a constant matrix, $\underline{x}=(x_{1},\ldots,x_{n})$ and $\overline{\partial}=(\partial_{1},\ldots,\partial_{n})^{t}$. The same is true for a formal derivation $\delta\in\widehat{\operatorname{Der}}$. We say $\delta\in\operatorname{Der}$ (resp. $\delta\in\widehat{\operatorname{Der}}$) is singular if $\delta_{-1}=0$ or, equivalently, $\delta\in\mathfrak{m}\operatorname{Der}$ (resp. $\delta\in\mathfrak{m}\widehat{\operatorname{Der}}=\widehat{\mathfrak{m}}\widehat{\operatorname{Der}}$).
We are going to make use of the Jordan-Chevalley decomposition for a singular formal derivation developed by R. Gérard and A. Levelt [GL, Théorèmes 1.5, 2.2 and 2.3]. To that end, we need to extend the classical concepts of semisimple and nilpotent endomorphisms of a finite dimensional vector space. This can be done in a proper way for some vector spaces of infinite dimension [GL, §1]. However, we will give here equivalent definitions that will be more practical for our purpose.
A formal singular derivation $\delta\in\mathfrak{m}\widehat{\operatorname{Der}}$ leaves invariant every $\widehat{\mathfrak{m}}^{k}$, so it induces a map in each quotient $\delta^{(k)}:\widehat{\mathcal{O}}/\widehat{\mathfrak{m}}^{k}\to\widehat{\mathcal{O}}/\widehat{\mathfrak{m}}^{k}$.
Definition 2.8.
Let $\delta\in\mathfrak{m}\widehat{\operatorname{Der}}$. We say $\delta$ is:
  • (1)
    semisimple if the induced map $\delta^{(k)}$ is diagonalizable for all $k\in\mathbb{N}$.
  • (2)
    topologically nilpotent if the induced map $\delta^{(k)}$ is nilpotent for all $k\in\mathbb{N}$.
Theorem 2.9 (Jordan-Chevalley decomposition).
Let $\delta$ be a singular derivation of $\widehat{\mathcal{O}}$. Then, there exist two unique commuting singular derivations $\delta_{S}$, $\delta_{N}$ such that $\delta_{S}$ is semisimple, $\delta_{N}$ is topologically nilpotent and $\delta=\delta_{S}+\delta_{N}$. Moreover, there exists a regular system of parameters of $\widehat{\mathcal{O}}$ (i.e. a formal coordinate system) in which $\delta_{S}$ is diagonal (i.e. $\delta_{S}$ is of the form $\sum_{i=1}^{n}\lambda_{i}x_{i}\partial_{i}$).
The following proposition collects some useful facts about semisimple and topologically nilpotent derivations (see [dVR24, Section 3 and Lemma 5.1]):
Proposition 2.10.
Let $\delta\in\mathfrak{m}\widehat{\operatorname{Der}}$ and $f\in\widehat{\mathcal{O}}$, $f\neq 0$. The following holds:
  • (1)
    $\delta$ is topologically nilpotent if and only if the matrix of its linear part is nilpotent. In particular, if $\delta$ has no linear part, then it is topologically nilpotent.
  • (2)
    If $\delta$ is topologically nilpotent and $\delta(f)=\alpha f$ for some $\alpha\in\widehat{\mathcal{O}}$, then $\alpha\in\widehat{\mathfrak{m}}$.
  • (3)
    If $\delta$ is logarithmic for $f$, then so are $\delta_{S}$ and $\delta_{N}$.
  • (4)
    If $\delta$ is semisimple and $\delta(f)=\alpha f$ for some $\alpha\in\widehat{\mathcal{O}}$, then there exists a formal unit $u$ such that, for $g=uf$, $\delta(g)=\alpha_{0}g$, where $\alpha_{0}$ is the constant term of $\alpha$.
We define the trace of a singular (convergent or formal) derivation as the trace of the linear map induced in the quotient $\mathfrak{m}/\mathfrak{m}^{2}$ (or $\widehat{\mathfrak{m}}/\widehat{\mathfrak{m}}^{2}$). In coordinates, this is just the trace of the matrix of its linear part. A key point in the proof of the known cases of Conjecture 1.4 is the existence of singular derivations with non-zero trace when LCT holds. This will also be important in the proof of the new cases:
Theorem 2.11.
If $D$ is a germ of free divisor in $\mathbb{C}^{n}$ satisfying LCT and $f\in\mathcal{O}$ is a reduced local equation of $D$ at 0, then there exist singular derivations in $\operatorname{Der}_{f}$ (and also in $\widehat{\operatorname{Der}}_{f}$) with non-zero trace.
Proof.
It can be deduced from [2002, Sections 2 and 3] and [GS, Lemma 7.5]. See [dVR24, Section 4] for a $\mathscr{D}$-module argument.
3 Strong Euler-homogeneity around a point
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$. We first introduce some analytic sets that will be used throughout the rest of the text and will play a crucial role in the characterization of strong Euler-homogeneity and Saito-holonomicity properties.
Let $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ be a generating set of $\operatorname{Der}_{f}$ (note that $m\geq n$ by coherence, since at a smooth point there cannot be less than $n$ generators). Let $A=(\delta_{i}(x_{j}))_{i,j}$ be the Saito matrix with respect to $\mathcal{S}$. Let $\alpha_{i}\in\mathcal{O}$ be such that $\delta_{i}(f)=\alpha_{i}f$ for $i=1,\ldots,m$ and let $\tilde{A}=(A|-\overline{\alpha})$ be the extended Saito matrix with respect to $\mathcal{S}$ and $f$.
For an $\mathcal{O}$-module $M$, let $\operatorname{Fitt}_{i}(M)$ denote the $i$-th Fitting ideal of $M$. Consider the Jacobian ideal $\mathcal{J}_{f}=\mathcal{O}(\partial_{1}(f),\ldots,\partial_{n}(f),f)$ and the $\mathcal{O}$-module $\mathcal{J}_{f}/(f)$ generated by the classes of the partial derivatives $\partial_{1}(f)],\ldots,[\partial_{n}(f)]$. We set $I_{i}:=\operatorname{Fitt}_{n-i}\left(\mathcal{J}_{f}/(f)\right)$ and $D_{i}:=V(I_{i+1})$.
The rows of $A$ are a generating set of the syzygies of $\partial_{1}(f)],\ldots,[\partial_{n}(f)]$. A free presentation of $\mathcal{J}_{f}/(f)$ as an $\mathcal{O}$-module is then:
$\mathcal{O}^{m}\stackrel{{\scriptstyle\cdot A}}{{\longrightarrow}}\mathcal{O}^{n}\stackrel{{\scriptstyle\underline{e}_{i}\mapsto[\partial_{i}(f)]}}{{\xrightarrow{\hskip 28.45274pt}}}\mathcal{J}_{f}/(f)\to 0.$
Thus, $I_{i}$ is the ideal generated by the minors of order $i$ of $A$ for $1\leq i\leq n$, $I_{i}=\mathcal{O}$ for $i\leq 0$ and $I_{i}=0$ for $i>n$. The set $D_{i}$ is then the vanishing locus of all minors of order $(i+1)$, that is, the set of those points at which $A$ has rank less or equal than $i$:
$D_{i}=\{p\in(\mathbb{C}^{n},0)\mid\operatorname{rank}A(p)\leq i\}.$
Let us see some properties of these sets:
Proposition 3.1.
The following properties hold:
  • (1)
    $D_{0}\subset D_{1}\subset\ldots\subset D_{n-1}\subset D_{n}$.
  • (2)
    $D_{n}=(\mathbb{C}^{n},0)$ and $D_{n-1}=D$.
  • (3)
    If $p\in D\setminus D_{k}$ for some $0\leq k\leq n-2$, then $(D,p)\cong(D^{\prime},0)\times(\mathbb{C}^{k+1},0)$, where $(D^{\prime},0)$ is a germ of divisor in $(\mathbb{C}^{n-k-1},0)$.
    In particular, if $p\in D\setminus D_{0}$, then $(D,p)\cong(D^{\prime},0)\times(\mathbb{C},0)$ and $D$ is a product with a smooth factor at $p$.
Proof.
As $I_{i+1}\subset I_{i}$ for all $i$, $D_{i-1}=V(I_{i})\subset V(I_{i+1})=D_{i}$, which proves $(1)$. Since $A$ is an $m\times n$ matrix with $m\geq n$, it is clear that $\operatorname{rank}A(p)\leq n$ for all $p$, so $D_{n}=(\mathbb{C}^{n},0)$. In [HauserMuller, Proposition 2.2], it is shown in a more general setting that $D=\{p\in(\mathbb{C}^{n},0)\mid\dim\operatorname{Der}_{f}(p)\leq n-1\}$, where $\operatorname{Der}_{f}(p)$ stands for the $\mathbb{C}$-vector subspace of $T_{p}(\mathbb{C}^{n},0)$ generated by $\delta_{1}(p),\ldots,\delta_{m}(p)$. It is clear that $\dim\operatorname{Der}_{f}(p)=\operatorname{rank}A(p)$, so $D=D_{n-1}$ and we get $(2)$.
Let $p\in D\setminus D_{k}$, so $\operatorname{rank}A(p)>k$. Let us see by induction on $k$ that, in a neighbourhood of $p$, at least $k+1$ partial derivatives of $f$ belong to the ideal generated by $f$ and the rest of them.
For $k=0$, at least one of the entries $a_{ij}$ (that corresponds to the coefficient of $\partial_{j}$ in $\delta_{i}$) of $A$ cannot vanish at $p$, so it is a unit in $\mathcal{O}_{X,p}$. Thus, we can isolate $\partial_{j}(f)$ in $\delta_{i}(f)\in(f)$ and see that $\partial_{j}(f)$ belongs to the ideal generated by $f$ and $\partial_{k}(f)$, $k\neq j$.
Suppose the result is true for $k\geq 0$ and let $\operatorname{rank}A(p)>k+1$. As before, we can find $i_{0},j_{0}$ such that $a_{i_{0}j_{0}}(p)\neq 0$. Now set $\delta^{\prime}_{k}=\delta_{k}-\dfrac{a_{kj_{0}}}{a_{i_{0}j_{0}}}\delta_{i_{0}}$ for $k\neq i_{0}$ and $\delta^{\prime}_{i_{0}}=\delta_{i_{0}}$. Let $A^{\prime}$ be the Saito matrix with respect to $\{\delta^{\prime}_{1},\ldots,\delta^{\prime}_{m}\}$ (which is still a generating set) and note that $a^{\prime}_{kj_{0}}(p)=0$ for $k\neq i_{0}$. Since $\operatorname{rank}A^{\prime}(p)=\operatorname{rank}A(p)>k+1$ and $a^{\prime}_{i_{0}j_{0}}(p)=a_{i_{0}j_{0}}(p)\neq 0$, the rank of $A^{\prime}(p)$ is exactly $1+\operatorname{rank}B(p)$, where $B$ is the submatrix of $A^{\prime}$ formed by the entries $a^{\prime}_{ij}$ with $i\neq i_{0}$, $j\neq j_{0}$. Thus, $\operatorname{rank}B(p)>k$ and, by induction hypothesis, at least $k+1$ partial derivatives of $f$ belong to the ideal generated by $f$ and the rest of them except $\partial_{j_{0}}(f)$. But, reasoning as in the case $k=0$, $\partial_{j_{0}}(f)$ also belongs to this ideal, so we conclude that at least $k+2$ partial derivatives of $f$ belong to the ideal generated by $f$ and the rest of them.
Now, by [Saito80, Lemma 3.5], there exist a unit $u$ and a local coordinate change such that $g=uf$ (which is another equation of $D$) only depends on nk1n-k-1 variables. If we set $D^{\prime}=V(g)\subset\mathbb{C}^{n-k-1}$, we get $(3)$.
Remark 3.2.
  • (a)
    For free divisors, the equality $D_{n-1}=D$ can be deduced from Saito’s criterion.
  • (b)
    Recall that strong Euler-homogeneity is preserved under smooth factors. Therefore, if $p\in D\setminus D_{0}$, so that $(D,p)\cong(D^{\prime},0^{\prime})\times(\mathbb{C},0)$, then $D$ is strongly Euler-homogeneous at $p$ if and only if so is $D^{\prime}$ at 0.[GS, Lemma 3.2].
Now, call $\tilde{I}_{i}:=\operatorname{Fitt}_{n+1-i}\left(\mathcal{J}_{f}\right)$ and $\tilde{D}_{i}:=V(\tilde{I}_{i+1})$. The rows of $\tilde{A}$ generate the syzygies of $\partial_{1}(f),\ldots,\partial_{n}(f),f$, so a free presentation of $\mathcal{J}_{f}$ as an $\mathcal{O}$-module is:
$\mathcal{O}^{m}\stackrel{{\scriptstyle\cdot\tilde{A}}}{{\longrightarrow}}\mathcal{O}^{n+1}\stackrel{{\scriptstyle\begin{subarray}{c}\underline{e}_{i}\mapsto\partial_{i}(f),\hskip 2.84544pti=1,\ldots,n\\ \underline{e}_{n+1}\mapsto f\end{subarray}}}{{\xrightarrow{\hskip 56.9055pt}}}\mathcal{J}_{f}\to 0.$
Therefore, $\tilde{I}_{i}$ is the ideal generated by the minors of order $i$ of $\tilde{A}$ for $1\leq i\leq\min\{m,n+1\}$, $\tilde{I}_{i}=\mathcal{O}$ for $i\leq 0$ and $\tilde{I}_{i}=0$ for $i>\min\{m,n+1\}$. As before, $\tilde{D}_{i}$ is the set of those points at which $\tilde{A}$ has rank less or equal than $i$:
$\tilde{D}_{i}=\{p\in(\mathbb{C}^{n},0)\mid\operatorname{rank}\tilde{A}(p)\leq i\}.$
Remark 3.3.
  • (a)
    Note that, if $g$ is another equation of $D$, then $\mathcal{J}_{g}=\mathcal{J}_{f}$ and $\mathcal{J}_{f}/(f)=\mathcal{J}_{g}/(g)$. Also, Fitting ideals do not depend on the choice of presentation. Thus, $D_{i}$ (resp. $\tilde{D}_{i}$) and $I_{i}$ (resp. $\tilde{I}_{i}$), which in principle might depend on the equation of $D$ or on the Saito matrix with respect to the chosen basis, only depend on $D$ itself.
  • (b)
    Note that the ideal generated by the $\alpha_{i}$ can be written as the colon ideal $(\partial_{1}(f),\ldots,\partial_{n}(f)):f$ (from now on, we will just write $\partial f:f$ for short). Whereas $I_{i}$ and $\tilde{I}_{i}$ do not depend on the choice of $f$, the ideal $\partial f:f$ does depend on such a choice. In fact, if $g=uf$ for some unit $u$, then
    $\partial g:g=(\alpha_{1}+\delta_{1}(u)u^{-1},\ldots,\alpha_{m}+\delta_{m}(u)u^{-1}).$
The sets $\tilde{D}_{i}$ and ideals $\tilde{I}_{i}$ are related with the previous ones in the following way:
Proposition 3.4.
The following properties hold:
  • (1)
    $\tilde{I}_{i+1}\subset I_{i+1}+(\alpha_{1},\ldots,\alpha_{m})I_{i}\subset I_{i}\subset\tilde{I}_{i}$   for all $i=1,\ldots,n-1$.
  • (2)
    $\tilde{D}_{i}\subset D_{i}\subset\tilde{D}_{i+1}$   for all $i=0,\ldots,n-2$.
  • (3)
    $\tilde{D}_{n-1}=D_{n-2}=\operatorname{Sing}D$.
Proof.
The minors of order $i+1$ of $\tilde{A}$ are the ones of $A$ together with those involving the last column $\overline{\alpha}$, which are (expanding them by this column) linear combinations of minors of order $i$ of $A$ with coefficients in the $\alpha_{i}$. This justifies the first inclusion of $(1)$. The second inclusion is due to the fact that $I_{i+1}\subset I_{i}$ for all $i$. The last inclusion holds because, $A$ being a submatrix of $\tilde{A}$, the minors of order $i$ of $A$ are also minors of order $i$ of $\tilde{A}$. Since the map $V$ is inclusion-reversing, $(2)$ is immediate from $(1)$.
Let us prove $(3)$. Since $\dim(\operatorname{Sing}D)\leq n-2$ (or $\operatorname{ht}\mathcal{J}_{f}\geq 2$), by [EisenbudFitting, Theorem 5.1], we have that $\sqrt{\mathcal{J}_{f}}=\sqrt{\operatorname{Fitt}_{1}(\mathcal{J}_{f})}=\sqrt{\tilde{I}_{n}}$. Thus,
$\tilde{D}_{n-1}=V(\tilde{I}_{n})=V\left(\sqrt{\tilde{I}_{n}}\right)=V\left(\sqrt{\mathcal{J}_{f}}\right)=V(\mathcal{J}_{f})=\operatorname{Sing}D.$
By $(2)$, $D_{n-2}\subset\tilde{D}_{n-1}=\operatorname{Sing}D$. Now, suppose that there exists $p\in\operatorname{Sing}D\setminus D_{n-2}$. Then, by $(3)$ in Proposition 3.1, $(D,p)\cong(\mathbb{C}^{n-1},0)$. But this is a contradiction because $(\mathbb{C}^{n-1},0)$ is non-singular, so $D_{n-2}=\operatorname{Sing}D$ and we have $(3)$.
Remark 3.5.
Since $D_{n-1}=D$ and $\tilde{D}_{n-1}=\operatorname{Sing}D$, we have that $D_{i}$ and $\tilde{D}_{i}$ are closed analytic subsets of $D$ for $0\leq i\leq n-1$.
Now, we are going to give algebraic and geometric characterizations of strong Euler-homogeneity for points in a neighbourhood of 0.
As $\tilde{D}_{n-1}=D_{n-2}$, we may wonder if it is always true that $\tilde{D}_{i}=D_{i-1}$ for all $i=0,\ldots,n-1$ (where $D_{-1}=V(I_{0})=V(\mathcal{O})=\varnothing$). We are going to see that this property is, in fact, equivalent to strong Euler-homogeneity in a neighbourhood of the origin.
This criterion is based on [GMNS, Lemma 7.5], where the authors give a characterization of strong Euler-homogeneity for germs of free divisors that are strongly Euler-homogeneous at the origin. We have adapted this result to the case in which we do not know a priori whether there exists an Euler vector field.
First, we give a characterization for strong Euler-homogeneity at a point $p$ in terms of the rank of the matrices $A(p)$ and $\tilde{A}(p)$:
Proposition 3.6.
$D$ is strongly Euler-homogeneous at the point $p\in D$ if and only if $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)+1$.
Proof.
Let us note that, as $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ is a generating set of $\operatorname{Der}_{f}$, in a neighbourhood of the origin, $D$ is strongly Euler-homogeneous at $p$ if and only if there exists $a_{1},\ldots,a_{m}\in\mathcal{O}$ such that $\chi=\sum_{i=1}^{m}a_{i}\delta_{i}$ vanishes at $p$ and $\chi(f)=\gamma f$ with $\gamma(p)\neq 0$ (that is, $\sum_{i=1}^{m}a_{i}(p)\alpha_{i}(p)\neq 0$). Moreover, the $\mathbb{C}$-vector space $\widetilde{N}=\{\bar{x}\in\mathbb{C}^{m}\mid\tilde{A}(p)^{t}\bar{x}=\bar{0}\}$ is always contained in the $\mathbb{C}$-vector space $N=\{\bar{x}\in\mathbb{C}^{m}\mid A(p)^{t}\bar{x}=\bar{0}\}$, with equality if and only if $\dim\widetilde{N}=m-\operatorname{rank}\tilde{A}(p)=m-\operatorname{rank}A(p)=\dim N$, that is, if and only if $\operatorname{rank}A(p)=\operatorname{rank}\tilde{A}(p)$.
Observe that we always have $\operatorname{rank}A(p)\leq\operatorname{rank}\tilde{A}(p)\leq\operatorname{rank}A(p)+1$. Let us suppose that $\operatorname{rank}\tilde{A}(p)\neq\operatorname{rank}A(p)+1$, so $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)$ and $\widetilde{N}=N$. Now, if a vector field $\chi=\sum_{i=1}^{m}a_{i}\delta_{i}$ verifies $\chi(p)=0$, then $\bar{a}(p)=(a_{1}(p),\ldots,a_{m}(p))^{t}$ belongs to $N=\widetilde{N}$, so $\sum_{i=1}^{m}a_{i}(p)\alpha_{i}(p)=0$ and there cannot be a strong Euler vector field at $p$.
Reciprocally, if $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)+1$, then $\widetilde{N}$ is strictly contained in $N$. Therefore, there exists $\bar{a}=(a_{1},\ldots,a_{m})^{t}\in\mathbb{C}^{m}$ such that $\sum_{i=1}^{m}a_{i}\delta_{i}(p)=0$ and $\sum_{i=1}^{m}a_{i}\alpha_{i}(p)\neq 0$. Taking $\chi=\sum_{i=1}^{m}a_{i}\delta_{i}$ we get the desired conclusion.
As a consequence of this result, we get a criterion for strong Euler-homogeneity in terms of the analytic sets $D_{i}$ and $\tilde{D}_{i}$:
Theorem 3.7 (Geometric criterion for strong Euler-homogeneity).
Let $D$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$. Then:
  • (1)
    $D$ is always strongly Euler-homogeneous on $D_{0}\setminus\tilde{D}_{0}$ and non-strongly Euler-homogeneous at any point of $\tilde{D}_{0}$. Thus, $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\tilde{D}_{0}=\varnothing$.
  • (2)
    $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\tilde{D}_{i}=D_{i-1}\text{ for }i=1,\ldots,n-2$.
  • (3)
    $D$ is strongly Euler-homogeneous if and only if $\tilde{D}_{0}=\varnothing$ and $\tilde{D}_{i}=D_{i-1}\text{ for }i=1,\ldots,n-2$.
    In particular, for $n=2$, $D$ is strongly Euler-homogeneous if and only if $\tilde{D}_{0}=\varnothing$.
Proof.
For $p\in\tilde{D}_{0}$, $\operatorname{rank}A(p)=\operatorname{rank}\tilde{A}(p)=0$, so $D$ is not strongly Euler-homogeneous at $p$. For $p\in D_{0}\setminus\tilde{D}_{0}$, $A(p)$ vanishes but $\tilde{A}(p)$ does not, so the last column of $\tilde{A}(p)$ is not zero. This means that $\operatorname{rank}A(p)=0$ and $\operatorname{rank}\tilde{A}(p)=1$, so $D$ is strongly Euler-homogeneous at $p$ and we get $(1)$.
Let us prove $(2)$. First, suppose that $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$. As we always have $D_{i-1}\subset\tilde{D}_{i}\subset D_{i}$ for $i=1,\ldots,n-1$ (Proposition 3.4), it is enough to prove $\tilde{D}_{i}\subset D_{i-1}$. Let $p\in\tilde{D}_{i}$ with $1\leq i\leq n-2$ (in particular, $p\in\tilde{D}_{n-2}\subset D$). If $p\in D_{0}$, then $p\in D_{i-1}$ (Proposition 3.1) and we are done. Otherwise, $p\in D\setminus D_{0}$, so $D$ is strongly Euler-homogeneous at $p$ and $\operatorname{rank}A(p)+1=\operatorname{rank}\tilde{A}(p)\leq i$, meaning $\operatorname{rank}A(p)\leq i-1$ and $p\in D_{i-1}$. Reciprocally, suppose $D$ is not strongly Euler-homogeneous at $p\in D\setminus D_{0}$ and let $i_{0}=\operatorname{rank}A(p)=\operatorname{rank}\tilde{A}(p)$. Note that $1\leq i_{0}\leq n-1$ as $p\in D\setminus D_{0}=D_{n-1}\setminus D_{0}$. We have $\operatorname{rank}\tilde{A}(p)\leq i_{0}$ and $\operatorname{rank}A(p)>i_{0}-1$, which implies $p\in\tilde{D}_{i_{0}}\setminus D_{i_{0}-1}$. But we have already proved that $\tilde{D}_{n-1}=D_{n-2}$, so $i_{0}\neq n-1$. Thus, we get $\tilde{D}_{i_{0}}\neq D_{i_{0}-1}$ for some $1\leq i_{0}\leq n-2$.
Finally, $(3)$ is straightforward from $(1)$ and $(2)$.
Remark 3.8.
One might ask why we should care about strong Euler-homogeneity only on $D_{0}$ or $D\setminus D_{0}$. Here is one reason:
Suppose that we want to prove that a divisor satisfying a certain property $P$ is always strongly Euler-homogeneous. Assume we know how to prove it for plane curves and that the property $P$ is preserved under smooth factors (as it happens with strong Euler-homogeneity). Then, we can apply an inductive argument: suppose the result is true for ambient dimension k1k-1 and consider a divisor $D$ in ambient dimension $k$ satisfying $P$. As $D$ is a product with a smooth factor at every point outside $D_{0}$, by induction hypothesis, we deduce that $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$. Thus, we will only have to show that $P$ implies strong Euler-homogeneity on $D_{0}$ (or $\tilde{D}_{0}=\varnothing$ by Theorem 3.7). And any equivalent or necessary condition for strong Euler-homogeneity outside $D_{0}$ could be useful for this purpose.
This strategy will be used in the proof of Theorems 4.12 and 6.10.
Hilbert’s Nullstellensatz gives us an immediate algebraic translation of the developed criterion:
Corollary 3.9 (Algebraic criterion for strong Euler-homogeneity).
Let $D$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$. Then:
  • (1)
    $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\tilde{I}_{1}=\mathcal{O}$.
  • (2)
    $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,\ldots,n-2$.
  • (3)
    $D$ is strongly Euler-homogeneous if and only if $\tilde{I}_{1}=\mathcal{O}$ and $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,\ldots,n-2$.
Remark 3.10.
We have formulated this criterion in terms of ideals of minors of $A$ and $\tilde{A}$, which is more practical for our purpose. In terms of Fitting ideals, this criterion reads as:
  • $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\sqrt{\operatorname{Fitt}_{n}(\mathcal{J}_{f})}=\sqrt{\operatorname{Fitt}_{n}\left(\mathcal{J}_{f}/(f)\right)}\hskip 5.69046pt(=\mathcal{O})$.
  • $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\sqrt{\operatorname{Fitt}_{i}(\mathcal{J}_{f})}=\sqrt{\operatorname{Fitt}_{i}\left(\mathcal{J}_{f}/(f)\right)}$ for all $i=2,\ldots,n-1$.
  • $D$ is strongly Euler-homogeneous if and only if $\sqrt{\operatorname{Fitt}_{i}(\mathcal{J}_{f})}=\sqrt{\operatorname{Fitt}_{i}\left(\mathcal{J}_{f}/(f)\right)}$ for all $i=2,\ldots,n$.
We present now two examples in which we apply the developed criterion to determine if they are strongly Euler-homogeneous or not. Both of them are strongly Euler-homogeneous at the origin but the first one is also strongly Euler-homogeneous outside the origin and the second one is not. Calculations have been made with Macaulay2 [M2]:
Example 3.11.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z+yt)$. A minimal generating set of $\operatorname{Der}_{f}$ is:
[…]
As $\delta_{1}(f)=6f$, $\delta_{2}(f)=yf$, $\delta_{3}(f)=-zf$, $\delta_{4}(f)=(xy+3y^{2}+yz)f$ and $\delta_{5}(f)=(x^{2}+xz+3yz-z^{2})f$, the extended Saito matrix is:
$\tilde{A}=\left(\begin{array}[]{cccc|c}x&y&z&0&-6\\ 0&0&0&z+yt&-y\\ 0&y(x+z)&-z(x+z)&-2t(x+z)&z\\ 0&y^{2}(y+z)&y(x^{2}-z^{2})&-x^{2}+yz+2z^{2}&-xy-3y^{2}-yz\\ 0&yz(y+z)&z(x^{2}-z^{2})&x^{2}t-2z^{2}t+z^{2}&-x^{2}-xz-3yz+z^{2}\\ \end{array}\right).$
By Corollary 3.9, $D$ is strongly Euler-homogeneous if and only if $\tilde{I}_{1}=\mathcal{O}$ (this holds because the last column of $\tilde{A}$ contains a unit) and $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,2$. We have:
[…]
We conclude that $D$ is strongly Euler-homogeneous.
Example 3.12.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z^{2}+yt)$ (note that the only difference with the previous example is that the $z$ in the last term of the product is squared). A minimal generating set of $\operatorname{Der}_{f}$ is:
[…]
We have $\delta_{1}(f)=7f$, $\delta_{2}(f)=yf$, $\delta_{3}(f)=-(x+2z)f$, $\delta_{4}(f)=(xy+2y^{2}-2yz)f$ and $\delta_{5}(f)=(2x^{2}+xz+2yz-2z^{2}+3yt)f$. The extended Saito matrix is then:
$\tilde{A}=\left(\begin{array}[]{cccc|c}x&y&z&t&-7\\ 0&0&0&z^{2}+yt&-y\\ 0&y(x+z)&-z(x+z)&-3t(x+z)&x+2z\\ 0&y^{2}(y+z)&y(x^{2}-z^{2})&-(2x^{2}z+y^{2}t+3yzt)&-xy-2y^{2}+2yz\\ 0&yz(y+z)&z(x^{2}-z^{2})&2x^{2}t-yzt+3yt^{2}&-2x^{2}-xz-2yz+2z^{2}-3yt\\ \end{array}\right)$.
As before, $D$ is strongly Euler-homogeneous if and only if $\tilde{I}_{1}=\mathcal{O}$ (which holds because of the 7-7 in the last column) and $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ for $i=1,2$. But we have:
[…]
This means that $D_{0}\subsetneq\tilde{D}_{1}$ and $D$ is not strongly Euler-homogeneous at any point of $\tilde{D}_{1}\setminus D_{0}=\{(0,0,0,t)\mid t\neq 0\}$, where both matrices $A$ and $\tilde{A}$ have rank $1$.
The following lemma states that the vanishing locus of $\partial f:f$ is always inside the singular locus. This will give us a necessary condition for a divisor to be strongly Euler-homogeneous outside $D_{0}$.
Lemma 3.13.
$V(\partial f:f)\subset\operatorname{Sing}D$.
Proof.
Recall that, with the previous notation, $\partial f:f=(\alpha_{1},\ldots,\alpha_{m})$. As $f$ is integral over the ideal generated by its partial derivatives [Scheja70, Satz 5.2], there exists a logarithmic derivation $\eta$ and an integer $r\geq 1$ such that $\eta(f)=f^{r}$. Let us write $\eta=\sum_{i=1}^{m}c_{i}\delta_{i}$, so that we get $f^{r-1}=\sum_{i=1}^{m}c_{i}\alpha_{i}$ when we apply $\eta$ to $f$. If $r=1$, then $V(\alpha_{1},\ldots,\alpha_{m})=\varnothing$ and there is nothing to prove. Otherwise, it is clear that $V(\alpha_{1},\ldots,\alpha_{m})\subset D$. As $D=D_{n-1}$, if $p\in V(\alpha_{1},\ldots,\alpha_{m})$, then $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)\leq n-1$, so $p\in\tilde{D}_{n-1}=\operatorname{Sing}D$.
Corollary 3.14 (Necessary condition for strong Euler-homogeneity outside $D_{0}$).
If $D=V(f)$ is a germ of divisor in $(\mathbb{C}^{n},0)$ that is strongly Euler-homogeneous on $D\setminus D_{0}$, then $\tilde{D}_{0}=V(\partial f:f)$ and $\sqrt{\tilde{I}_{1}}=\sqrt{\partial f:f}$.
Proof.
It is clear that $\tilde{D}_{0}\subset V(\alpha_{1},\ldots,\alpha_{m})$. For the converse, let $p\in V(\alpha_{1},\ldots,\alpha_{m})$. As the last column of $\tilde{A}$ vanishes at $p$, we have that $\tilde{A}(p)$ and $A(p)$ have the same rank. By Lemma 3.13, $p\in\operatorname{Sing}D=D_{n-2}$, so $\operatorname{rank}A(p)\leq n-2$. Let $i_{0}=\operatorname{rank}\tilde{A}(p)=\operatorname{rank}A(p)\leq n-2$. Then, $i_{0}$ cannot be greater than zero, as then $p\in\tilde{D}_{i_{0}}\setminus D_{i_{0}-1}$ and $D$ would not be strongly Euler-homogeneous outside $D_{0}$ (Theorem 3.7). Therefore, $\operatorname{rank}\tilde{A}(p)=0$ and $p\in\tilde{D}_{0}$. The equality $\sqrt{\tilde{I}_{1}}=\sqrt{\partial f:f}$ follows from the first one by Hilbert’s Nullstellensatz.
This result tells us that, when $D$ is strongly Euler-homogeneous outside $D_{0}$, the radical of $\partial f:f$ does not depend on the choice of $f$.The following example shows that, if strong Euler-homogeneity outside $D_{0}$ is not assumed, this may no longer be true:
Example 3.15.
Let $D=V(f)$ in $(\mathbb{C}^{3},0)$ with $f=xyz(x^{3}+xyz+y^{3}z^{3})$. This divisor is not strongly Euler-homogeneous outside $D_{0}$, as $\sqrt{I_{1}}=(x,y,z)\neq(x,yz)=\sqrt{\tilde{I_{2}}}$. And, if we set $g=(1+y)f$, we have $\sqrt{\partial g:g}=(x,y)\neq(5x^{2}+2yz,5y^{2}z^{2}+2x)=\sqrt{\partial f:f}$.
4 Saito-holonomic type properties
Let us begin this section by recalling the different notions of Koszul-freeness defined in [NarvDual, Definitions 1.10 and 1.13] and [SymmetryGS, Remark 7.3]. We consider the graded ring $\operatorname{gr}\mathscr{D}_{X,p}$ associated with the filtration induced by the order. For a differential operator $P\in\mathscr{D}_{X,p}$, we will denote by $\sigma(P)$ the principal symbol of $P$ in $\operatorname{gr}\mathscr{D}_{X,p}$. It is well-known that $\operatorname{gr}\mathscr{D}_{X,p}$ can be identified with the commutative ring $\mathcal{O}_{X,p}[\underline{\xi}]$, where $\underline{\xi}=(\xi_{1},\ldots,\xi_{n})$ and $\xi_{i}=\sigma(\partial_{i})$.
Definition 4.1.
Let $D\subset X$ be a free divisor. We say $D$ is:
  • weakly Koszul-free at $p\in D$ if, for some (or any) basis $\{\delta_{1},\ldots,\delta_{n}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$, the sequence
    $\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{n})-\alpha_{n}s,\text{ with }\alpha_{i}\in\mathcal{O}_{X,p}\text{ being such that }\delta_{i}(f)=\alpha_{i}f,$
    is regular in $\operatorname{gr}\mathscr{D}_{X,p}[s]$.
  • Koszul-free at $p\in D$ if, for some (or any) basis $\{\delta_{1},\ldots,\delta_{n}\}$ of $\operatorname{Der}_{X,p}(-\log D)$, the sequence
    $\sigma(\delta_{1}),\ldots,\sigma(\delta_{n}),$
    is regular in $\operatorname{gr}\mathscr{D}_{X,p}$.
  • strongly Koszul-free at $p\in D$ if, for some (or any) basis $\{\delta_{1},\ldots,\delta_{n}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$, the sequence
    $f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{n})-\alpha_{n}s,\text{ with }\alpha_{i}\in\mathcal{O}_{X,p}\text{ being such that }\delta_{i}(f)=\alpha_{i}f,$
    is regular in $\operatorname{gr}\mathscr{D}_{X,p}[s]$.
We say $D$ is weakly Koszul-free, Koszul-free or strongly Koszul-free if it is so at each $p\in D$.
Our purpose is to define and study analogous properties in the general case in which our divisor is not necessarily free. In order to do this, we need first to make some algebraic considerations:
In a Cohen-Macaulay local ring, a sequence of $r$ elements is regular if and only if the height of the ideal they generate is $r$ [MatsumuraGris, Theorem 17.4]. The ring $\mathcal{O}_{X,p}[\underline{\xi}]=\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ (resp. $\mathcal{O}_{X,p}[\underline{\xi},s]$) is not local but we will see that, for homogeneous elements with respect to $\underline{\xi}$ (resp. with respect to $\underline{\xi},s$), this is still true. In order to do so, we will study the extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}\{\underline{x},\underline{\xi}\}$.
If a ring map $A\to B$ is faithfully flat, then a sequence of elements is regular in $A$ if and only if the corresponding sequence is regular in $B$. Moreover, if $A$ and $B$ are Noetherian, then the height of an ideal $I\subset A$ (denoted by $\operatorname{ht}I$) coincides with the height of its extension $B\cdot I$ in $B$ [MatsumuraRosa, (4.C), (5.D) and (13.B.3)]. The extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}\{\underline{x},\underline{\xi}\}$ is not faithfully flat but, for homogeneous elements and ideals, these statements still hold. For the sake of completeness, we include the proof here:
Proposition 4.2.
Let $a_{1},\ldots,a_{r}\in\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ be homogeneous elements with respect to $\underline{\xi}$ and consider the ring extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}\{\underline{x},\underline{\xi}\}$. Then:
  • (1)
    $a_{1},\ldots,a_{r}$ is a regular sequence in $\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ if and only if it is so in $\mathbb{C}\{\underline{x},\underline{\xi}\}$.
  • (2)
    If $I=(a_{1},\ldots,a_{r})$, then $\operatorname{ht}I=\operatorname{ht}I^{e}$, where $I^{e}$ denotes the extended ideal.
Consequently, $a_{1},\ldots,a_{r}$ is a regular sequence in $\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ if and only if $\operatorname{ht}(a_{1},\ldots,a_{r})=r$.
Proof.
Let $A=\mathbb{C}\{\underline{x}\}[\underline{\xi}]$. As the extension $\mathbb{C}\{\underline{x},\underline{\xi}\}\subset\mathbb{C}[[\underline{x},\underline{\xi}]]$ is faithfully flat, we just need to prove the statements for the extension $\mathbb{C}\{\underline{x}\}[\underline{\xi}]\subset\mathbb{C}[[\underline{x},\underline{\xi}]]$. Let $M=(\underline{x},\underline{\xi})$ be the maximal homogeneous ideal of $A$, so that the $M$-adic completion of $A$ is $\hat{A}=\mathbb{C}[[\underline{x},\underline{\xi}]]$, and let $S=1+M$. By [MatsumuraRosa, (24.A) and (24.B)], the canonical map $A\to\hat{A}$ factors through $S^{-1}A$ and the extension $S^{-1}A\to\hat{A}$ is faithfully flat. Thus, it remains to show the results for the extension $A\to S^{-1}A$:
It is easy to see that, if $J$ is a homogeneous ideal, $a\in A$, $s\in S$ and $as\in J$, then $a\in J$. From this, we obtain that $(S^{-1}J)^{c}=J$, where the superindex $c$ means the contraction of the ideal. This allows us to deduce that $a_{1},\ldots,a_{r}$ is a regular sequence in $\mathbb{C}\{\underline{x}\}[\underline{\xi}]$ if and only if so is $a_{1}/1,\ldots,a_{r}/1$ in $S^{-1}A$, which gives $(1)$.
Let us prove $(2)$. Consider a homogeneous (with respect to $\underline{\xi}$) prime ideal $\mathfrak{p}$. Note that $S$ cannot meet $\mathfrak{p}$ because the 0-th degree part of an element in the intersection would be a unit belonging to $\mathfrak{p}$. Consequently, we have $\operatorname{ht}\mathfrak{p}=\operatorname{ht}(S^{-1}\mathfrak{p})$ [BouAC89, VIII.9, Corollaire after Proposition 7]. As minimal primes over homogeneous ideals are homogeneous [BouAC89, VIII.64, Lemme 1.c)], we deduce that $\operatorname{ht}I=\operatorname{ht}(S^{-1}I)$, which finishes the proof.
The last statement follows from the previous ones and the fact that $\mathbb{C}\{\underline{x},\underline{\xi}\}$ is a Cohen-Macaulay local ring.
Now, we are ready to generalize Koszul type properties:
Definition 4.3.
Let $D\subset X$ be a divisor. We say:
  • $D$ is weakly Saito-holonomic at $p\in D$ if, for some (or any) generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$,
    $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n\text{ in }\operatorname{gr}\mathscr{D}_{X,p}[s],$
    where $\alpha_{i}\in\mathcal{O}_{X,p}$ is such that $\delta_{i}(f)=\alpha_{i}f.$
  • $D$ is Saito-holonomic at $p\in D$ if, for some (or any) generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$,
    $\operatorname{ht}(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))=n\text{ in }\operatorname{gr}\mathscr{D}_{X,p}.$
  • $D$ is strongly Saito-holonomic at $p\in D$ if, for some (or any) generating set $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ of $\operatorname{Der}_{X,p}(-\log D)$ and some (or any) reduced local equation $f$ of $D$ at $p$,
    $\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1\text{ in }\operatorname{gr}\mathscr{D}_{X,p}[s],$
    where $\alpha_{i}\in\mathcal{O}_{X,p}$ is such that $\delta_{i}(f)=\alpha_{i}f.$
We say $D$ is weakly Saito-holonomic, Saito-holonomic or strongly Saito-holonomic if it is so at each $p\in D$.
Remark 4.4.
  • (a)
    For free divisors, these definitions coincide with those of Koszul-freeness given in Definition 4.1 by virtue of Proposition 4.2.
  • (b)
    In [GMNS, Theorem 7.4], the authors state that holonomic free divisors in the sense of K. Saito (see [Saito80, Definition 3.8]) are exactly Koszul-free divisors. The same argument is valid in the general case to conclude that our definition of holonomicity is equivalent to the one given by K. Saito. This explains the name we have chosen for this property, in agreement with [Uli17, Definition 2.5].
By coherence, if $D$ satisfies any of the properties in Definition 4.3 at $p$, then it does so in a neighbourhood of $p$. This will allow us to characterize divisors in $(\mathbb{C}^{n},0)$ satisfying them in a neighbourhood of 0 just by looking at the dimensions of the germs $D_{i}$ and $\tilde{D}_{i}$ defined in the previous section. In order to do this, we need first a technical lemma:
Lemma 4.5.
Let $E$ be a non-empty analytic set in $(\mathbb{C}^{n},0)$ given by $k$ equations with $0\leq k\leq n$ and let $B(x)$ be an $m\times r$ matrix with entries in $\mathcal{O}$. Let $W=\{(x,y)\in(E,0)\times(\mathbb{C}^{r},0)\mid B(x)y=0\}$ and let $B_{i}=\{p\in(E,0)\mid\operatorname{rank}B(p)\leq i\}$. Then, $\dim W\leq r-k$ if and only if $B_{i}=\varnothing$ for all $i<k$ and $\dim B_{i}\leq i-k$ for all $k\leq i\leq s$, where $s=\min\{m,r\}$. In particular, as $B_{s}=E$, this can only hold when $s\geq n$.
Proof.
Let us consider the canonical projection $\pi:W\to E$. Note that it is surjective because $(p,0)\in W$ for every $p\in E$, so $\pi(\pi^{-1}(B_{i}))=B_{i}$ for all $i$. Let $0\leq i\leq s$ be such that $B_{i}\neq\varnothing$, so that $\dim B_{i}\geq 0$. For every $p\in B_{i}$, the dimension of $\pi^{-1}(p)$ is at least rir-i (because the dimension of $\{y\in(\mathbb{C}^{r},0)\mid B(p)y=0\}$ is $r-\operatorname{rank}B(p)$ and $\operatorname{rank}B(p)\leq i$), so $\dim\pi^{-1}(B_{i})\geq\dim B_{i}+r-i$. Now, if $\dim W\leq r-k$, then
$r-k\geq\dim W\geq\dim\pi^{-1}(B_{i})\geq\dim B_{i}+r-i\hskip 5.69046pt(\geq r-i).$
In particular, $i\geq k$ (so $B_{i}=\varnothing$ for $i<k$) and $\dim B_{i}\leq i-k$.
Note that $B_{s}=E$. Since $E$ is given by $k$ equations, $\dim E\geq n-k$ so we get $n-k\leq\dim E=\dim B_{s}\leq s-k$ and $s\geq n$.
To prove the converse, consider an irreducible component $W^{\prime}$ of $W$. We can distinguish two cases:
  • If $W^{\prime}\subset\pi^{-1}(B_{0})$, in particular, $B_{0}\neq\varnothing$, so $k=0$ and $\dim B_{0}=0$ (by hypothesis). As $p\in B_{0}$ implies $\operatorname{rank}B(p)=0$ and $\dim\pi^{-1}(p)=r$, we get $\dim W^{\prime}\leq\dim\pi^{-1}(B_{0})=r=r-k$.
  • If $W^{\prime}\not\subset\pi^{-1}(B_{0})$, as $W^{\prime}\subset W=\pi^{-1}(E)=\pi^{-1}(B_{s})$, there exists $i\leq s$ for which $W^{\prime}\subset\pi^{-1}(B_{i})\setminus\pi^{-1}(B_{i-1})$. In particular, $B_{i}\setminus B_{i-1}$ (which is the set of points $p\in E$ such that $\operatorname{rank}B(p)=i$) is a non-empty open set of $B_{i}$ where the fibre dimension of $\pi:W^{\prime}\to B_{i}$ attains the minimum value $\lambda(\pi)=r-i$. Since $W^{\prime}$ is irreducible, we have that $\dim W^{\prime}=\lambda(\pi)+\dim\pi(W^{\prime})\leq r-i+\dim B_{i}\leq r-k$, where the last inequality holds by hypothesis.
We conclude that every irreducible component of $W^{\prime}$ verifies $\dim W^{\prime}\leq r-k$ and, thus, $\dim W\leq r-k$.
In what follows, we will need to consider the following map of graded algebras:
$\begin{array}[]{cccccc}\varphi:&\mathcal{O}[\underline{\xi},s]&\longrightarrow&\operatorname{\mathcal{R}}(\mathcal{J}_{f})\\ &\xi_{i}&\mapsto&\partial_{i}(f)\cdot t\\ &s&\mapsto&f\cdot t\end{array}$
where $\operatorname{\mathcal{R}}(\mathcal{J}_{f})=\bigoplus_{i=0}^{\infty}\mathcal{J}_{f}^{i}t^{i}$ is the Rees algebra of $\mathcal{J}_{f}$. As $\varphi$ is surjective and $\operatorname{\mathcal{R}}(\mathcal{J}_{f})$ is an integral domain, $\ker\varphi$ is a prime ideal. Since the dimension of $\mathcal{O}[\underline{\xi},s]$ is $2n+1$ and that of $\operatorname{\mathcal{R}}(\mathcal{J}_{f})$ is $n+1$ (cf. [ReesDim, Theorem 5.1.4]), we deduce that $\operatorname{ht}(\ker\varphi)=n$.
The homogeneous part of degree $1$ of $\ker\varphi$ is generated as an $\mathcal{O}$-module by the set $\{\sigma(\delta_{i})-\alpha_{i}s,i=1,\ldots,m\}$ (in the notation of Definition 4.3). The ideal of $\mathcal{O}[\underline{\xi},s]$ generated by these elements is denoted by $\ker^{(1)}\varphi$. A divisor $D$ is said to be of linear Jacobian type if $\ker\varphi=\ker^{(1)}\varphi$.
Theorem 4.6.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$. Then:
  • (1)
    $D$ is weakly Saito-holonomic if and only if $\dim\tilde{D}_{i}\leq i$ for all $i=0,\ldots,n-3$.
  • (2)
    $D$ is Saito-holonomic if and only if $\dim D_{i}\leq i$ for all $i=0,\ldots,n-3$.
  • (3)
    $D$ is strongly Saito-holonomic if and only if $\tilde{D}_{0}=\varnothing$ and $\dim\tilde{D}_{i}\leq i-1$ for all $i=1,\ldots,n-2$.
Proof.
Let $\mathcal{S}=\{\delta_{1},\ldots,\delta_{m}\}$ be a generating set of $\operatorname{Der}_{f}$ and let $A$ be the Saito matrix with respect to $\mathcal{S}$. Let $\alpha_{i}\in\mathcal{O}$ such that $\delta_{i}(f)=\alpha_{i}f$ for all $i=1,\ldots,m$ and let $\tilde{A}$ be the extended Saito matrix. By coherence, it is enough to show these properties hold at 0:
  • (1)
    $D$ is weakly Saito-holonomic at 0 if and only if $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n$ in $\mathcal{O}[\underline{\xi},s]$. As $(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)$ is a homogeneous ideal (with respect to $\xi_{1},\ldots,\xi_{n},s$), by Proposition 4.2, this is equivalent to $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n$ in $\mathbb{C}\{\underline{x},\underline{\xi},s\}$ or even $\dim W=n+1$, where $W=V(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)\subset(\mathbb{C}^{2n+1},0)$. As $(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=\ker^{(1)}\varphi\subset\ker\varphi$ and $\ker\varphi$ is a prime ideal of $\mathcal{O}[\underline{\xi},s]$ of height $n$, the inequality $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)\leq n$ (or $\dim W\geq n+1$) always holds. Thus, $D$ is weakly Saito-holonomic at 0 if and only if $\dim W\leq n+1$. Note that we can write
    $W=\{(x,(\xi,s))\in(\mathbb{C}^{n},0)\times(\mathbb{C}^{n+1},0)\mid\tilde{A}(x)(\xi,s)^{t}=0\}.$
    Now, we just need to apply Lemma 4.5. With the notation of the lemma, $E=\mathbb{C}^{n},k=0$, $B(x)=\tilde{A}(x)$, $r=n+1$, $s=\min\{m,n+1\}\in\{n,n+1\}$ (recall that $m\geq n$) and $B_{i}=\tilde{D}_{i}$. Thus, $\dim W\leq n+1$ if and only if $\dim\tilde{D}_{i}\leq i$ for all $i=0,\ldots,s$. The conditions $\dim\tilde{D}_{n}\leq n$ and $\dim\tilde{D}_{n+1}\leq n+1$ (if needed) always holds because $\tilde{D}_{n}\subset\tilde{D}_{n+1}\subset\mathbb{C}^{n}$. Moreover, since $\tilde{D}_{n-2}\subset\tilde{D}_{n-1}=\operatorname{Sing}D$ and the singular locus of $D$ is of dimension at most n2n-2, we have $\dim\tilde{D}_{n-2}\leq\dim\tilde{D}_{n-1}\leq n-2<n-1$. We conclude that $D$ is weakly Saito-holonomic if and only if $\dim\tilde{D}_{i}\leq i$ for all $i=0,\ldots,n-3$.
  • (2)
    $D$ is Saito-holonomic at 0 if and only if $\operatorname{ht}(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))=n$ in $\mathcal{O}[\underline{\xi}]$, or in $\mathbb{C}\{\underline{x},\underline{\xi}\}$ by Proposition 4.2. Since $(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))\subset(\xi_{1},\ldots,\xi_{n})$, we get $\operatorname{ht}(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))\leq n$. Thus, we just need to consider the other inequality, which is equivalent to $\dim W\leq n$, where $W=V(\sigma(\delta_{1}),\ldots,\sigma(\delta_{m}))=\{(x,\xi)\in(\mathbb{C}^{n},0)\times(\mathbb{C}^{n},0)\mid A(x)\xi^{t}=0\}$. We proceed as in $(1)$ with $E=\mathbb{C}^{n},k=0,B(x)=A(x)$, $r=n$, $s=\min\{m,r\}=n$ and $B_{i}=D_{i}$ to obtain that $\dim W\leq n$ if and only if $\dim D_{i}\leq i$ for all $0\leq i\leq n$. Since $D_{n}=\mathbb{C}^{n}$, $D_{n-1}=D$ and $D_{n-2}=\operatorname{Sing}D$, the last three conditions are automatically satisfied.
  • (3)
    $D$ is strongly Saito-holonomic at 0 if and only if $\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1$ in $\mathcal{O}[\underline{\xi},s]$ or, reasoning as before, if $\dim W=n$, where $W=V(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=\{(x,(\xi,s))\in D\times(\mathbb{C}^{n+1},0)\mid\tilde{A}(x)(\xi,s)^{t}=0\}$. Note that
    $\dim W\geq\dim V(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)-1\geq n,$
    where the last inequality was argued in $(1)$. Thus, $D$ is strongly Saito-holonomic if and only if $\dim W\leq n$, where $W=V(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=\{(x,(\xi,s))\in D\times(\mathbb{C}^{n+1},0)\mid\tilde{A}(x)(\xi,s)^{t}=0\}$. We apply again Lemma 4.5 with $E=D,k=1,B(x)=\tilde{A}(x)$, $r=n+1$ and $s=\min\{m,n+1\}\in\{n,n+1\}$ to obtain that $\dim W\leq n$ if and only if $B_{0}=\varnothing$ and $\dim B_{i}\leq i-1$ for all $i=1,\ldots,s$. Now, note that $B_{i}=\{p\in D\mid\operatorname{rank}\tilde{A}(p)\leq i\}=\tilde{D}_{i}$ for $i\leq n-1$, since $\tilde{D}_{i}\subset D$ for $i\leq n-1$ (observe that in $\tilde{D}_{i}$ we allow $p$ to be in principle any point of $(\mathbb{C}^{n},0)$, while in $B_{i}$ it must be a point of $D$). Since $B_{n-1}=\tilde{D}_{n-1}=\operatorname{Sing}D$ and $B_{n}\subset B_{n+1}=D$, the condition $\dim B_{i}\leq i-1$ is always satisfied for $i\in\{n-1,n,n+1\}$. We conclude that $D$ is strongly Saito-holonomic if and only if $\tilde{D}_{0}=\varnothing$ and $\dim\tilde{D}_{i}\leq i-1$ for $i=1,\ldots,n-2$.
Remark 4.7.
  • (a)
    As $\tilde{D}_{i}\subset D_{i}\subset\tilde{D}_{i+1}$, it is now clear that strongly Saito-holonomic $\Rightarrow$ Saito-holonomic $\Rightarrow$ weakly Saito-holonomic.
  • (b)
    If $D$ is Saito-holonomic or weakly Saito-holonomic in a punctured neighbourhood of a point $p$, then it is so at $p$. This is because outside the fibre at $p$, the set $W$ (in the notation of the proof of Theorem 4.6) has the required dimension ($n$ or $n+1$, respectively) and the fibre at $p$ (of maximal dimension $n$ or $n+1$, respectively) cannot increase that dimension.
    This argument cannot be applied to strong Saito-holonomicity, since we need $W$ to be $n$-dimensional and the fibre at $p$ could be of dimension $n+1$. In fact, any non-quasihomogeneous germ of plane curve (which is a non-strongly Euler-homogeneous isolated singularity) is a counterexample (see Theorem 4.12).
  • (c)
    The characterization given by K. Saito for holonomic divisors in terms of the dimension of the sets $D_{i}$ ($A_{r}$ in his notation) coincides with the one given in Theorem 4.6 for Saito-holonomic divisors (see [Saito80, Definition 3.12 and Lemma 3.13]). This confirms the equivalence of the two definitions.
Some straightforward consequences of Theorem 4.6 in low dimension are:
Corollary 4.8.
Let $D$ be a germ of divisor in $(\mathbb{C}^{n},0)$.
  • (1)
    If $n=2$, then $D$ is Saito-holonomic [Cald1999, Corollary 4.2.2].
  • (2)
    If $n=3$ and $D$ is strongly Euler-homogeneous, then it is weakly Saito-holonomic.
Proof.
For $n=2$, the conditions for Saito-holonomicity in Theorem 4.6 are automatically satisfied. For $n=3$, $D$ is weakly Saito-holonomic if and only if $\dim\tilde{D}_{0}\leq 0$. But strong Euler-homogeneity implies $\tilde{D}_{0}=\varnothing$, so this holds.
As it happens with strong Euler-homogeneity and Koszul-freeness [CN02, Proposition 1.10], the three versions of Saito-holonomicity are well-behaved with respect to smooth factors:
Proposition 4.9.
Let $D^{\prime}$ be a germ of divisor in $(\mathbb{C}^{n-1},0)$ and let $D=D^{\prime}\times\mathbb{C}\subset(\mathbb{C}^{n},0)$. Then $D$ is weakly Saito-holonomic, Saito-holonomic or strongly Saito-holonomic if and only if so is $D^{\prime}$.
Proof.
Let $D^{\prime}=V(g)$, where $g\in\mathbb{C}\{x_{2},\ldots,x_{n}\}$ and let $f(x_{1},\ldots,x_{n})=g(x_{2},\dots,x_{n})$, so that $D=V(f)$. If $\mathcal{S}^{\prime}=\{\delta_{1},\ldots,\delta_{m}\}$ is a generating set for $\operatorname{Der}_{g}$, then $\mathcal{S}=\{\partial_{1},\delta_{1},\ldots,\delta_{m}\}$ is a generating set for $\operatorname{Der}_{f}$ (where $\partial_{1}(f)=0$). Thus, if $A^{\prime}$ and $\tilde{A}^{\prime}$ are the Saito matrices for $D^{\prime}$ with respect to $\mathcal{S}^{\prime}$ (and $g$), then the Saito matrices for $D$ with respect to $\mathcal{S}$ (and $f$) are:
[…]
It is clear that $\operatorname{rank}A(p)=\operatorname{rank}A^{\prime}(p)+1$ and $\operatorname{rank}\tilde{A}(p)=\operatorname{rank}\tilde{A}^{\prime}(p)+1$. Consequently, we have the relations $D_{i}=D^{\prime}_{i-1}\times\mathbb{C}$ and $\tilde{D}_{i}=\tilde{D}^{\prime}_{i-1}\times\mathbb{C}$ for $i\geq 1$. Observe that $D_{i}=\varnothing$ if and only if $D^{\prime}_{i-1}=\varnothing$ and similarly for $\tilde{D}_{i}$ and $\tilde{D}^{\prime}_{i-1}$. Note also that $D_{0}=\tilde{D}_{0}=\varnothing$. Thus, if they are not empty, $\dim D_{i}\leq k$ if and only if $\dim D^{\prime}_{i-1}\leq k-1$ and $\dim\tilde{D}_{i}\leq k$ if and only if $\dim\tilde{D}^{\prime}_{i-1}\leq k-1$. This proves the result for weakly Saito-holonomic and Saito-holonomic divisors if we take $k=i$ by virtue of Theorem 4.6. Taking k=i1k=i-1 and noticing that if $\dim\tilde{D}_{1}=\dim(\tilde{D}_{0}^{\prime}\times\mathbb{C})\leq 0$, then $\tilde{D}_{0}^{\prime}=\varnothing$, we prove the result for strong Saito-holonomicity.
Remark 4.10.
An inductive argument shows that if $D=D^{\prime}\times\mathbb{C}^{k}$, where $D^{\prime}$ is a divisor in $\mathbb{C}^{n-k}$, then $D_{i}=D^{\prime}_{i-k}\times\mathbb{C}^{k}$ and $\tilde{D}_{i}=\tilde{D}^{\prime}_{i-k}\times\mathbb{C}^{k}$ for $i\geq k$. and $D_{i}=\tilde{D}_{i}=\varnothing$ for $i<k$.
A direct algebraic translation of Theorem 4.6 is the following:
Corollary 4.11.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$. Then:
  • (1)
    $D$ is weakly Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\mathcal{J}_{f})\geq i$ for all $i=3,\ldots,n$.
  • (2)
    $D$ is Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\mathcal{J}_{f}/(f))\geq i+1$ for all $i=2,\ldots,n-1$.
  • (3)
    $D$ is strongly Saito-holonomic if and only if $\hskip 2.84544pt(\tilde{I}_{1}=)\operatorname{Fitt}_{n}(\mathcal{J}_{f})=\mathcal{O}$ and $\operatorname{ht}\operatorname{Fitt}_{i}(\mathcal{J}_{f})\geq i+1$ for all $i=2,\ldots,n-1$.
Proof.
To prove $(1)$ and $(3)$ we use that $\tilde{D}_{i}=V(\tilde{I}_{i+1})=V(\operatorname{Fitt}_{n-i}(\mathcal{J}_{f}))$, so $\dim\tilde{D}_{i}\leq k$ if and only if $\operatorname{ht}\operatorname{Fitt}_{n-i}(\mathcal{J}_{f})\geq n-k$. We proceed similarly for $(2)$ with $D_{i}=V(I_{i+1})=V(\operatorname{Fitt}_{n-i-1}(\mathcal{J}_{f}/(f)))$.
In the free case, we know that, if a divisor is Koszul-free, then it is strongly Koszul-free if and only if it is strongly Euler-homogeneous [NarvDual, Theorem 4.7]. We are now able to generalize this result to the non-free case:
Theorem 4.12.
Let $D\subset X$ be a Saito-holonomic divisor. Then, it is strongly Saito-holonomic if and only if it is strongly Euler-homogeneous.
Proof.
These properties being local, we may assume $D$ is a germ of divisor in $(\mathbb{C}^{n},0)$. If $D$ is strongly Euler-homogeneous, then $\tilde{D}_{0}=\varnothing$ and $D_{i-1}=\tilde{D}_{i}$ for all $i=1,\ldots,n-1$. Since $D$ is also Saito-holonomic, $\dim D_{i}\leq i$ for all $i=0,\ldots,n-3$, so $\dim\tilde{D}_{i}=\dim D_{i-1}\leq i-1$ for all $i=1,\ldots,n-2$ and $D$ is strongly Saito-holonomic.
To prove the converse, we will proceed by induction on $n$. For $n=2$, by Theorem 4.6, a divisor is strongly Saito-holonomic if and only if $\tilde{D}_{0}=\varnothing$. By Theorem 3.7, this happens if and only if it is strongly Euler-homogeneous. Suppose the result is true for n1n-1 and consider a strongly Saito-holonomic divisor $D$ in $(\mathbb{C}^{n},0)$. Since strong Saito-holonomicity is well-behaved with respect to smooth factors (Proposition 4.9), by Remark 3.8, we only have to show that $\tilde{D}_{0}=\varnothing$. But this is guaranteed for a strongly Saito-holonomic divisor by Theorem 4.6.
Remark 4.13.
There are plenty of results in [Uli17] in which the divisor is assumed to be Saito-holonomic and strongly Euler-homogeneous. By Theorem 4.12, all of them can be reformulated in terms of strong Saito-holonomicity.
In the free case, it is known that being of linear Jacobian type is equivalent to being strongly Koszul-free [NarvDual, Proposition 1.11]. In the non-free case, we only have the following:
Proposition 4.14.
Let $D\subset X$ be a divisor of linear Jacobian type. Then, it is strongly Saito-holonomic.
Proof.
We may suppose $D$ is a germ of divisor in $(\mathbb{C}^{n},0)$. As $D$ is of linear Jacobian type, we have that $I:=\ker^{(1)}\varphi=\ker\varphi$. But $\ker\varphi$ is a prime ideal of height $n$ and $f\notin\ker\varphi$, so $\operatorname{ht}(I+(f))=\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1$. Thus, $D$ is strongly Saito-holonomic.
Example 4.15.
The converse is not true. Consider the hyperplane arrangement $D=V(f)$, where $f=xyz(x+t)(y+t)(z+t)(x+y+t)(x+z+t)(y+z+t)$ in $(\mathbb{C}^{4},0)$. It is well-known that all hyperplane arrangements are Saito-holonomic and obviously strongly Euler-homogeneous. Thus, they are strongly Saito-holonomic by Theorem 4.12. However, $D$ is an example of a hyperplane arrangement that is not of differential linear type (that is, the $\mathscr{D}[s]$-annihilator of $f^{s}$ is not generated by operators of order one) [Uli17, Example 5.7], so it cannot be of linear Jacobian type [CN09, Proposition 1.15].
Nevertheless, if we impose an extra condition on $\ker^{(1)}\varphi$, we have the converse:
Proposition 4.16.
Let $D$ be a germ of strongly Saito-holonomic divisor in $(\mathbb{C}^{n},0)$. If $\ker^{(1)}\varphi$ is unmixed (that is, all its associated primes have the same height), then $D$ is of linear Jacobian type.
Proof.
As $D$ is strongly Saito-holonomic, in particular, it is weakly Saito-holonomic. Thus, $\operatorname{ht}(f,\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n+1$ and $\operatorname{ht}(\sigma(\delta_{1})-\alpha_{1}s,\ldots,\sigma(\delta_{m})-\alpha_{m}s)=n$. That is, $\operatorname{ht}((f)+I)=n+1$ and $\operatorname{ht}I=n$, where $I=\ker^{(1)}\varphi$. A divisor is always of linear Jacobian type at smooth points, so the quotient $\ker\varphi/\ker^{(1)}\varphi$ is supported at the singular locus. In particular, for any $a\in\ker\varphi$, there exists $k\geq 1$ such that $f^{k}a\in I$. Thus, it remains to prove that $f$ is not a zero divisor modulo $I$. Suppose it were. Then $f$ would lie in some associated prime $\mathfrak{p}$ of $I$. In particular, $(f)+I\subset\mathfrak{p}$. But then $\operatorname{ht}\mathfrak{p}\geq\operatorname{ht}((f)+I)=n+1$, which is a contradiction with the fact that $I$ is unmixed.
Remark 4.17.
  • (a)
    The unmixedness theorem, which holds in any Cohen-Macaulay ring, states that any ideal of height $h$ that can be generated by $h$ elements is unmixed [MatsumuraRosa, 16.C]. Thus, Proposition 4.16 applies when $\ker^{(1)}\varphi$ can be generated by $n$ elements. This is what happens in the free case.
    This result suggests the introduction of stronger versions of the Saito-holonomic type properties, in which we also require that the ideal in question is unmixed. Those notions would coincide with the previous ones for free divisors, for which the unmixedness hypothesis is automatic. However, we have not explored the relations between them.
  • (b)
    Proposition 4.16 complements [Uli17, Corollary 3.23] where, instead of unmixedness, tameness is assumed.
Let us revisit Examples 3.11 and 3.12 to study their Saito-holonomic type properties:
Example 4.18.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z+yt)$. We have that
[…]
As $\dim D_{0}=1$, $D$ is not Saito-holonomic. Consequently, it is not strongly Saito-holonomic. However, $\dim\tilde{D}_{i}\leq i$ for $i=0,1$, so it is weakly Saito-holonomic.
Recall that $D$ is strongly Euler-homogeneous. This example tells us that we need to demand Saito-holonomicity in Theorem 4.12. Weak Saito-holonomicity is not enough to guarantee the equivalence between strong Euler-homogeneity and strong Saito-holonomicity.
Example 4.19.
Let $D=V(f)$ in $(\mathbb{C}^{4},0)$ with $f=xy(x+z)(x^{2}+yz)(z^{2}+yt)$. We have that:
$\begin{array}[]{l}\tilde{D}_{0}=\varnothing,\\[5.69046pt] \tilde{D}_{1}=V(x,y,z),\hskip 8.5359ptD_{0}=(x,y,z,t),\\[5.69046pt] \tilde{D}_{2}=D_{1}=V(t(y+z),t(x+z),z(z-t),z(y+t),z(x+t),xy-zt).\end{array}$
As $\dim D_{i}=i$ for $i=0,1$, $D$ is Saito-holonomic (and, therefore, also weakly Saito-holonomic). Since $D$ is not strongly Euler-homogeneous, Theorem 4.12 tells us that $D$ is not strongly Saito-holonomic either. And, indeed, we confirm this by noticing that $\dim\tilde{D}_{1}=1>0$.
5 Extending the criteria to formal power series
The objective of this section is to extend the characterizations developed in previous sections, in which the Saito matrices play an important role, to the case of formal power series. This will be useful to prove the main results of the next section.
[…]
Consider the formal Jacobian ideal $\widehat{\mathcal{J}}_{f}=\widehat{\mathcal{O}}(\partial_{1}(f),\ldots,\partial_{n}(f),f)$. Let $J_{i}=\operatorname{Fitt}_{n-i}(\widehat{\mathcal{J}}_{f}/(f))$. As in the convergent case, a free presentation of $\widehat{\mathcal{J}}_{f}/(f)$ as an $\widehat{\mathcal{O}}-$module is given by the matrix $B$, so $J_{i}$ is the ideal of $\widehat{\mathcal{O}}$ generated by the minors of order $i$ of $B$ (for $1\leq i\leq n$). Similarly, if we set $\tilde{J}_{i}=\operatorname{Fitt}_{n+1-i}(\widehat{\mathcal{J}}_{f})$, then $\tilde{J}_{i}$ is the ideal of $\widehat{\mathcal{O}}$ generated by the minors of order $i$ of $\tilde{B}$.
Once more, these ideals do not depend on the choice of basis, coordinates or equation. Note that any convergent generating set $\mathcal{S}$ of $\operatorname{Der}_{f}$ is still an $\widehat{\mathcal{O}}-$generating set of $\widehat{\operatorname{Der}}_{f}$, so we deduce that $J_{i}$ and $\tilde{J}_{i}$ are also the ideals of $\widehat{\mathcal{O}}$ generated by the minors of order $i$ of the Saito matrices with respect to $\mathcal{S}$ and $f$. That is, $J_{i}=I_{i}^{e}$ and $\tilde{J}_{i}=\tilde{I}_{i}^{e}$.
In particular, by $(3)$ in Proposition 3.4, we have that $\sqrt{J_{n-1}}=\sqrt{\tilde{J}_{n}}=\sqrt{\widehat{\mathcal{J}}_{f}}$.
Now, we can state the algebraic criterion for strong Euler-homogeneity in terms of $J_{i}$ and $\tilde{J}_{i}$:
Proposition 5.1 (Formal criterion for strong Euler-homogeneity).
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$ with $n\geq 2$. Then:
  • (1)
    $D$ is strongly Euler-homogeneous on $D_{0}$ if and only if $\tilde{J}_{1}=\widehat{\mathcal{O}}$.
  • (2)
    $D$ is strongly Euler-homogeneous on $D\setminus D_{0}$ if and only if $\sqrt{\tilde{J}_{i+1}}=\sqrt{J_{i}}$ for $i=1,\ldots,n-2$.
  • (3)
    $D$ is strongly Euler-homogeneous if and only if $\tilde{J}_{1}=\widehat{\mathcal{O}}$ and $\sqrt{\tilde{J}_{i+1}}=\sqrt{J_{i}}$ for $i=1,\ldots,n-2$.
Proof.
Due to faithful flatness of $\widehat{\mathcal{O}}$ over $\mathcal{O}$, given two ideals $I,J\subset\mathcal{O}$, we have that $\sqrt{I}=\sqrt{J}$ if and only if $\sqrt{I^{e}}=\sqrt{J^{e}}$ in $\widehat{\mathcal{O}}$. Thus, $\sqrt{\tilde{I}_{i+1}}=\sqrt{I_{i}}$ if and only if $\sqrt{\tilde{J}_{i+1}}=\sqrt{J_{i}}$. The result follows from Corollary 3.9.
Remark 5.2.
As before, this criterion can be read in terms of Fitting ideals. We just need to replace $\mathcal{J}_{f}$ by $\widehat{\mathcal{J}}_{f}$ and $\mathcal{J}_{f}/(f)$ by $\widehat{\mathcal{J}}_{f}/(f)$ in Remark 3.10.
The relevance of Proposition 5.1 is that, unexpectedly, formal power series, which only make sense at the origin, can be used to study strong Euler-homogeneity away from this point. Furthermore, the necessary condition for having strong Euler-homogeneity outside $D_{0}$ (Corollary 3.14) also has a formal analogue, although the proof is not so straightforward:
Corollary 5.3.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$ that is strongly Euler-homogeneous outside $D_{0}$. Let $g$ be a “formal equation” of $D$ (i.e. a formal unit multiplied by $f$). Then $\sqrt{\tilde{J}_{1}}=\sqrt{\partial g:g}$.
Proof.
Let $\{\hat{\eta}_{1},\ldots,\hat{\eta}_{m}\}$ be a generating set of $\widehat{\operatorname{Der}}_{f}$ such that $\hat{\eta}_{i}(g)=\hat{\beta}_{i}g$ for all $i=1,\ldots,m$, so that $\partial g:g=(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})$. Recall that the $J_{i}$ (resp. $\tilde{J}_{i}$) do not depend on the chosen presentation, so these ideals are the ones generated by the minors of order $i$ of the Saito matrix with respect to $\{\hat{\eta}_{1},\ldots,\hat{\eta}_{m}\}$ (resp. the extended Saito matrix with respect to $\{\hat{\eta}_{1},\ldots,\hat{\eta}_{m}\}$ and $g$). Thus, $(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})\subset\tilde{J}_{1}$ and it is clear that $\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}\subset\sqrt{\tilde{J}_{1}}$.
Let us prove the other inclusion. The proof starts as in Lemma 3.13: $g$ is integral over the ideal of its partial derivatives (convergent series with the trivial metric are exactly formal power series, so the proof of [Scheja70, Satz 5.2] is still valid), so either some $\hat{\beta}_{i}$ is a unit or a power of $g$ belongs to $(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})$.
In the first case, it is clear that $(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})=\tilde{J}_{1}=\widehat{\mathcal{O}}$ and we are done. In the second one, as $D_{n-1}=D$ and $f$ is reduced, $\sqrt{I_{n}}=(f)$ and $\sqrt{J_{n}}=\sqrt{I_{n}^{e}}=(f)^{e}=(g)\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$.
Note that, reasoning as in $(1)$ in Proposition 3.4, we have $\tilde{J}_{i}\subset J_{i}+(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})$, so $\sqrt{\tilde{J}_{i}}\subset\sqrt{J_{i}+(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}=\sqrt{\sqrt{J_{i}}+\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}}$ for all $i$. In particular, as $\sqrt{J_{n}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$, we also have $\sqrt{\tilde{J}_{n}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$. But, as $D$ is strongly Euler-homogeneous outside $D_{0}$, we have that $\sqrt{J_{i}}=\sqrt{\tilde{J}_{i+1}}$ for $i\leq n-1$ (Proposition 5.1). This implies $\sqrt{J_{n-1}}=\sqrt{\tilde{J}_{n}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$ and so
$\sqrt{\tilde{J}_{n-1}}\subset\sqrt{\sqrt{J_{n-1}}+\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}.$
By reverse induction, we finally get $\sqrt{\tilde{J}_{1}}\subset\sqrt{(\hat{\beta}_{1},\ldots,\hat{\beta}_{m})}$.
Remark 5.4.
As in the convergent case, we deduce that, when $D$ is strongly Euler-homogeneous outside $D_{0}$, the radical of the ideal $\partial g:g$ does not depend on the choice of formal equation $g$.
We have argued that the Fitting ideals of the $\widehat{\mathcal{O}}-$modules $\widehat{\mathcal{J}}_{f}$ and $\widehat{\mathcal{J}}_{f}/(f)$ are the extended ones of those of $\mathcal{J}_{f}$ and $\mathcal{J}_{f}/(f)$. Since the height of the extension is the same, we deduce from Corollary 4.11 that we can also check (weak or strong) Saito-holonomicity using formal data:
Corollary 5.5.
Let $D=V(f)$ be a germ of divisor in $(\mathbb{C}^{n},0)$. Then:
  • (1)
    $D$ is weakly Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\widehat{\mathcal{J}}_{f})\geq i$ for all $i=3,\ldots,n$.
  • (2)
    $D$ is Saito-holonomic if and only if $\hskip 2.84544pt\operatorname{ht}\operatorname{Fitt}_{i}(\widehat{\mathcal{J}}_{f}/(f))\geq i+1$ for all $i=2,\ldots,n-1$.
  • (3)
    $D$ is strongly Saito-holonomic if and only if $\hskip 2.84544pt(\tilde{J}_{1}=)\operatorname{Fitt}_{n}(\widehat{\mathcal{J}}_{f})=\widehat{\mathcal{O}}$ and $\operatorname{ht}\operatorname{Fitt}_{i}(\widehat{\mathcal{J}}_{f})\geq i+1$ for all $i=2,\ldots,n-1$.
6 Proving new cases of Conjecture 1.4
6.1 Assuming strong Euler-homogeneity in a punctured neighbourhood
In general, there is no hope that strong Euler-homogeneity outside a point implies that the point also has this property. In order to give a counterexample, note that a divisor is always strongly Euler-homogeneous at smooth points (because a local equation is $x_{1}=0$, which verifies $x_{1}\partial_{1}(x_{1})=x_{1}$). So we just need to consider an isolated singularity that is not quasihomogeneous (and so not strongly Euler-homogeneous by [Saito71, Satz 4.1]) at the origin. Take, for example, $V(x^{4}+y^{5}+y^{4}x)\subset\mathbb{C}^{2}$. However, we will see that, for free divisors, under mild assumptions, this becomes true.
Theorem 6.1.
Let $D=V(f)$ be a germ of free divisor in $(\mathbb{C}^{n},0)$ that is strongly Euler-homogeneous outside the origin. Then, the following are equivalent:
  • (1)
    $D$ is strongly Euler-homogeneous at the origin.
  • (2)
    There exists a non-topologically nilpotent singular derivation in $\widehat{\operatorname{Der}}_{f}$.
  • (3)
    There exists a non-zero semisimple derivation in $\widehat{\operatorname{Der}}_{f}$.
Proof.
$(1)\Rightarrow(2)$: If $D$ is strongly Euler-homogeneous at 0, then there exists a singular derivation $\delta\in\widehat{\operatorname{Der}}_{f}$ such that $\delta(f)=f$. By statement $(2)$ in Proposition 2.10, $\delta$ cannot be topologically nilpotent and we are done.
$(2)\Rightarrow(3)$: If there exists a non-topologically nilpotent singular derivation $\delta\in\widehat{\operatorname{Der}}_{f}$, in particular, its semisimple part $\delta_{S}$ is not zero and it is still logarithmic (Proposition 2.10, $(3)$).
$(3)\Rightarrow(1)$: Suppose $D$ is not strongly Euler-homogeneous at 0. In particular, it cannot be a product at 0 (otherwise, it would be non-strongly Euler homogeneous on a set of positive dimension), so $\widehat{\operatorname{Der}}_{f}\subset\mathfrak{m}\widehat{\operatorname{Der}}$ by Proposition 2.7. On the other hand, by Proposition 3.6, the origin must be the only point $p$ at which $\operatorname{rank}A(p)=\operatorname{rank}\tilde{A}(p)=0$. That is, $\tilde{D}_{0}=\{0\}$ and $\mathfrak{m}=\sqrt{\tilde{I}_{1}}$. Taking extensions, we have $\widehat{\mathfrak{m}}=\sqrt{\tilde{J}_{1}}$.
If there exists a non-zero semisimple derivation $\delta\in\widehat{\operatorname{Der}}_{f}$, then its class in $\widehat{\operatorname{Der}}_{f}/\mathfrak{m}\widehat{\operatorname{Der}}_{f}$ is non-zero. Indeed, if $\delta\in\mathfrak{m}\widehat{\operatorname{Der}}_{f}\subset\mathfrak{m}^{2}\widehat{\operatorname{Der}}$, then $\delta$, not having linear part, would also be topologically nilpotent by $(1)$ in Proposition 2.10, something that can only happen if $\delta=0$ by uniqueness.
Thus, by Nakayama’s Lemma, we can construct a formal basis $\{\delta_{1},\ldots,\delta_{n}\}$ of $\widehat{\operatorname{Der}}_{f}$ with $\delta_{1}=\delta$. Now, we can find a formal unit $u$ such that, for $g=uf$, we have $\delta_{1}(g)=\beta g$ with $\beta\in\mathbb{C}$ (Proposition 2.10, $(4)$). Since we are assuming $D$ is not strongly Euler-homogeneous at the origin, $\beta=0$ (otherwise, $\tilde{J}_{1}=\widehat{\mathcal{O}}$ and $D$ would be strongly Euler-homogeneous at 0 by Proposition 5.1). Let $\beta_{i}\in\widehat{\mathcal{O}}$ be such that $\delta_{i}(g)=\beta_{i}g$ (note that $\beta_{1}=\beta=0$). By Corollary 5.3, we have $\widehat{\mathfrak{m}}=\sqrt{\tilde{J}_{1}}=$ $\sqrt{(\beta_{1},\ldots,\beta_{n})}=\sqrt{(\beta_{2},\ldots,\beta_{n})}$, but this is impossible since $\operatorname{ht}\widehat{\mathfrak{m}}=n$ and $\operatorname{ht}\sqrt{(\beta_{2},\ldots,\beta_{n})}=\operatorname{ht}(\beta_{2},\ldots,\beta_{n})\leq n-1$.
Remark 6.2.
  • (a)
    In particular, if $D$ is a non-quasihomogeneous germ of plane curve (which is an isolated singularity), then all its logarithmic derivations are topologically nilpotent. This generalizes [dVR24, Proposition 5.2], where we proved that, for plane curves $D=V(f)$ with $f\in\mathfrak{m}^{3}$, non-Euler logarithmic derivations (those $\delta$ for which $\delta(f)\in\widehat{\mathfrak{m}}f$) are topologically nilpotent.
  • (b)
    As we said before, $D=V(f)$ with $f=x^{4}+y^{5}+y^{4}x$ is a non-quasihomogeneous plane curve. A basis of $\operatorname{Der}_{f}$ is:
[…]
    And, indeed, as they have no linear part, every logarithmic derivation is topologically nilpotent.
A direct consequence of Theorem 6.1 is that a weak version of Conjecture 1.4 holds in arbitrary dimension:
Theorem 6.3.
If $D=V(f)$ is a germ of free divisor in $(\mathbb{C}^{n},0)$ satisfying LCT and being strongly Euler-homogeneous outside 0, then it is so at 0.
Proof.
Suppose $D$ is not strongly Euler-homogeneous at the origin. By Theorem 6.1, all singular logarithmic derivations are topologically nilpotent. In particular, all of them have zero trace so, by Theorem 2.11, LCT cannot hold.
Remark 6.4.
This gives an alternative proof of the conjecture for $n=2$, since singularities are isolated and a divisor is always strongly Euler-homogeneous at smooth points.
6.2 Assuming weak Koszul-freeness
The characterization given in Theorem 4.6 allows us to prove that the conjecture holds for weakly Koszul-free divisors:
Theorem 6.5.
Let $D\subset X$ be a weakly Koszul-free divisor. If $D$ satisfies LCT, then it is strongly Euler-homogeneous.
Proof.
Since these are local properties, we may assume $D$ is a germ of free divisor in $(\mathbb{C}^{n},0)$. The statement is trivially true for $n=1$ (we also know that it is true for $n=2$, but the following argument gives a new proof of this fact). Suppose it is true for dimension $n$ and let us prove it for dimension $n+1$.
If $D$ is a product at a point, then $D$ is locally isomorphic to $D^{\prime}\times\mathbb{C}$, where $D^{\prime}$ is a germ of weakly Koszul-free divisor in $\mathbb{C}^{n}$ that satisfies LCT by Proposition 4.9 and Lemma 2.4. By induction hypothesis, $D^{\prime}$ is strongly Euler-homogeneous and so is $D$. This proves that $D$ is strongly Euler-homogeneous outside $D_{0}$ (where $D$ is a product). As $D$ is weakly Koszul-free, by Theorem 4.6, $\dim\tilde{D}_{0}\leq 0$, so $\tilde{D}_{0}\subset\{0\}$. By Proposition 3.6, $D$ is always strongly Euler-homogeneous on $D_{0}\setminus\tilde{D}_{0}$, so we have strong Euler-homogeneity on a punctured neighbourhood of 0. Since $D$ satisfies LCT, by Theorem 6.3, $D$ must also be strongly Euler-homogeneous at 0.
It is well-known that, even for free divisors, LCT does not imply Koszul-freeness. The four lines is a classical counterexample. However, this divisor is still weakly Koszul-free. From Theorem 4.6 we can easily deduce an interesting consequence: this is a general fact for free divisors in ambient dimension $n=3$.
Corollary 6.6.
Let $D\subset X$ be a free divisor in a complex analytic manifold of dimension $3$. If $D$ satisfies LCT, then it is weakly Koszul-free.
Proof.
These properties being local, we may assume $D$ is a germ of free divisor in $(\mathbb{C}^{3},0)$. By Theorem 4.6, $D$ is weakly Koszul-free if and only if $\dim\tilde{D}_{0}\leq 0$. As $D$ satisfies LCT, it is strongly Euler-homogeneous (because we know Conjecture 1.4 is true for $n=3$). In particular, $\tilde{D}_{0}=\varnothing$ and we are done.
6.3 Assuming ambient dimension $n=4$
M. Granger and M. Schulze proved in [GS, Theorem 5.4] the so-called formal structure theorem for logarithmic vector fields. Given a germ $f\in\widehat{\mathcal{O}}$, they call $s$ the maximal dimension of the vector space of diagonal derivations $\sigma\in\operatorname{Der}_{f}$ such that $\sigma(f)\in\mathbb{C}\cdot f$ for $f$ varying in a formal contact equivalence class (that is, $s$ is maximal for all formal coordinate systems and changes of $f$ by unit factors). Then, they prove that there exists a minimal generating set of $\operatorname{Der}_{f}$ in which $s$ derivations are diagonal in some formal coordinate system and the rest of them are (topologically) nilpotent satisfying some additional properties.
They use the notion of semisimplicity given by K. Saito (for them, a derivation is semisimple if it only has linear part and the associated matrix is semisimple), which depends on the choice of coordinates. In this subsection, we first develop a version of this theorem in which we use the concepts of semisimplicity and topological nilpotency from Definition 2.8, following the Gérard-Levelt approach. This allows us to define $s$ intrinsically as the maximal number of commuting semisimple derivations in a minimal generating set of $\operatorname{Der}_{f}$, in such a way that it is completely independent of the choice of equation or coordinate system. Then, we use this version of the formal structure theorem to show that Conjecture 1.4 holds unconditionally in the $4$-dimensional case.
First, we need to generalize the statement $(4)$ in Proposition 2.10 to the case of multiple diagonal derivations:
Proposition 6.7.
Let $f\in\mathcal{O}$, $f\neq 0$, and let $\delta_{1},\ldots,\delta_{m}\in\mathfrak{m}\widehat{\operatorname{Der}}$ be diagonal derivations (in some formal coordinate system) such that $\delta_{i}(f)=\alpha_{i}f$ for some $\alpha_{i}\in\widehat{\mathcal{O}}$. Then, there exists a formal unit $u$ such that, for $g=uf$, we have $\delta_{i}(g)=\alpha_{i0}g$, where $\alpha_{i0}$ is the constant term of $\alpha_{i}$.
Proof.
Let us write $\delta_{i}=\underline{x}D_{i}\overline{\partial}$, where $D_{i}=\operatorname{diag}(\lambda_{i})$ and $\lambda_{i}$ is a vector of (possibly complex) weights. Also write $\alpha_{i}=\sum_{\beta\in\mathbb{N}^{n}}\alpha_{i\beta}{x}^{\beta}$.
Now, set
$u=\exp\left(-\sum_{{\beta\cdot\lambda_{1}}\neq 0}\dfrac{\alpha_{1\beta}x^{\beta}}{\beta\cdot\lambda_{1}}-\ldots-\sum_{\begin{subarray}{c}\beta\cdot\lambda_{1}=0\\ \cdots\\ \beta\cdot\lambda_{m}\neq 0\end{subarray}}\dfrac{\alpha_{m\beta}x^{\beta}}{\beta\cdot\lambda_{m}}\right).$
Note that $u$ is a well-defined formal power series because the argument of the exponential, call it $b$, vanishes at the origin ($\beta=0$ is not in the sum), so we can compose the two series and the result is another formal power series that does not vanish at the origin. Thus, $u$ is a formal unit and
$\delta_{i}(u)=\delta_{i}(e^{b})=e^{b}\delta_{i}(b)=u\delta_{i}(b).$
In order to compute $\delta_{i}(b)$, take into account that $\delta_{i}$ applied to a monomial $x^{\beta}$ returns $(\beta\cdot\lambda_{i})x^{\beta}$. Moreover, diagonal derivations always commute, so $0=[\delta_{i},\delta_{j}](f)=(\delta_{i}(\alpha_{j})-\delta_{j}(\alpha_{i}))f$ and we deduce that $\delta_{i}(\alpha_{j})=\delta_{j}(\alpha_{i})$. Therefore:
$\delta_{i}(\alpha_{j})=\sum_{\beta}(\beta\cdot\lambda_{i})\alpha_{j\beta}x^{\beta}=\sum_{\beta}(\beta\cdot\lambda_{j})\alpha_{i\beta}x^{\beta}=\delta_{j}(\alpha_{i}),$
that is, $(\beta\cdot\lambda_{i})\alpha_{j\beta}=(\beta\cdot\lambda_{j})\alpha_{i\beta}$ for all $i,j=1,\ldots,m$ and $\beta\in\mathbb{N}^{n}$. Thus,
$\delta_{i}(b)$ $=-\sum_{{\beta\cdot\lambda_{1}}\neq 0}\dfrac{\alpha_{1\beta}(\beta\cdot\lambda_{i})}{\beta\cdot\lambda_{1}}x^{\beta}-\ldots-\sum_{\begin{subarray}{c}\beta\cdot\lambda_{1}=0\\ \cdots\\ \beta\cdot\lambda_{i}\neq 0\end{subarray}}\dfrac{\alpha_{i\beta}(\beta\cdot\lambda_{i})}{\beta\cdot\lambda_{i}}x^{\beta}$
$=-\sum_{{\beta\cdot\lambda_{1}}\neq 0}\alpha_{i\beta}x^{\beta}-\ldots-\sum_{\begin{subarray}{c}\beta\cdot\lambda_{1}=0\\ \cdots\\ \beta\cdot\lambda_{i}\neq 0\end{subarray}}\alpha_{i\beta}x^{\beta}.$
Let $g=uf$, so that
$\delta_{i}(g)$ $=\delta_{i}(u)f+u\delta_{i}(f)=u\delta_{i}(b)f+u\alpha_{i}f=(\alpha_{i}+\delta_{i}(b))uf=(\alpha_{i}+\delta_{i}(b))g$
$=\left(\sum_{\beta}\alpha_{i\beta}x^{\beta}-\sum_{{\beta\cdot\lambda_{1}}\neq 0}\alpha_{i\beta}x^{\beta}-\ldots-\sum_{\begin{subarray}{c}\beta\cdot\lambda_{1}=0\\ \cdots\\ \beta\cdot\lambda_{i}\neq 0\end{subarray}}\alpha_{i\beta}x^{\beta}\right)g=\left(\sum_{\begin{subarray}{c}\beta\cdot\lambda_{1}=0\\ \cdots\\ \beta\cdot\lambda_{i}=0\end{subarray}}\alpha_{i\beta}x^{\beta}\right)g.$
Let us see that $\alpha^{\prime}_{i}:=\displaystyle\sum_{\begin{subarray}{c}\beta\cdot\lambda_{1}=0\\ \cdots\\ \beta\cdot\lambda_{i}=0\end{subarray}}\alpha_{i\beta}x^{\beta}$ must be $\alpha_{i0}$ and we will have the result:
Note that $\delta_{i}(\alpha^{\prime}_{i})=0$ (as each summand $x^{\beta}$ of $\alpha^{\prime}_{i}$ satisfies $\beta\cdot\lambda_{i}=0$), so if we decompose $g$ as a sum of eigenvectors of $\delta_{i}$ (which can be done in a unique way by [Saito71, Lemma 2.3]), $g=\sum_{\mu\in\mathbb{C}}g_{\mu}$ with $\delta_{i}(g_{\mu})=\mu g_{\mu}$, we have
$\delta_{i}(g)=\sum_{\mu\in\mathbb{C}}\mu g_{\mu}=\alpha^{\prime}_{i}g=\sum_{\mu\in\mathbb{C}}\alpha^{\prime}_{i}g_{\mu}.$
Since $\mu g_{\mu}$ and $\alpha^{\prime}_{i}g_{\mu}$ are both eigenvectors of $\delta_{i}$ for $\mu$, by the uniqueness of the decomposition, we must have $\mu g_{\mu}=\alpha^{\prime}_{i}g_{\mu}$ for every $\mu\in\mathbb{C}$. As $g\neq 0$, some $g_{\mu}\neq 0$. But then $\alpha^{\prime}_{i}=\mu$ is constant and the only possibility for this is $\alpha^{\prime}_{i}=\alpha_{i0}$, as that is the value of $\alpha^{\prime}_{i}$ at the origin.
Theorem 6.8.
Let $f\in\widehat{\mathcal{O}}$ be such that $\operatorname{Der}_{f}\subset\mathfrak{m}\widehat{\operatorname{Der}}$ and let $s\geq 0$ be the maximal number of commuting semisimple derivations in a minimal generating set of $\operatorname{Der}_{f}$. If $s=0$, then each minimal generating set of $\operatorname{Der}_{f}$ is formed by topologically nilpotent derivations. Otherwise, there exist a minimal generating set $\{\sigma_{1},\ldots,\sigma_{s},\nu_{1},\ldots,\nu_{r}\}$ of $\operatorname{Der}_{f}$, a formal coordinate system $(y_{1},\ldots,y_{n})$ and a formal unit $u$ such that
  • (1)
    $\sigma_{1},\ldots,\sigma_{s}$ are diagonal with rational weights: $\sigma_{i}=\sum_{j=1}^{n}w_{ij}y_{j}\dfrac{\partial}{\partial y_{j}}$ with $w_{ij}\in\mathbb{Q}$ for all $i=1,\ldots,s$ and $j=1,\ldots,n$.
  • (2)
    For $g=uf$ we have $\sigma_{i}(g)=\alpha_{i}g$ with $\alpha_{i}\in\mathbb{Q}$ for all $i=1,\ldots,s$.
  • (3)
    $\nu_{1},\ldots,\nu_{r}$ are topologically nilpotent.
  • (4)
    $\sigma_{i},\nu_{j}]=c_{ij}\nu_{j}$ with $c_{ij}\in\mathbb{Q}$ (that is, $\nu_{j}$ is $\sigma_{i}$-homogeneous of weight $c_{ij}$) for all $i=1,\ldots,s$, $j=1,\ldots,r$.
Proof.
Suppose $s=0$ and let $\{\delta_{1},\ldots,\delta_{k}\}$ be a minimal generating set of $\operatorname{Der}_{f}$. Their semisimple and topologically nilpotent parts are still logarithmic (Proposition 2.10), so $\{\delta_{1S},\delta_{1N},\ldots,\delta_{kS},\delta_{kN}\}$ is a generating set. We claim that $\delta_{iS}=0$ for all $i$: the class of each $\delta_{iS}$ modulo $\widehat{\mathfrak{m}}\operatorname{Der}_{f}$ must be zero (otherwise, it could be extended to a basis of $\operatorname{Der}_{f}/\widehat{\mathfrak{m}}\operatorname{Der}_{f}$, giving, by Nakayama’s Lemma, a minimal generating set of $\operatorname{Der}_{f}$ with a semisimple derivation and $s\geq 1$). Thus, $\delta_{iS}\subset\widehat{\mathfrak{m}}\operatorname{Der}_{f}\subset\mathfrak{m}^{2}\widehat{\operatorname{Der}}$. But any derivation in $\mathfrak{m}^{2}\widehat{\operatorname{Der}}$ is topologically nilpotent (as it has no linear part) so $\delta_{iS}$ is both semisimple and topologically nilpotent. By the uniqueness of the decomposition, we must have $\delta_{iS}=0$. This means $\delta_{i}=\delta_{iN}$ is topologically nilpotent for all $i$ and we are done.
Now, if $s\geq 1$, consider a minimal generating set $\{\sigma_{1},\ldots,\sigma_{s},\eta_{1},\ldots,\eta_{r}\}$ of $\operatorname{Der}_{f}$, where $\sigma_{1},\ldots,\sigma_{s}$ are commuting semisimple derivations. With a bit of linear algebra, [GL, Théorème 2.3] can be easily generalized to the case of $s$ commuting derivations, so we deduce that there exists a regular system of parameters $y_{1},\ldots,y_{n}$ such that $\sigma_{i}=\sum_{j=1}^{n}w_{ij}y_{j}\partial_{j}$ with $w_{ij}\in\mathbb{C}$ for all $i=1,\ldots,s$, $j=1,\ldots,n$. By Proposition 6.7, there exists a formal unit $u$ such that $g=uf$ verifies $\sigma_{i}(g)=\alpha_{i}g$ with $\alpha_{i}\in\mathbb{C}$ for all $i=1,\ldots,s$. Also, we may suppose that $w_{ij}$ and $\alpha_{i}$ are rational for all $i,j$ by [Saito71, Lemma 1.4].
Write $\underline{\sigma}=(\sigma_{1},\ldots,\sigma_{s})$. Now, we decompose $\eta_{i}=\sum_{\underline{\lambda}\in\mathbb{Q}^{s}}\eta_{i\underline{\lambda}}$, where $\eta_{i\underline{\lambda}}$ is $\underline{\sigma}$-homogeneous of degree $\underline{\lambda}$ for all $i=1,\ldots,r$. Let $\beta\in\widehat{\mathcal{O}}$ be such that $\eta_{i}(g)=\beta g$ (recall that $\eta_{i}$ is also logarithmic for $g$) and write in a similar way $\beta=\sum_{\underline{\lambda}\in\mathbb{Q}^{s}}\beta_{\underline{\lambda}}$. We get
$\sum_{\underline{\lambda}\in\mathbb{Q}^{s}}\eta_{i\underline{\lambda}}(g)=\sum_{\underline{\lambda}\in\mathbb{Q}^{s}}\beta_{\underline{\lambda}}g,$
where $\eta_{i\underline{\lambda}}(g)$ and $\beta_{\underline{\lambda}}g$ are both $\underline{\sigma}$-homogeneous of the same degree ($\underline{\lambda}+\underline{\alpha}$, where we are writing $\underline{\alpha}=(\alpha_{1},\ldots,\alpha_{s})$). By the uniqueness of this decomposition, we get $\eta_{i\underline{\lambda}}(g)=\beta_{\underline{\lambda}}g$, so $\eta_{i\underline{\lambda}}$ is logarithmic for all $i=1,\ldots,r$ and $\underline{\lambda}\in\mathbb{Q}^{s}$.
Note that each derivation $\sigma_{i}\in\mathfrak{m}\widehat{\operatorname{Der}}$ induces a linear map $\bar{\sigma_{i}}$ in $\operatorname{Der}_{f}/\widehat{\mathfrak{m}}\operatorname{Der}_{f}$ by $\bar{\sigma_{i}}(\bar{\eta})=\overline{[\sigma_{i},\eta]}$ (it is well-defined since $\eta\in\widehat{\mathfrak{m}}\operatorname{Der}_{f}$ implies $\sigma_{i},\eta]\in\widehat{\mathfrak{m}}\operatorname{Der}_{f}$). We claim that, for each $\delta\in\operatorname{Der}_{f}$, the number of $\delta_{\underline{\lambda}}$ whose class in $\operatorname{Der}_{f}/\widehat{\mathfrak{m}}\operatorname{Der}_{f}$ is not zero is finite: suppose the class of $\delta_{\underline{\lambda}}$ in $\operatorname{Der}_{f}/\widehat{\mathfrak{m}}\operatorname{Der}_{f}$ is not zero. Then, it is a common eigenvector of the induced maps $\overline{\sigma_{i}}$ with $\underline{\lambda}$ as vector of eigenvalues. But since $\operatorname{Der}_{f}/\widehat{\mathfrak{m}}\operatorname{Der}_{f}$ is of finite dimension $s+r$, the number of different $s$-tuples of eigenvalues is also finite.
In particular, the class of each $\eta_{i}$ is a (finite) linear combination of the classes of the $\eta_{i\underline{\lambda}}$, so these classes together with those of the $\sigma_{i}$ (a finite number) generate $\operatorname{Der}_{f}/\widehat{\mathfrak{m}}\operatorname{Der}_{f}$. By Nakayama’s Lemma, the corresponding derivations generate $\operatorname{Der}_{f}$. Thus, we get that a finite subset $S$ of $\{\sigma_{1},\ldots,\sigma_{s},\eta_{1\underline{\lambda}},\ldots,\eta_{r\underline{\lambda}},\mid\underline{\lambda}\in\mathbb{Q}^{s}\}$ containing $\sigma_{1},\ldots,\sigma_{s}$ is a generating set of $\operatorname{Der}_{f}$.
Now, for each $\eta\in S\setminus\{\sigma_{1},\ldots,\sigma_{s}\}$ we do the following: if its $\underline{\sigma}$-degree is $\underline{\lambda}\neq\underline{0}$, then it is topologically nilpotent [GS, Lemma 2.6] and we do nothing. If $\underline{\lambda}=\underline{0}$, then we replace $\eta$ by $\eta_{S},\eta_{N}$ (which still commute with the $\sigma_{i}$ by [GL, Théorème 1.6]) and we still have a generating set of $\operatorname{Der}_{f}$. Denote again by $S$ the generating set obtained from this process (which is formed exclusively by $\underline{\sigma}$-homogeneous topologically nilpotent and commuting semisimple derivations) and extract a minimal generating set $S^{\prime}$ from $S$ containing $\sigma_{1},\ldots,\sigma_{s}$. Since $s$ is maximal, the only semisimple derivations in $S^{\prime}$ must be $\sigma_{1},\ldots,\sigma_{s}$. Thus, the rest of them, $\nu_{1},\ldots,\nu_{r}$, must be topologically nilpotent and we get the desired result.
Remark 6.9.
  • (a)
    If $f\in\mathcal{O}$, Theorem 6.8 is valid for $\widehat{\operatorname{Der}}_{f}$.
  • (b)
    As previously mentioned, this version of the formal structure theorem gives an intrinsic definition of the number $s$.
We are now ready to prove Conjecture 1.4 for $n=4$, following a similar strategy to that used in the $n=3$ case in [GS]:
Theorem 6.10.
Let $X$ be a $4$-dimensional analytic manifold and let $D\subset X$ be a free divisor in it. If $D$ satisfies LCT, then $D$ must be strongly Euler-homogeneous.
Proof.
As these properties are local, we may assume $D=V(f)$ is a germ of free divisor in $(\mathbb{C}^{4},0)$ and it is enough to prove that if it satisfies LCT, then it is strongly Euler-homogeneous at 0.
Note that, since $D$ is a product outside $D_{0}$, by Lemma 2.4, it must be strongly Euler-homogeneous at those points (as we know LCT implies strong Euler-homogeneity for free divisors in dimension $3$). Let us suppose that it is not strongly Euler-homogeneous at 0. In particular, it cannot be a product at 0.
As LCT holds for $D$, by Theorem 2.11, there must be at least one logarithmic derivation with non-zero trace (in particular, it cannot be topologically nilpotent). Thus, in the notation of Theorem 6.8, $s\geq 1$ and we can find a formal coordinate system $x_{1},\ldots,x_{4}$ and a basis of $\widehat{\operatorname{Der}}_{f}$, $\{\sigma,\delta_{2},\delta_{3},\delta_{4}\}$, such that $\sigma=\sum_{i=1}^{4}\lambda_{i}x_{i}\partial_{i}$ with $\lambda_{i}\in\mathbb{Q}$, $\operatorname{tr}(\sigma)=\sum_{i=1}^{4}\lambda_{i}\neq 0$ and $\delta_{i}$ is $\sigma$-homogeneous of degree $c_{i}\in\mathbb{Q}$ for $i=2,3,4$ (if $s>1$ so that some of the $\delta_{i}$ are diagonal, then they trivially commute with $\sigma$ and $c_{i}=0$). Moreover, there exists a formal unit $u$ such that $g=uf$ verifies $\sigma(g)=\alpha_{1}g$, with $\alpha_{1}\in\mathbb{C}$. Let $\alpha_{i}\in\widehat{\mathcal{O}}$ be such that $\delta_{i}(g)=\alpha_{i}g$ for $i=2,3,4$. As we are assuming $D$ is not strongly Euler-homogeneous at 0, by Proposition 2.7, $\alpha_{i}\in\widehat{\mathfrak{m}}$ for all $i=1,\ldots,4$. In particular, $\alpha_{1}=0$.
Now we distinguish cases according to the number of $\lambda_{i}$ that are zero:
If every $\lambda_{i}\neq 0$, then each $x_{i}$ belongs to $J_{1}$ (recall that this is the ideal of $\widehat{\mathcal{O}}$ generated by the entries of the formal Saito-matrix). Moreover, this ideal is proper because $D$ is not a product at 0 (and, thus, $\widehat{\operatorname{Der}}_{g}\subset\mathfrak{m}\widehat{\operatorname{Der}}$ by statement $(3)$ Proposition 2.7), so $J_{1}=\widehat{\mathfrak{m}}$, $I_{1}=J_{1}^{c}=\mathfrak{m}$ and $D_{0}=\{0\}$. Therefore, $D$ is strongly Euler-homogeneous outside 0 but not at 0, which contradicts Theorem 6.3.
If three of the $\lambda_{i}$ are zero, say $\lambda_{1}=\lambda_{2}=\lambda_{3}=0$, then $\lambda_{4}\neq 0$ and $\partial_{4}(g)=0$, so $D$ is (formally and then convergently by Proposition 2.7) a product at 0, reaching a contradiction.
If exactly two of the $\lambda_{i}$ are zero, say $\lambda_{1}=\lambda_{2}=0$, $\lambda_{3},\lambda_{4}\neq 0$, then, as $\sigma(g)=0$, $g$ is of the form $\sum_{\lambda_{3}j+\lambda_{4}k=0}a_{jk}(x_{1},x_{2})x_{3}^{j}x_{4}^{k}$. Equivalently, $g=\sum_{\mu\in\mathbb{N}}a_{\mu}(x_{1},x_{2})x_{3}^{\mu p}x_{4}^{\mu q}$ where $p,q\geq 1$ are coprime and $\lambda_{3}p+\lambda_{4}q=0$ (if $p$ or $q$ vanishes, $g$ would be a product because it would not depend on all variables). By formal Saito’s criterion (cf. [dVR24, Corollary 3.4]), $g$ is a unit times the determinant of the Saito matrix, which belongs to $(x_{3},x_{4})$ as $\lambda_{1}=\lambda_{2}=0$. This implies that $a_{0}(x_{1},x_{2})=0$ and we can extract $x_{3}^{p}x_{4}^{q}$ as a common factor. But $g$ is reduced, so we must have $p=q=1$ and $\lambda_{3}+\lambda_{4}=0$. Thus, $\operatorname{tr}(\sigma)=0$ and we get a contradiction.
The only case left (and the hardest to discard) is that only one of the $\lambda_{i}$ is zero, say $\lambda_{1}=0$. Now, we have:
$(x_{2},x_{3},x_{4})\subset J_{1}\subset\tilde{J}_{1}\subset\sqrt{\tilde{J}_{1}}=\sqrt{(\alpha_{2},\alpha_{3},\alpha_{4})},$
where the last equality is due to Corollary 5.3. As $\operatorname{ht}\sqrt{(\alpha_{2},\alpha_{3},\alpha_{4})}=\operatorname{ht}(\alpha_{2},\alpha_{3},\alpha_{4})\leq 3$, the inclusions cannot be strict. So we deduce $(x_{2},x_{3},x_{4})=J_{1}=\sqrt{(\alpha_{2},\alpha_{3},\alpha_{4})}$.
Since $\sigma(g)=0$, $\sum_{i=1}^{4}\lambda_{i}\beta_{i}=0$ for any monomial $x^{\beta}$ appearing in $g$, so not all $\lambda_{i}$ can have the same sign (as $g$ is not a product, all variables must appear in some monomial of $g$). Let us assume without loss of generality that $\lambda_{2}>0$ and $\lambda_{3},\lambda_{4}<0$.
Let us also note that $\alpha_{i}$ is $\sigma$-homogeneous of degree $c_{i}$:
$c_{i}\alpha_{i}g=c_{i}\delta_{i}(g)=[\sigma,\delta_{i}](g)=\sigma(\alpha_{i}g)=\sigma(\alpha_{i})g,$
so $\sigma(\alpha_{i})=c_{i}\alpha_{i}$ for $i=2,3,4$.
As there exists some $\beta^{\prime}\geq 1$ such that $x_{2}^{\beta^{\prime}}\in(\alpha_{2},\alpha_{3},\alpha_{4})$, some $\alpha_{i}$, say $\alpha_{2}$, must have a monomial of the form $x_{2}^{\beta}$ for some $\beta\geq 1$ (otherwise, $(\alpha_{2},\alpha_{3},\alpha_{4})$ would be contained in $(x_{1},x_{3},x_{4})$ and clearly $x_{2}^{\beta^{\prime}}\not\in(x_{1},x_{3},x_{4})$). Then, $c_{2}=\beta\lambda_{2}>0$.
Let $A$ be the Saito matrix with respect to $\{\sigma,\delta_{2},\delta_{3},\delta_{4}\}$ and let $a_{ij}=\delta_{i}(x_{j})$ (the entries of $A$ corresponding to the $\delta_{i}$). As $\delta_{i}$ is $\sigma$-homogeneous of degree $c_{i}$ and $x_{j}$ is $\sigma$-homogeneous of degree $\lambda_{j}$, it follows that $a_{ij}$ is $\sigma$-homogeneous of degree $c_{i}+\lambda_{j}$.
Now, we claim:
  • Every $\sigma$-homogeneous formal power series $h$ of strictly positive degree (with respect to $\sigma$) belongs to $(x_{2})$.
    Proof: As $x_{1}$ has degree $\lambda_{1}=0$ and $x_{3},x_{4}$ have negative degree ($\lambda_{3}$ and $\lambda_{4}$, respectively), a monomial in which $x_{2}$ does not appear must necessarily have a non-positive degree, so cannot be part of $h$.
  • Every $\sigma$-homogeneous formal power series $h$ of degree zero belongs to $(x_{2})+\mathbb{C}[[x_{1}]]$.
    Proof: If $x_{2}$ does not appear in a monomial of degree zero, then neither $x_{3}$ nor $x_{4}$, having a strictly negative degree, can appear, so the monomial must be a power of $x_{1}$.
Since $g=v\cdot\det(A)$ for some formal unit $v$, it follows that $g\in(x_{2},x_{3},x_{4})$. But then $g$ must be a multiple of $x_{2}$, since its degree is zero and a power of $x_{1}$ cannot be a monomial of $g$. This implies $\operatorname{Der}_{g}\subset\operatorname{Der}_{x_{2}}=\langle\partial_{1},x_{2}\partial_{2},\partial_{3},\partial_{4}\rangle$ [GS, Lemma 3.4] and so $a_{i2}$ is a multiple of $x_{2}$ for all $i$. Thus, we can consider $\delta^{\prime}_{i}=\delta_{i}-\frac{a_{i2}}{\lambda_{2}x_{2}}\sigma$ that is still $\sigma$-homogeneous of degree $c_{i}$ (as $a_{i2}/x_{2}$ is $\sigma$-homogeneous of degree $c_{i}+\lambda_{2}-\lambda_{2}=c_{i}$) and verifies $\delta^{\prime}_{i}(g)=\alpha_{i}g$. It is clear that $\{\sigma,\delta^{\prime}_{2},\delta^{\prime}_{3},\delta^{\prime}_{4}\}$ is a basis of $\operatorname{Der}_{g}$ whose Saito matrix has the same determinant that $A$. Note that $\delta_{i}^{\prime}(x_{2})=\delta_{i}(x_{2})-\frac{a_{i2}}{\lambda_{2}x_{2}}\sigma(x_{2})=a_{i2}-a_{i2}=0$. For the sake of simplicity, this new basis will be denoted again by $\{\sigma,\delta_{2},\delta_{3},\delta_{4}\}$ and its Saito matrix by $A$. With this new notation, $a_{i2}=0$ for $i=2,3,4$. Thus, expanding the determinant by the second column we get $g=v\cdot\det(A)=v\cdot\lambda_{2}x_{2}h$, where $h=\left|\begin{array}[]{ccc}a_{21}&a_{23}&a_{24}\\ a_{31}&a_{33}&a_{34}\\ a_{41}&a_{43}&a_{44}\\ \end{array}\right|$.
Note that $\det(A)$ is $\sigma$-homogeneous of degree $\sum_{i=2}^{4}\lambda_{i}+\sum_{i=2}^{4}c_{i}$ because so are all of its summands, so the unit $v$ must also be $\sigma$-homogeneous. But a unit, having a non-zero constant term, can only be $\sigma$-homogeneous of degree 0. Thus, we deduce that $\det(A)$ is also $\sigma$-homogeneous of degree 0 and $\sum_{i=2}^{4}\lambda_{i}+\sum_{i=2}^{4}c_{i}=0$.
Since $f$ (and so $g$), is reduced, $h$ cannot be a multiple of $x_{2}$. By the previous discussion, $x_{2}h$ is $\sigma$-homogeneous of degree 0 and, thus, $h$ must be $\sigma$-homogeneous of degree $\lambda_{2}$. Therefore, it has at least a monomial of the form $x_{1}^{i}x_{3}^{j}x_{4}^{k}$ with $j\lambda_{3}+k\lambda_{4}=-\lambda_{2}$ (or, equivalently, $\lambda_{2}=j|\lambda_{3}|+k|\lambda_{4}|$).
As $a_{21}$ and $\alpha_{2}$ are $\sigma$-homogeneous of degree $c_{2}>0$, by the first claim, they must be a multiple of $x_{2}$. Let us write $a_{21}=x_{2}a^{\prime}_{21}$ and $\alpha_{2}=x_{2}\alpha^{\prime}_{2}$.
Now consider $a_{23}$, which is $\sigma$-homogeneous of degree $c_{2}+\lambda_{3}$ and belongs to $J_{1}=(x_{2},x_{3},x_{4})$. If $c_{2}+\lambda_{3}\geq 0$, then $a_{23}$ must also be a multiple of $x_{2}$ by the second claim. The same applies to $a_{24}$ if $c_{2}+\lambda_{4}\geq 0$. But $a_{23}$ and $a_{24}$ cannot be a multiple of $x_{2}$ at the same time, as this would imply that $h$ is also a multiple of $x_{2}$, so $c_{2}+\lambda_{3}<0$ or $c_{2}+\lambda_{4}<0$:
  • If $c_{2}+\lambda_{3}=\beta\lambda_{2}+\lambda_{3}<0$, then $|\lambda_{3}|>\beta\lambda_{2}$ and:
    $\lambda_{2}=j|\lambda_{3}|+k|\lambda_{4}|>j\beta\lambda_{2}+k|\lambda_{4}|\geq j\lambda_{2},$
    so $j=0$ and $\lambda_{2}=k|\lambda_{4}|$ (in particular, $k\neq 0$).
  • If $c_{2}+\lambda_{4}=\beta\lambda_{2}+\lambda_{4}<0$, then $|\lambda_{4}|>\beta\lambda_{2}$ and:
    $\lambda_{2}=j|\lambda_{3}|+k|\lambda_{4}|>j|\lambda_{3}|+k\beta\lambda_{2}\geq k\lambda_{2},$
    so $k=0$ and $\lambda_{2}=j|\lambda_{3}|$ (in particular, $j\neq 0$).
Therefore, one of the degrees is negative and the other one is non-negative. Without loss of generality, we may assume that $c_{2}+\lambda_{3}<0$ and $c_{2}+\lambda_{4}\geq 0$ (thus, $a_{24}$ is a multiple of $x_{2}$ by the previous discussion). Moreover, any monomial of $h$ in which $x_{2}$ does not appear is of the form $x_{1}^{i}x_{4}^{k}$ (recall that $j=0$) with $k$ fixed ($k=\lambda_{2}/|\lambda_{4}|$).
We proceed in the same way with $x_{3}$: belonging to $\sqrt{(\alpha_{2},\alpha_{3},\alpha_{4})}$, as before, there exists $\gamma\geq 1$ such that $x_{3}^{\gamma}$ is a monomial of some $\alpha_{i}$. It cannot be a monomial of $\alpha_{2}=x_{2}\alpha^{\prime}_{2}$, so we may assume that it is a monomial of $\alpha_{3}$. Thus, $c_{3}=\gamma\lambda_{3}$.
As $\sum_{i=2}^{4}\lambda_{i}+\sum_{i=2}^{4}c_{i}=0$, we deduce:
$c_{4}=$ $-\lambda_{2}-\lambda_{3}-\lambda_{4}-c_{2}-c_{3}$
$=$ $-(1+\beta)\lambda_{2}+(1+\gamma)|\lambda_{3}|+|\lambda_{4}|$
$>$ $\hskip 2.84544pt[-(1+\beta)+(1+\gamma)\beta]\lambda_{2}+|\lambda_{4}|$
$=$ $\hskip 2.84544pt(\gamma\beta-1)\lambda_{2}+|\lambda_{4}|\geq|\lambda_{4}|>0,$
where we have used that $c_{2}=\beta\lambda_{2}$, $c_{3}=\gamma\lambda_{3}$, $\lambda_{3},\lambda_{4}<0$ in the second equality, $|\lambda_{3}|>\beta\lambda_{2}$ in the following inequality and $\beta,\gamma\geq 1$ in the next one.
But $c_{4}$ is the degree of $a_{41}$ and $\alpha_{4}$, so, again by the first claim, they are both a multiple of $x_{2}$: $a_{41}=x_{2}a^{\prime}_{41}$ and $\alpha_{4}=x_{2}\alpha^{\prime}_{4}$.
On the other hand, $\partial_{2}(g)=v\lambda_{2}h+\lambda_{2}x_{2}\partial_{2}(v\cdot h)$. If we evaluate it at $x_{2}=0$ we get $\partial_{2}(g)|_{x_{2}=0}=\lambda_{2}v|_{x_{2}=0}\cdot h|_{x_{2}=0}$.
By Cramer’s rule we also have
$\partial_{2}(g)=v\left|\begin{array}[]{cccc}0&0&\lambda_{3}x_{3}&\lambda_{4}x_{4}\\ a_{21}&\alpha_{2}&a_{23}&a_{24}\\ a_{31}&\alpha_{3}&a_{33}&a_{34}\\ a_{41}&\alpha_{4}&a_{43}&a_{44}\\ \end{array}\right|=v\left|\begin{array}[]{cccc}0&0&\lambda_{3}x_{3}&\lambda_{4}x_{4}\\ x_{2}a^{\prime}_{21}&x_{2}\alpha^{\prime}_{2}&a_{23}&x_{2}a^{\prime}_{24}\\ a_{31}&\alpha_{3}&a_{33}&a_{34}\\ x_{2}a^{\prime}_{41}&x_{2}\alpha^{\prime}_{4}&a_{43}&a_{44}\\ \end{array}\right|.$
Finally, evaluating $\partial_{2}(g)$ at $x_{2}=0$:
$\partial_{2}(g)|_{x_{2}=0}=$ $v|_{x_{2}=0}\cdot\left|\begin{array}[]{cccc}0&0&\lambda_{3}x_{3}&\lambda_{4}x_{4}\\ x_{2}a^{\prime}_{21}&x_{2}\alpha^{\prime}_{2}&a_{23}&x_{2}a^{\prime}_{24}\\ a_{31}&\alpha_{3}&a_{33}&a_{34}\\ x_{2}a^{\prime}_{41}&x_{2}\alpha^{\prime}_{4}&a_{43}&a_{44}\\ \end{array}\right|_{x_{2}=0}=0.$
Thus, $\lambda_{2}v|_{x_{2}=0}\cdot h|_{x_{2}=0}=0$, which implies $h|_{x_{2}=0}=0$ because $v$ is a unit. But then $h$ is a multiple of $x_{2}$, which is a contradiction.
We conclude that $D$ must be strongly Euler-homogeneous at 0.
7 The case of linear free divisors
Recall that linear free divisors are those free divisors $D\subset\mathbb{C}^{n}$ that have a global basis of linear logarithmic derivations (i.e. derivations that only have linear part). By Saito’s criterion, these divisors are always defined by homogeneous polynomials of degree $n$, so there is always a strong Euler vector field at the origin: $\chi=\sum_{i=1}^{n}x_{i}\partial_{i}$. In [GMNS, Theorem 7.10], Granger-Mond-Nieto-Schulze prove:
Theorem 7.1.
Every linear free divisor in dimension $n\leq 4$ is locally quasihomogeneous and hence LCT and GLCT (Global Logarithmic Comparison Theorem) hold.
They conjecture that this is true in any dimension. However, the following is an example of a linear free divisor in dimension $5$ such that it is not even strongly Euler-homogeneous and that does not verify LCT:
Example 7.2.
Let $D=V(f)$ with $f=x(8x^{3}u-y(8tx^{2}-4xyz+y^{3}))$.
A basis of its logarithmic derivations is:
$\begin{array}[]{l}\delta_{1}=x\partial_{x}+y\partial_{y}+z\partial_{z}+t\partial_{t}+u\partial_{u},\\ \delta_{2}=-4x\partial_{x}+y\partial_{y}+6z\partial_{z}+11t\partial_{t}+16u\partial_{u},\\ \delta_{3}=x\partial_{y}+y\partial_{z}+z\partial_{t}+t\partial_{u},\\ \delta_{4}=2x\partial_{z}+y\partial_{t},\\ \delta_{5}=x\partial_{t}+y\partial_{u}.\par\end{array}$
Here, $\delta_{1}$ is the Euler vector field ($\delta_{1}(f)=5f$) and $\delta_{i}(f)=0$ for $i=2,\ldots,5$.
The extended Saito matrix is then:
$\tilde{A}=\left(\begin{array}[]{ccccc|c}x&y&z&t&u&-5\\ -4x&y&6z&11t&16u&0\\ 0&x&y&z&t&0\\ 0&0&2x&y&0&0\\ 0&0&0&x&y&0\\ \end{array}\right).$
We have that:
[…]
By Theorem 3.7, $D$ is strongly Euler-homogeneous if and only if $\tilde{D}_{0}=\varnothing$ and $\tilde{D}_{i}=D_{i-1}$ for $i=1,\ldots,3$, so we conclude that $D$ is not strongly Euler-homogeneous at any point of $\tilde{D}_{3}\setminus D_{2}$. However, $\dim D_{i}=i$ for $i=0,1,2$, so $D$ is a Koszul-free divisor (Theorem 4.6). Since every Koszul-free divisor satisfying LCT must be strongly Euler-homogeneous, this divisor cannot satisfy LCT.
Moreover, its $b$-function is
$b(s)=(s+1)^{3}(2s+1)^{2}(3s+2)(3s+4)(4s+3)(4s+5)(6s+5)(6s+7).$
In [SymmetryGS, Theorem 1.4 and Conjecture 1.5], M. Granger and M. Schulze prove that all reductive linear free divisors have $b$-functions with symmetric roots about 1-1 and conjecture that this is true for any linear free divisor. However, this $b$-function is not symmetric about 1-1 due to the factor $(2s+1)$. We deduce that $D$ is not reductive and that this conjecture is also false for non-reductive linear free divisors.
Nevertheless, an immediate consequence of Theorem 6.10 is that Conjecture 1.4 holds for linear free divisors in $n=5$:
Corollary 7.3.
Let $D$ be a linear free divisor in $\mathbb{C}^{5}$. If $D$ satisfies LCT, then $D$ is strongly Euler-homogeneous.
Proof.
Since we know the result is true for any free divisor in dimension $4$, we have that $D$ is strongly Euler-homogeneous outside $D_{0}$, where it is a product. But $D$ is linear, so $D_{0}$ is only the origin, where it is clearly strongly Euler-homogeneous. Thus, $D$ is strongly Euler-homogeneous.
Remark 7.4.
In fact, the same proof is valid to argue that if Conjecture 1.4 is true in dimension $n$, then it is true for linear free divisors in dimension $n+1$.
Acknowledgements
The author would like to thank his PhD advisors Luis Narváez Macarro and Alberto Castaño Domínguez for their support and guidance while writing this paper.